### v1

# Classification of abelian Schur groups I

Abstract.
A finite group $G$ is called a Schur group if every $S$-ring over $G$ is schurian, i.e. associated in a natural way with a subgroup of $\operatorname{Sym}(G)$ that contains all right translations of $G$. The list of all possible abelian Schur groups was obtained by Evdokimov, Kovács, and Ponomarenko in 2016. In two papers, we finish a classification of abelian Schur groups. In the present paper, we study schurity of several groups from the list. Namely, we prove that a direct product of the elementary abelian group of order 4 and a cyclic group, whose order is an odd prime power or a product of two distinct odd primes, is a Schur group and establish nonschurity of some other groups from the list.
Keywords: $S$-rings, Schur groups, permutation groups.
MSC: 05E30, 20B25.
The author was supported by the state contract of the Sobolev Institute of Mathematics (project number FWNF-2026-0011).
1. Introduction
Let $G$ be a finite group. A subring of the group ring $\mathbb{Z}G$ is called an $S$-ring (Schur ring) over $G$ if it is a free $\mathbb{Z}$-module spanned by a partition of $G$ closed under taking inverse and containing the identity element $e$ of $G$ as a class. The notion of an $S$-ring was introduced by Schur [30]. Using the $S$-ring approach, Schur proved that every primitive permutation group containing a regular cyclic subgroup of composite order is $2$-transitive which generalizes the Burnside theorem [2, p. 339]. The general theory of $S$-rings was developed by Wielandt (see [32, Chapter IV]). To date, $S$-rings are realized as a powerful tool for studying permutation groups containing regular subgroups, combinatorial Cayley objects, especially, isomorphisms of them, representations of finite groups. For more details on $S$-rings and their applications, we refer the readers to [3, Section 2.4] and [20].
To study permutation groups having regular subgroups, Schur used $S$-rings whose partition of the underlying group arises from the action of an appropriate permutation group (see Section 2.3 for the exact definition). As it was shown by Wielandt [32], not all $S$-rings can be constructed in such way. An $S$-ring is said to be schurian if it arises from the action of a suitable permutation group [25]. The schurity property is one of the crucial properties of $S$-rings closely related to the isomorphism problem for Cayley graphs (see [20, Section 5.3] for details).
A finite group $G$ is defined to be a Schur group if every $S$-ring over $G$ is schurian [25]. The main problem considered in the present paper is the following one.
Problem (Pöschel, 1974).
Determine all Schur groups.
In general, the above problem seems to be hard, in particular, because the number of $S$-rings over a given group is exponential in the order of the group. There are several results on schurity of nonabelian groups [21, 24, 26, 28, 29]. However, most of the results on the schurity problem are concerned with abelian groups. The first result on schurity of abelian groups was obtained in [25]. Namely, it was proved that cyclic $p$-groups of odd order are Schur. This result was extended to the case $p=2$ in [11]. In [16], a schurity of cyclic groups of order $pq$, where $p$ and $q$ are distinct primes, was proved. The above results were used later for solving the isomorphism problem for Cayley graphs over the corresponding groups [15, 16].
A complete classification of cyclic Schur groups was obtained in [9]. Namely, it was proved that a cyclic group of order $n$ is Schur if and only if $n$ belongs to one of the following families of integers:
$p^{k},~pq^{k},~2pq^{k},~pqr,~2pqr,$
where $p$, $q$, and $r$ are distinct primes and $k\geq 0$ is an integer.
Investigations of schurity of abelian groups was continued in [10]. At first, necessary and sufficient condition of schurity for elementary abelian groups was obtained in this paper. Namely, an elementary abelian noncyclic group of order $n$ is Schur if and only if
$n\in\{4,8,9,16,27,32\}.$
Further, it was shown that an abelian Schur group which is neither cyclic nor elementary abelian belongs to one of the explicitly given infinite families of groups. Given a positive integer $n$, the cyclic and elementary abelian groups of order $n$ are denoted by $C_{n}$ and $E_{n}$, respectively.
Theorem 1.1.
[10, Theorem 1.3] An abelian Schur group, which is neither cyclic nor elementary abelian, is isomorphic to a group from one of the following nine families:
  1. (1)
    $C_{2}\times C_{2^{k}}$, $C_{2p}\times C_{2^{k}}$, $E_{4}\times C_{p^{k}}$, $E_{4}\times C_{pq}$, $E_{16}\times C_{p}$,
  2. (2)
    $C_{3}\times C_{3^{k}}$, $C_{6}\times C_{3^{k}}$, $E_{9}\times C_{q}$, $E_{9}\times C_{2q}$,
where $p$ and $q$ are distinct primes, $p\neq 2$, and $k\geq 1$ is an integer.
In the same paper, it was proved that the groups $E_{4}\times C_{p}$, where $p$ is an odd prime, are Schur groups. A schurity property of the groups $C_{2}\times C_{2^{k}}$ and $C_{3}\times C_{3^{k}}$ was established in [21] and [27], respectively. In [23], it was checked that $E_{9}\times C_{q}$ is a Schur group for every prime $q$.
In two papers, we study the shurity problem for all groups from Theorem 1.1 whose schurity is unknown. In the present paper, we deal with the most laborious case, namely, with the groups $E_{4}\times C_{p^{k}}$ and $E_{4}\times C_{pq}$. We prove the following theorem.
Theorem 1.2.
The groups $E_{4}\times C_{p^{k}}$ and $E_{4}\times C_{pq}$, where $p$ and $q$ are odd primes and $k\geq 1$, are Schur groups.
To prove Theorem 1.2, we characterize all $S$-rings over the groups from this theorem (see Theorem 7.1). We prove that every nontrivial $S$-ring over each of these groups arises from the action of an appropriate automorphism group or can be obtained as a tensor or generalized wreath product with special additional properties. To establish these properties is one of the crucial points in the proof because they allow to verify a sufficient condition of schurity for a generalized wreath product.
Perhaps, the most surprising fact concerning the groups from Theorem 1.1 is that not all of them are Schur groups. This is demonstrated by two theorems below.
Theorem 1.3.
The groups $C_{2p}\times C_{2^{k}}$, where $p$ is an odd prime and $k\geq 3$, are not Schur groups.
Since the class of Schur groups is closed under taking subgroups, it suffices to prove Theorem 1.3 only for $k=3$. Note that the groups $E_{4}\times C_{p}$ and $C_{2p}\times C_{4}$, where $p$ is an odd prime, are Schur groups. A schurity of the first group follows from [10, Theorem 1.4] as it was mentioned before, whereas a schurity of the second group will be verified in the subsequent paper together with a schurity of other abelian groups of order $8p$.
One more family of non-Schur groups from Theorem 1.1 is provided by the theorem below.
Theorem 1.4.
Let $p$ be an odd prime. The group $E_{16}\times C_{p}$ is a Schur group if and only if $p=3$.
In the proofs of Theorems 1.3 and 1.4, we construct new nonschurian $S$-rings over the groups $C_{2p}\times C_{8}$ and $E_{16}\times C_{p}$, respectively. These nonschurian $S$-rings are generalized wreath products of $S$-rings over proper subgroups of the above groups (see Sections 9 and 10). The nonschurian $S$-ring over $C_{2p}\times C_{8}$ from Section 9 in case $p=3$ was found by computer calculations using [14] and it is listed in [33]. The nonschurian $S$-ring over $E_{16}\times C_{p}$ from Section 10 in case $p=5$ was also found using [14].
We finish the introduction with a brief outline of the paper. Section 2 contains a necessary background of $S$-rings, especially, of the multiplier theorems, isomorphisms, and minimal $S$-rings. In Section 3, we provide an information on three constructions of $S$-rings, namely, tensor, generalized wreath, and star products. Several required facts on dual $S$-rings over the group of all irreducible complex characters of an abelian group are given in Section 4. Recent results on $S$-rings over an abelian group having a Sylow subgroup of prime order are provided in Section 5. In Section 6, we give a background of $S$-rings over cyclic groups, especially, over cyclic $p$-groups. In Section 7, we obtain a description of $S$-rings over the groups $E_{4}\times C_{p^{k}}$ and $E_{4}\times C_{pq}$ (Theorem 7.1). In Sections 8, 9, and 10, we prove Theorems 1.21.3 and 1.4, respectively.
Notation.
The identity element and the set of all nonidentity elements of a group $G$ are denoted by $e$ and $G^{\#}$, respectively.
The projections of $X\subseteq G_{1}\times G_{2}$ to $G_{1}$ and $G_{2}$ are denoted by $X_{G_{1}}$ and $X_{G_{2}}$, respectively.
If $X\subseteq G$, then the element $\sum_{x\in X}{x}$ of the group ring $\mathbb{Z}G$ is denoted by $\underline{X}$.
The set $\{x^{-1}:~x\in X\}$ is denoted by $X^{-1}$.
The subgroup of $G$ generated by $X$ is denoted by $\langle X\rangle$; we also set $\operatorname{rad}(X)=\{g\in G:\ gX=Xg=X\}$.
The set $\{(g,xg):~x\in X,~g\in G\}\subseteq G^{2}$ is denoted by $r(X)$.
If $m\in\mathbb{Z}$, then the set $\{x^{m}:~x\in X\}$ is denoted by $X^{(m)}$.
If $G$ is abelian and $m\in\mathbb{Z}$ coprime to $|G|$, then the automorphism of $G$ which maps every $g\in G$ to $g^{m}$ is denoted by $\sigma_{m}$.
The group of all permutations of a set $\Omega$ is denoted by $\operatorname{Sym}(\Omega)$.
The subgroup of $\operatorname{Sym}(G)$ induced by all right multiplications of $G$ is denoted by $G_{r}$.
If $f\in\operatorname{Sym}(\Omega)$ and $\Delta\subseteq\Omega$ is $f$-invariant, then the permutation of $\Delta$ induced by $f$ is denoted $f^{\Delta}$.
For a set $\Delta\subseteq\operatorname{Sym}(G)$ and a section $S=U/L$ of $G$ we set
$\Delta^{S}=\{f^{S}:~f\in\Delta,~S^{f}=S\},$
where $S^{f}=S$ means that $f$ permutes the $L$-cosets in $U$ and $f^{S}$ denotes the bijection of $S$ induced by $f$.
If $K\leq\operatorname{Sym}(\Omega)$ and $\alpha\in\Omega$, then the set of all orbits of $K$ on $\Omega$ and the stabilizer of $\alpha$ in $K$ are denoted by $\operatorname{Orb}(K,\Omega)$ and $K_{\alpha}$, respectively.
The set of all positive divisors of $n\in\mathbb{Z}$ is denoted by $[n]$.
Given a cyclic group $G$ of order $n$ and $l\in[n]$, the subgroup of $G$ of order $l$ is denoted by $G_{l}$ and the set of all generators of $G_{l}$ is denoted by $G_{l}^{*}$.
If $X\subseteq H\times G$, where $G$ is a cyclic group of order $n$, $h\in H$, and $l\in[n]$, then put $X_{h}=h^{-1}X\cap G$ and $X_{h,l}=h^{-1}X\cap G_{l}^{*}$.
2. $S$-rings
In this section, we provide a necessary background of $S$-rings. In general, we follow [10, 21, 27], where most of the material is contained.
2.1. Definitions and basic facts
A subring $\mathcal{A}\subseteq\mathbb{Z}G$ is called an $S$-ring (a Schur ring) over $G$ if there exists a partition $\mathcal{S}=\mathcal{S}(\mathcal{A})$ of $G$ such that:
$(1)$ $\{e\}\in\mathcal{S}$;
$(2)$ if $X\in\mathcal{S}$, then $X^{-1}\in\mathcal{S}$;
$(3)$ $\mathcal{A}=\operatorname{Span}_{\mathbb{Z}}\{\underline{X}:\ X\in\mathcal{S}\}$.
The elements of $\mathcal{S}$ are called the basic sets of $\mathcal{A}$. The number of basic sets is called the rank of $\mathcal{A}$ and denoted by $\operatorname{rk}(\mathcal{A})$. Clearly, $\mathbb{Z}G$ is an $S$-ring. The partition $\{\{e\},G^{\#}\}$ defines the $S$-ring $\mathcal{T}_{G}$ of rank $2$ over $G$. It is easy to check that if $X,Y\in\mathcal{S}(\mathcal{A})$, then $XY\in\mathcal{S}(\mathcal{A})$ whenever $|X|=1$ or $|Y|=1$. The $S$-ring $\mathcal{A}$ is called symmetric if $X=X^{-1}$ for every $X\in\mathcal{S}(\mathcal{A})$ and antisymmetric if for every $X\in\mathcal{S}(\mathcal{A})$ the equality $X=X^{-1}$ implies that $X=\{e\}$.
Let $X,Y\in\mathcal{S}$. If $Z\in\mathcal{S}$, then the number of distinct representations of $z\in Z$ in the form $z=xy$ with $x\in X$ and $y\in Y$ does not depend on the choice of $z\in Z$. Denote this number by $c^{Z}_{XY}$. One can see that $\underline{X}\cdot\underline{Y}=\sum_{Z\in\mathcal{S}(\mathcal{A})}c^{Z}_{XY}\underline{Z}$. Therefore the numbers $c^{Z}_{XY}$ are the structure constants of $\mathcal{A}$ with respect to the basis $\{\underline{X}:\ X\in\mathcal{S}\}$.
A set $T\subseteq G$ is called an $\mathcal{A}$-set if $\underline{T}\in\mathcal{A}$. If $T$ is an $\mathcal{A}$-set, then put
$\mathcal{S}(\mathcal{A})_{T}=\{X\in\mathcal{S}(\mathcal{A}):~X\subseteq T\}.$
A subgroup $H\leq G$ is called an $\mathcal{A}$-subgroup if $H$ is an $\mathcal{A}$-set. The $S$-ring $\mathcal{A}$ is called primitive if there is no a nontrivial proper $\mathcal{A}$-subgroup of $G$ and imprimitive otherwise. One can verify that for every $\mathcal{A}$-set $X$, the groups $\langle X\rangle$ and $\operatorname{rad}(X)$ are $\mathcal{A}$-subgroups. The set of all $\mathcal{A}$-subgroups of $G$ is denoted by $\mathcal{H}(\mathcal{A})$. It is easy to see that if $X\subseteq\operatorname{rad}(Y)$, then
(1) $\underline{X}\cdot\underline{Y}=\underline{Y}\cdot\underline{X}=|X|\underline{Y}.$
Lemma 2.1.
[10, Lemma 2.1] Let $\mathcal{A}$ be an $S$-ring over a group $G$, $H$ an $\mathcal{A}$-subgroup of $G$, and $X\in\mathcal{S}(\mathcal{A})$. Then the number $|X\cap Hx|$ does not depend on $x\in X$.
Lemma 2.2.
[21, Theorem 2.6] Let $X$ be a basic set of an $S$-ring $\mathcal{A}$ over a group $G$. Suppose that $H\leq\operatorname{rad}(X\setminus H)$ for some subgroup $H$ of $G$ such that $X\cap H\neq\varnothing$ and $X\setminus H\neq\varnothing$. Then $\operatorname{rad}(X)\leq H$ and $X=\langle X\rangle\setminus\operatorname{rad}(X)$.
Let $\{e\}\leq L\unlhd U\leq G$. A section $U/L$ is called an $\mathcal{A}$-section if $U$ and $L$ are $\mathcal{A}$-subgroups. If $S=U/L$ is an $\mathcal{A}$-section, then the module
$\mathcal{A}_{S}=\operatorname{Span}_{\mathbb{Z}}\left\{\underline{X}^{\pi}:~X\in\mathcal{S}(\mathcal{A}),~X\subseteq U\right\},$
where $\pi:U\rightarrow U/L$ is the canonical epimorphism, is an $S$-ring over $S$.
2.2. Multiplier theorems
Let $G$ be an abelian group. If $X,Y\subseteq G$ are such that $Y^{(m)}=X$ for some $m\in\mathbb{Z}$ coprime to $|G|$, then we say that $X$ and $Y$ are rationally conjugate. In this case, if $X\in\operatorname{Orb}(K,G)$ for some $K\leq\operatorname{Aut}(G)$, then $Y\in\operatorname{Orb}(K,G)$. The following two statements are known as the first and second Schur theorems on multipliers (see [32, Theorem 23.9, (a)-(b)]).
Lemma 2.3.
Let $\mathcal{A}$ be an $S$-ring over an abelian group $G$. Then $X^{(m)}\in\mathcal{S}(\mathcal{A})$ for every $X\in\mathcal{S}(\mathcal{A})$ and every $m\in\mathbb{Z}$ coprime to $|G|$.
Lemma 2.4.
Let $\mathcal{A}$ be an $S$-ring over an abelian group $G$, $p$ a prime divisor of $|G|$, and $H=\{g\in G:g^{p}=e\}$. Then for every $\mathcal{A}$-set $X$, the set
$X^{[p]}=\{x^{p}:x\in~X,~|X\cap Hx|\not\equiv 0\mod p\}$
is an $\mathcal{A}$-set.
Below, we deduce from Lemma 2.3 one more lemma which will be used further in the proof of Theorem 1.2.
Lemma 2.5.
Let $G=H\times D$, where $H$ is an abelian group and $D\cong C_{n}$ with $n$ coprime to $|H|$, $\mathcal{A}$ an $S$-ring over $G$, and $X\in\mathcal{S}(\mathcal{A})$. Then there is $K\leq\operatorname{Aut}(D)$ such that every nonempty set $X_{h,l}=h^{-1}X\cap D_{l}^{*}$, where $h\in H$ and $l\in[n]$, is an orbit of $K$.
Proof.
Let $K$ be the setwise stabilizer of $X$ in the group $M\leq\operatorname{Aut}(G)$ such that $M^{H}$ is trivial and $M^{D}=\operatorname{Aut}(D)$. By the definition of $K$, each set $X_{h,l}$ is $K$-invariant. On the other hand, for $x,y\in X_{h,l}$, there is an integer $m$ coprime to $n$ such that $x^{m}=y$ and $m\equiv 1\mod|H|$ and hence $(hx)^{m}=hy$. The latter implies that $X^{(m)}\cap X\neq\varnothing$. So $X^{(m)}=X$ by Lemma 2.3. Therefore $\sigma_{m}\in K$. Thus, $K$ is transitive on each $X_{h,l}$ and consequently $X_{h,l}$ is an orbit of $K$. ∎
2.3. Isomorphisms and schurity
Put $\mathcal{R}(\mathcal{A})=\{r(X):~X\in\mathcal{S}(\mathcal{A})\}$. Let $\mathcal{A}^{\prime}$ be an $S$-ring over a group $G^{\prime}$. A bijection $f$ from $G$ to $G^{\prime}$ is called a (combinatorial) isomorphism from $\mathcal{A}$ to $\mathcal{A}^{\prime}$ if $\mathcal{R}(\mathcal{A})^{f}=\mathcal{R}(\mathcal{A}^{\prime})$, where $\mathcal{R}(\mathcal{A})^{f}=\{r(X)^{f}:~X\in\mathcal{S}(\mathcal{A})\}$ and $r(X)^{f}=\{(g^{f},h^{f}):~(g,h)\in r(X)\}$. If there exists an isomorphism from $\mathcal{A}$ to $\mathcal{A}^{\prime}$, then $\mathcal{A}$ and $\mathcal{A}^{\prime}$ are said to be isomorphic.
A bijection $f\in\operatorname{Sym}(G)$ is defined to be a (combinatorial) automorphism of $\mathcal{A}$ if $r(X)^{f}=r(X)$ for every $X\in\mathcal{S}(\mathcal{A})$. The set of all automorphisms of $\mathcal{A}$ forms the group called the automorphism group of $\mathcal{A}$ and denoted by $\operatorname{Aut}(\mathcal{A})$. One can see that $\operatorname{Aut}(\mathcal{A})\geq G_{r}$. If $f\in\operatorname{Aut}(\mathcal{A})$, then
(2) $(Xy)^{f}=Xy^{f}$
for every $X\in\mathcal{S}(\mathcal{A})$ and $y\in G$. If $H$ is an $\mathcal{A}$-subgroup of $G$, then the set of all right $H$-cosets is an imprimitivity system of $\operatorname{Aut}(\mathcal{A})$.
Let $K$ be a subgroup of $\operatorname{Sym}(G)$ containing $G_{r}$. Schur proved in [30] that the $\mathbb{Z}$-submodule
$V(K,G)=\operatorname{Span}_{\mathbb{Z}}\{\underline{X}:~X\in\operatorname{Orb}(K_{e},G)\}$
is an $S$-ring over $G$. An $S$-ring $\mathcal{A}$ over $G$ is called schurian if $\mathcal{A}=V(K,G)$ for some $K\leq\operatorname{Sym}(G)$ with $K\geq G_{r}$. The group $G$ is called a Schur group if every $S$-ring over $G$ is schurian.
If $S$ is an $\mathcal{A}$-section, then $\mathcal{A}_{S}=V(K^{S},S)$. This yields that $\mathcal{A}_{S}$ is also schurian. Therefore a section (in particular, a subgroup) of a Schur group is also a Schur group.
Clearly, $\operatorname{Aut}(\mathcal{A})\geq K$. One can verify that $\mathcal{A}$ is schurian if and only if
$\mathcal{A}=V(\operatorname{Aut}(\mathcal{A}),G)$
or, equivalently, $\mathcal{S}(\mathcal{A})=\operatorname{Orb}(\operatorname{Aut}(\mathcal{A})_{e},G)$. It is easy to see that $\mathcal{T}_{G}=V(K,G)$ for every $2$-transitive group $K\leq\operatorname{Sym}(G)$ containing $G_{r}$.
The $S$-ring $\mathcal{A}$ is said to be normal if $G_{r}$ is normal in $\operatorname{Aut}(\mathcal{A})$ or, equivalently, $\operatorname{Aut}(\mathcal{A})\leq\operatorname{Hol}(G)=G_{r}\rtimes\operatorname{Aut}(G)$.
Let $K\leq\operatorname{Aut}(G)$. Then $\operatorname{Orb}(K,G)$ forms a partition of $G$ that defines an $S$-ring $\mathcal{A}$ over $G$. In this case, $\mathcal{A}$ is called cyclotomic and denoted by $\operatorname{Cyc}(K,G)$. One can see that $\mathcal{A}=V(G_{r}K,G)$. So every cyclotomic $S$-ring is schurian. If $S$ is an $\mathcal{A}$-section, then $\mathcal{A}_{S}=\operatorname{Cyc}(K^{S},G)$, i.e. $\mathcal{A}_{S}$ is also cyclotomic. Note that a normal $S$-ring can be nonschurian and hence noncyclotomic as well as a cyclotomic $S$-ring can be nonnormal. However, if a normal $S$-ring is schurian, then it is also cyclotomic.
2.4. Minimal $S$-rings
Two permutation groups $K_{1}$ and $K_{2}$ on a set $\Omega$ are called $2$-equivalent if $\operatorname{Orb}(K_{1},\Omega^{2})=\operatorname{Orb}(K_{2},\Omega^{2})$ (here we assume that $K_{1}$ and $K_{2}$ act on $\Omega^{2}$ componentwise). In this case, we write $K_{1}\approx_{2}K_{2}$. The relation $\approx_{2}$ is an equivalence relation on the set of all subgroups of $\operatorname{Sym}(\Omega)$. Every $\approx_{2}$-equivalence class has the largest element with respect to inclusion. Given $K\leq\operatorname{Sym}(\Omega)$, this element from the class containing $K$ is called the $2$-closure of $K$ and denoted by $K^{(2)}$. It should be noted that an $\approx_{2}$-equivalence class may have several minimal with respect to inclusion elements (see Example $1$).
Let $K\leq\operatorname{Sym}(G)$ with $K\geq G_{r}$ and $\mathcal{A}=V(K,G)$. Then $K^{(2)}=\operatorname{Aut}(\mathcal{A})$ and $V(K^{\prime},G)=\mathcal{A}$ if and only if $K^{\prime}\approx_{2}K$ for every $K^{\prime}\leq\operatorname{Sym}(G)$ with $K^{\prime}\geq G_{r}$. A schurian $S$-ring $\mathcal{A}$ over $G$ is said to be $2$-minimal if
$\{K\leq\operatorname{Sym}(G):~K\geq G_{r},~K\approx_{2}\operatorname{Aut}(\mathcal{A})\}=\{\operatorname{Aut}(\mathcal{A})\}.$
The set of all minimal with respect to inclusion permutation groups $K^{\prime}\leq\operatorname{Sym}(G)$ such that $K^{\prime}\geq G_{r}$ and $V(K^{\prime},G)=\mathcal{A}$ is denoted by $\mathcal{K}^{\min}(\mathcal{A})$.
Example 1. This example demonstrates that $\mathcal{K}^{\min}(\mathcal{A})$ may contain more than one element. Let $p\geq 5$ be an odd prime, $G\cong C_{p}$, and $\mathcal{A}=\mathcal{T}_{G}$. Let $K_{1}=\operatorname{Hol}(G)$. Clearly, $K_{1}\geq G_{r}$ and $\mathcal{A}=V(K_{1},G)$. One can see that the one-point stabilizer $(K_{1})_{e}=\operatorname{Aut}(G)$ acts regularly on $G^{\#}$. So $K_{1}\in\mathcal{K}^{\min}(\mathcal{A})$.
Let $K_{2}=\operatorname{Alt}(G)$. Since $p$ is odd, every element from $G_{r}$ is an even permutation of $G$ and hence $K_{2}\geq G_{r}$. The group $K_{2}$ is (p2)(p-2)-transitive. Together with $p\geq 5$, this implies that $K_{2}$ is $2$-transitive and consequently $\mathcal{A}=V(K_{2},G)$. As $p$ is odd, a generator of $\operatorname{Aut}(G)$ is an odd permutation of $G$. So $K_{1}\nleq K_{2}$. Therefore $K_{2}^{\prime}\neq K_{1}$, where $K_{2}^{\prime}\in\mathcal{K}^{\min}(\mathcal{A})$ with $K_{2}^{\prime}\leq K_{2}$. Thus, $|\mathcal{K}^{\min}(\mathcal{A})|\geq 2$.
The next lemma and remark after it easily follow from computer calculations using [14].
Lemma 2.6.
Let $\mathcal{A}$ be an $S$-ring over a group $G$. If $|G|\leq 3$ or $|G|=4$ and $\mathcal{A}\neq\mathcal{T}_{G}$, then $\mathcal{A}$ is $2$-minimal.
Remark 2.7.
Let $|G|=4$. Then $\operatorname{Aut}(\mathcal{T}_{G})=\operatorname{Sym}(G)\cong S_{4}$ and $\mathcal{T}_{G}=V(\operatorname{Sym}(G),G)=V(\operatorname{Alt}(G),G)$. This implies that $\mathcal{T}_{G}$ is not $2$-minimal. The group $\operatorname{Alt}(G)\cong A_{4}$ of order $12$ is a unique proper subgroup of $\operatorname{Aut}(\mathcal{T}_{G})=\operatorname{Sym}(G)$ which is $2$-equivalent to $\operatorname{Aut}(\mathcal{T}_{G})$ and hence $\operatorname{Alt}(G)\in\mathcal{K}^{\min}(\mathcal{T}_{G})$.
Two groups $K_{1},K_{2}\leq\operatorname{Aut}(G)$ are said to be Cayley equivalent if $\operatorname{Orb}(K_{1},G)=\operatorname{Orb}(K_{2},G)$. In this case, we write $K_{1}\approx_{\operatorname{Cay}}K_{2}$. If $\mathcal{A}=\operatorname{Cyc}(K,G)$ for some $K\leq\operatorname{Aut}(G)$, then $\operatorname{Aut}_{G}(\mathcal{A})=\operatorname{Aut}(\mathcal{A})\cap\operatorname{Aut}(G)$ is the largest group which is Cayley equivalent to $K$. A cyclotomic $S$-ring $\mathcal{A}$ over $G$ is called Cayley minimal if
$\{K\leq\operatorname{Aut}(G):~K\approx_{\operatorname{Cay}}\operatorname{Aut}_{G}(\mathcal{A})\}=\{\operatorname{Aut}_{G}(\mathcal{A})\}.$
It is easy to see that $\mathbb{Z}G$ is $2$-minimal and Cayley minimal.
3. Tensor, generalized wreath, and star products
This section contains a necessary information on three main constructions for producing $S$-rings.
3.1. Tensor product
Let $\mathcal{A}$ be an $S$-ring over a group $G$. Suppose that $G_{1}$ and $G_{2}$ are $\mathcal{A}$-subgroups such that $G=G_{1}\times G_{2}$. The $S$-ring $\mathcal{A}$ is defined to be a tensor product of $\mathcal{A}_{G_{1}}$ and $\mathcal{A}_{G_{2}}$ if
$\mathcal{S}(\mathcal{A})=\mathcal{S}(\mathcal{A}_{G_{1}})\otimes\mathcal{S}(\mathcal{A}_{G_{2}})=\{X_{1}\times X_{2}:~X_{1}\in\mathcal{S}(\mathcal{A}_{G_{1}}),~X_{2}\in\mathcal{S}(\mathcal{A}_{G_{2}})\}.$
In this case, we write $\mathcal{A}=\mathcal{A}_{G_{1}}\otimes\mathcal{A}_{G_{2}}$. The tensor product is called nontrivial if $\{e\}<G_{1}<G$ and $\{e\}<G_{2}<G$ and trivial otherwise.
Let $\mathcal{A}$ be an $S$-ring over a group $G$ and $L$ an $\mathcal{A}$-subgroup of $G$. We say that $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}$ if there exists an $\mathcal{A}$-subgroup $U$ such that $G=L\times U$ and $\mathcal{A}=\mathcal{A}_{L}\otimes\mathcal{A}_{U}$.
It is easy to check that
(3) $\operatorname{Aut}(\mathcal{A}_{G_{1}}\otimes\mathcal{A}_{G_{2}})=\operatorname{Aut}(\mathcal{A}_{G_{1}})\otimes\operatorname{Aut}(\mathcal{A}_{G_{2}}).$
The lemma below immediately follows from Eq. (3).
Lemma 3.1.
A tensor product of two $S$-rings is schurian if and only if each of the operands so is.
If $\mathcal{A}_{G_{1}}=\operatorname{Cyc}(K_{1},G_{1})$ and $\mathcal{A}_{G_{2}}=\operatorname{Cyc}(K_{2},G_{2})$, then
(4) $\mathcal{A}_{G_{1}}\otimes\mathcal{A}_{G_{2}}=\operatorname{Cyc}(K_{1}\times K_{2},G_{1}\times G_{2}).$
Lemma 3.2.
[10, Lemma 2.3] Let $\mathcal{A}$ be an $S$-ring over an abelian group $G=G_{1}\times G_{2}$. Suppose that $G_{1}$ and $G_{2}$ are $\mathcal{A}$-subgroups. Then
  1. (1)
    $X_{G_{i}}\in\mathcal{S}(\mathcal{A})$ for all $X\in\mathcal{S}(\mathcal{A})$ and $i\in\{1,2\}$;
  2. (2)
    $\mathcal{A}\geq\mathcal{A}_{G_{1}}\otimes\mathcal{A}_{G_{2}}$, and the equality is attained whenever $\mathcal{A}_{G_{i}}=\mathbb{Z}G_{i}$ for some $i\in\{1,2\}$.
Lemma 3.3.
Let $K\leq\operatorname{Sym}(G)$ such that $K\geq G_{r}$ and $\mathcal{A}=V(K,G)$. Suppose that $\mathcal{A}=\mathcal{A}_{G_{1}}\otimes\mathcal{A}_{G_{2}}$ for some $\mathcal{A}$-subgroups $G_{1}$ and $G_{2}$ such that $G=G_{1}\times G_{2}$ and $K^{G_{i}}\in\mathcal{K}^{\min}(\mathcal{A}_{G_{i}})$ for some $i\in\{1,2\}$. Then $K=K^{G_{1}}\times K^{G_{2}}$.
Proof.
One can see that
$K\leq K^{G_{1}}\times K^{G_{2}}\leq\operatorname{Aut}(\mathcal{A}_{1})\times\operatorname{Aut}(\mathcal{A}_{2})\leq\operatorname{Aut}(\mathcal{A})$
and hence $K$ is a subdirect product of $K^{G_{1}}$ and $K^{G_{2}}$. This implies that there exist groups $K_{1}\trianglelefteq K^{G_{1}}$, $K_{2}\trianglelefteq K^{G_{2}}$, and an isomorphism $\psi$ from $K^{G_{1}}/K_{1}$ to $K^{G_{2}}/K_{2}$ such that
$K=\{(f_{1},f_{2})\in K^{G_{1}}\times K^{G_{2}}:~(K_{1}f_{1})^{\psi}=K_{2}f_{2}\}.$
Since $V(K,G)=\mathcal{A}_{G_{1}}\otimes\mathcal{A}_{G_{2}}$, we conclude that $\mathcal{A}_{G_{1}}=V(K_{1},G_{1})$ and $\mathcal{A}_{G_{2}}=V(K_{2},G_{2})$. By the condition of the lemma, $K^{G_{i}}\in\mathcal{K}^{\min}(\mathcal{A}_{G_{i}})$ which yields that $K_{i}=K^{G_{i}}$. Thus, $K=K^{G_{1}}\times K^{G_{2}}$ as required. ∎
Example 2. The following example demonstrates that the condition $K^{G_{i}}\in\mathcal{K}^{\min}(\mathcal{A}_{G_{i}})$ for some $i\in\{1,2\}$ in Lemma 3.3 is essential. Let $\mathcal{A}_{G_{1}}=\mathcal{T}_{G_{1}}$ and $\mathcal{A}_{G_{2}}=\mathcal{T}_{G_{2}}$, where $|G_{i}|=p$ is an odd prime for each $i\in\{1,2\}$, and
$K=\operatorname{Alt}(G_{1})\times\operatorname{Alt}(G_{2})\cup(\operatorname{Sym}(G_{1})\setminus\operatorname{Alt}(G_{1}))\times(\operatorname{Sym}(G_{2})\setminus\operatorname{Alt}(G_{2})).$
Then $K\geq\operatorname{Alt}(G_{1})\times\operatorname{Alt}(G_{2})\geq(G_{1})_{r}\times(G_{2})_{r}=G_{r}$ because $p$ is an odd prime, $\mathcal{A}=V(K,G)$, $K^{G_{1}}=\operatorname{Sym}(G_{1})$, $K^{G_{2}}=\operatorname{Sym}(G_{2})$, and $K\neq K^{G_{1}}\times K^{G_{2}}$. In this case, $K^{G_{i}}\notin\mathcal{K}^{\min}(\mathcal{A}_{G_{i}})$ for each $i\in\{1,2\}$ because $\operatorname{Alt}(G_{i})<K^{G_{i}}$ and $\operatorname{Alt}(G_{i})\approx_{2}K^{G_{i}}$.
The corollary below immediately follows from Lemma 3.3 and Eq. (3).
Corollary 3.4.
A tensor product of two $2$-minimal $S$-rings is $2$-minimal.
3.2. Generalized wreath product
Let $S=U/L$ be an $\mathcal{A}$-section of $G$. The $S$-ring $\mathcal{A}$ is called the $S$-wreath product or generalized wreath product of $\mathcal{A}_{U}$ and $\mathcal{A}_{G/L}$ if $L\trianglelefteq G$ and every basic set $X$ of $\mathcal{A}$ outside $U$ is a union of some $L$-cosets or, equivalently, $L\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus U}$. In this case, we write $\mathcal{A}=\mathcal{A}_{U}\wr_{S}\mathcal{A}_{G/L}$. The $S$-wreath product is called nontrivial if $L\neq\{e\}$ and $U\neq G$ and trivial otherwise. The construction of a generalized wreath product of $S$-rings was introduced in [4]. If $L=U$, then the $S$-wreath product coincides with the wreath product $\mathcal{A}_{L}\wr\mathcal{A}_{G/L}$ of $\mathcal{A}_{L}$ and $\mathcal{A}_{G/L}$. The $S$-ring $\mathcal{A}$ is said to be decomposable if $\mathcal{A}$ is a nontrivial $S$-wreath product for some $\mathcal{A}$-section $S$ and indecomposable otherwise.
Given a section $S=U/L$ of a group $G$ such that $L\trianglelefteq G$ and $S$-rings $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ over $U$ and $G/L$, respectively, such that $S$ is both an $\mathcal{A}_{1}$- and an $\mathcal{A}_{2}$-section, and $(\mathcal{A}_{1})_{S}=(\mathcal{A}_{2})_{S}$, there is a unique $S$-ring $\mathcal{A}$ over $G$ that is the $S$-wreath product with $\mathcal{A}_{U}=\mathcal{A}_{1}$ and $\mathcal{A}_{G/L}=\mathcal{A}_{2}$ (see [4]).
Lemma 3.5.
[6, Corollary 5.7] Let $\mathcal{A}$ be an $S$-ring over an abelian group $G$ and $S=U/L$ an $\mathcal{A}$-section. Suppose that $\mathcal{A}$ is the $S$-wreath product and $\mathcal{A}_{U}$ and $\mathcal{A}_{G/L}$ are schurian. Then $\mathcal{A}$ is schurian if and only if there exist two groups $K_{1}$ and $K_{0}$ such that
$K_{1}\geq U_{r},~K_{0}\geq(G/L)_{r},~K_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{U}),~K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{G/L}),~\text{and}~K_{0}^{S}=K_{1}^{S}.$
Lemma 3.6.
[21, Corollary 10.3] Under the hypothesis of Lemma 3.5, the $S$-ring $\mathcal{A}$ is schurian if $\mathcal{A}_{S}$ is $2$-minimal. In particular, $\mathcal{A}$ is schurian if $U=L$.
The lemma below for cyclic groups was proved in [9, Theorem 7.5]. In fact, the proof for abelian groups is the same. To make the text self-contained, we provide a proof here.
Lemma 3.7.
Under the hypothesis of Lemma 3.5, the $S$-ring $\mathcal{A}$ is schurian whenever $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{U}$ or $\mathcal{A}_{S}$ is $\otimes$-complemented in $\mathcal{A}_{G/L}$.
Proof.
Suppose that the first condition of the lemma holds, i.e. $\mathcal{A}_{U}=\mathcal{A}_{L}\otimes\mathcal{A}_{H}$ for some $\mathcal{A}_{U}$-subgroup $H$. Clearly, $\mathcal{A}_{S}\cong\mathcal{A}_{H}$. Eq. (3) implies that $\operatorname{Aut}(\mathcal{A}_{U})=\operatorname{Aut}(\mathcal{A}_{L})\times\operatorname{Aut}(\mathcal{A}_{H})$. Let $K_{0}=\operatorname{Aut}(\mathcal{A}_{G/L})$. Observe that $K_{0}^{S}\approx_{2}\operatorname{Aut}(\mathcal{A}_{S})$ because $\mathcal{A}_{G/L}$ is schurian. Since $\mathcal{A}_{S}\cong\mathcal{A}_{H}$, there is $K_{2}\leq\operatorname{Sym}(H)$ such that $K_{2}\geq H_{r}$ and $K_{2}^{S}=K_{0}^{S}$. Put $K_{1}=\operatorname{Aut}(\mathcal{A}_{L})\times K_{2}$. One can see that:
  1. (1)
    $K_{1}\geq U_{r}$ because $K_{2}\geq H_{r}$;
  2. (2)
    $K_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{U})$ because $K_{2}^{S}\approx_{2}\operatorname{Aut}(\mathcal{A}_{S})$ and $\operatorname{Aut}(\mathcal{A}_{U})=\operatorname{Aut}(\mathcal{A}_{L})\times\operatorname{Aut}(\mathcal{A}_{H})$;
  3. (3)
    $K_{1}^{S}=K_{2}^{S}=K_{0}^{S}$.
Therefore all the conditions of Lemma 3.5 hold for $K_{1}$ and $K_{0}$. Thus, $\mathcal{A}$ is schurian.
Now suppose that the second condition of the lemma holds, i.e. $\mathcal{A}_{G/L}=\mathcal{A}_{S}\otimes\mathcal{A}_{H}$ for some $\mathcal{A}_{G/L}$-subgroup $H$. Eq. (3) yields that $\operatorname{Aut}(\mathcal{A}_{G/L})=\operatorname{Aut}(\mathcal{A}_{S})\times\operatorname{Aut}(\mathcal{A}_{H})$. Let $K_{1}=\operatorname{Aut}(\mathcal{A}_{U})$. Note that $K_{1}^{S}\approx_{2}\operatorname{Aut}(\mathcal{A}_{S})$ because $\mathcal{A}_{U}$ is schurian and $K_{1}^{S}\geq S_{r}$ because $K_{1}\geq U_{r}$. Put $K_{0}=K_{1}^{S}\times\operatorname{Aut}(\mathcal{A}_{H})$. Then:
  1. (1)
    $K_{0}\geq(G/L)_{r}$ because $K_{1}^{S}\geq S_{r}$;
  2. (2)
    $K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{G/L})$ because $K_{1}^{S}\approx_{2}\operatorname{Aut}(\mathcal{A}_{S})$ and $\operatorname{Aut}(\mathcal{A}_{G/L})=\operatorname{Aut}(\mathcal{A}_{S})\times\operatorname{Aut}(\mathcal{A}_{H})$;
  3. (3)
    $K_{0}^{S}=K_{1}^{S}$.
Thus, all the conditions of Lemma 3.5 hold for $K_{1}$ and $K_{0}$ and consequently $\mathcal{A}$ is schurian. ∎
3.3. Star product
Let $L$ and $U$ be normal $\mathcal{A}$-subgroups. The $S$-ring $\mathcal{A}$ is called the star product of $\mathcal{A}_{L}$ and $\mathcal{A}_{U}$ if the following conditions hold:
$(1)$ each $X\in\mathcal{S}(\mathcal{A})$ with $X\subseteq(U\setminus L)$ is a union of some $(L\cap U)$-cosets;
$(2)$ for each $X\in\mathcal{S}(\mathcal{A})$ with $X\subseteq G\setminus(L\cup U)$ there exist $Y\in\mathcal{S}(\mathcal{A}_{L})$ and $Z\in\mathcal{S}(\mathcal{A}_{U})$ such that $X=YZ$.
In this case, we write $\mathcal{A}=\mathcal{A}_{L}\star\mathcal{A}_{U}$. The construction of a star product of $S$-rings was introduced in [12] and extended to association schemes under the name “crested product” in [1]. A star product is called nontrivial if $L\neq\{e\}$ and $U\neq G$ and trivial otherwise.
Let $L$ and $U$ be proper nontrivial normal $\mathcal{A}$-subgroups of $G$ and $\mathcal{A}=\mathcal{A}_{L}\star\mathcal{A}_{U}$. It follows immediately from the definitions that:
  1. (1)
    if $|L\cap U|=1$, then $\mathcal{A}=\mathcal{A}_{L}\otimes\mathcal{A}_{U}$;
  2. (2)
    if $|L\cap U|>1$, then $\mathcal{A}$ is the nontrivial $L/(L\cap U)$-wreath product.
4. Duality
Let $G$ be an abelian group. Denote by $\widehat{G}$ the group dual to $G$, i.e., the group of all irreducible complex characters of $G$. It is well known that $\widehat{G}\cong G$ and there is a uniquely determined lattice antiisomorphism between the subgroups of $G$ and $\widehat{G}$. The image of a subgroup $H$ of $G$ with respect to this antiisomorphism is denoted by $H^{\bot}$.
For any $S$-ring $\mathcal{A}$ over the group $G$, one can define the dual $S$-ring $\widehat{\mathcal{A}}$ over $\widehat{G}$ as follows: two irreducible characters of $G$ belong to the same basic set of $\widehat{\mathcal{A}}$ if they have the same value on each basic set of $\mathcal{A}$ (for the exact definition, we refer the readers to [7, 31]). One can verify that
$\operatorname{rk}(\widehat{\mathcal{A}})=\operatorname{rk}(\mathcal{A})$
and the $S$-ring dual to $\widehat{\mathcal{A}}$ is equal to $\mathcal{A}$. The following lemma collects some facts on the dual $S$-ring taken from [8, Section 2.3].
Lemma 4.1.
Let $G$ be an abelian group and $\mathcal{A}$ an $S$-ring over $G$. Then
  1. (1)
[…]
  2. (2)
    $\widehat{\mathcal{A}_{H}}=\widehat{\mathcal{A}}_{\widehat{G}/H^{\bot}}$ and $\widehat{\mathcal{A}_{G/H}}=\widehat{\mathcal{A}}_{H^{\bot}}$ for every $H\in\mathcal{H}(\mathcal{A})$;
  3. (3)
    $\mathcal{A}=\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ if and only if $\widehat{\mathcal{A}}=\widehat{\mathcal{A}_{1}}\otimes\widehat{\mathcal{A}_{2}}$;
  4. (4)
    $\mathcal{A}$ is the $U/L$-wreath product for some $\mathcal{A}$-section $U/L$ if and only if $\widehat{\mathcal{A}}$ is the $L^{\bot}/U^{\bot}$-wreath product;
  5. (5)
[…]
5. $S$-rings over $H\times C_{p}$
Throughout this section, $G=H\times P$, where $H$ is an abelian group and $P\cong C_{p}$ with prime $p$ coprime to $|H|$, and $\mathcal{A}$ is an $S$-ring over $G$. Let $H_{1}$ be a maximal $\mathcal{A}$-subgroup contained in $H$ and $P_{1}$ the least $\mathcal{A}$-subgroup containing $P$. Note that $H_{1}P_{1}$ is an $\mathcal{A}$-subgroup.
Lemma 5.1.
[17, Lemma 6.3] If $H_{1}\neq(H_{1}P_{1})_{p^{\prime}}$, the Hall $p^{\prime}$-subgroup of $H_{1}P_{1}$, then $\mathcal{A}_{H_{1}P_{1}}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$.
Lemma 5.2.
[22, Proposition 15] If $\mathcal{A}_{H_{1}P_{1}/H_{1}}\cong\mathbb{Z}C_{p}$, then $\mathcal{A}_{H_{1}P_{1}}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$.
Lemma 5.3.
Suppose that $H_{1}<H$. Then one of the following statements holds:
  1. (1)
    $H_{1}P_{1}=G$, $P_{1}<G$, and $\mathcal{A}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$;
  2. (2)
    $\mathcal{A}=\mathcal{A}_{H_{1}}\wr\mathcal{A}_{G/H_{1}}$ with $\mathcal{A}_{G/H_{1}}=\mathcal{T}_{G/H_{1}}$;
  3. (3)
    $\mathcal{A}$ is the nontrivial $(H_{1}P_{1})/P_{1}$-wreath product.
Proof.
From [10, Lemma 6.2] it follows that Statement $(2)$ of the lemma holds or $\mathcal{A}$ is the $(H_{1}P_{1})/P_{1}$-wreath product with $P_{1}<G$. If the above wreath product is nontrivial, then Statement $(3)$ of the lemma holds, whereas if it is trivial, i.e. if $H_{1}P_{1}=G$, then $H_{1}\neq(H_{1}P_{1})_{p^{\prime}}=G_{p^{\prime}}=H$. So $\mathcal{A}=\mathcal{A}_{H_{1}P_{1}}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$ by Lemma 5.1 and Statement $(1)$ of the lemma holds. ∎
Lemma 5.4.
If $p=2$, then $\mathcal{A}_{H_{1}P_{1}}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$.
Proof.
If $H_{1}\neq(H_{1}P_{1})_{p^{\prime}}$, then the lemma follows from Lemma 5.1. If $H_{1}=(H_{1}P_{1})_{p^{\prime}}$, then $(H_{1}P_{1})/H_{1}\cong C_{2}$ and hence $\mathcal{A}_{(H_{1}P_{1})/H_{1}}\cong\mathbb{Z}C_{2}$. So in this case, the lemma follows from Lemma 5.2. ∎
6. $S$-rings over cyclic groups
In this section, we provide several results on $S$-rings over cyclic groups required for the proofs of the main results. An $S$-ring over a cyclic group is said to be circulant.
6.1. General results
Lemma 6.1.
[32, Theorem 25.4] If $G$ is an abelian group of composite order having a cyclic Sylow subgroup, then $\mathcal{T}_{G}$ is the only primitive $S$-ring over $G$.
The description of $S$-rings over a cyclic group was obtained in [18, 19]. Below, we provide this description in a convenient for us form taken from [6, Theorem 4.1]. If $\mathcal{A}$ is an $S$-ring over a cyclic group $G$, then put $\operatorname{rad}(\mathcal{A})=\operatorname{rad}(X)$, where $X$ is a basic set of $\mathcal{A}$ containing a generator of $G$. Due to Lemma 2.3, all basic sets of $\mathcal{A}$ containing a generator of $G$ are rationally conjugate and hence have the same radical. Therefore $\operatorname{rad}(\mathcal{A})$ does not depend on the choice of $X$.
Lemma 6.2.
Let $\mathcal{A}$ be an $S$-ring over a cyclic group. Then the following statements hold:
  1. (1)
    $|\operatorname{rad}(\mathcal{A})|=1$ if and only if $\mathcal{A}$ is a tensor product of a normal cyclotomic $S$-ring with trivial radical and trivial $S$-rings;
  2. (2)
    $|\operatorname{rad}(\mathcal{A})|>1$ if and only if $\mathcal{A}$ is a nontrivial generalized wreath product.
Lemma 6.3.
[6, Theorem 4.2(1)] Every normal $S$-ring over a cyclic group is cyclotomic and, in particular, schurian.
Lemma 6.4.
Every cyclotomic $S$-ring over a cyclic group is Cayley minimal.
Proof.
Let $G$ be a cyclic group and $\mathcal{A}=\operatorname{Cyc}(K,G)$ for some $K\leq\operatorname{Aut}(G)$. Assume that there exists $K^{\prime}<K$ such that $\mathcal{A}=\operatorname{Cyc}(K^{\prime},G)$. Let $X$ be a basic set of $\mathcal{A}$ containing a generator $x$ of $G$. Clearly, the one-point stabilizer of $x$ in $\operatorname{Aut}(G)$ is trivial. Since $X\in\operatorname{Orb}(K,G)\cap\operatorname{Orb}(K^{\prime},G)$, we conclude that $|K|=|K_{x}||X|=|X|=|K^{\prime}_{x}||X|=|K^{\prime}|$, a contradiction to $K^{\prime}<K$. ∎
Lemma 6.5.
Every normal $S$-ring over a cyclic group is $2$-minimal.
Proof.
Let $\mathcal{A}$ be a normal $S$-ring over a cyclic group $G$. Clearly, $\operatorname{Aut}(\mathcal{A})\leq\operatorname{Hol}(G)$. As $\mathcal{A}$ is normal, $\mathcal{A}$ is schurian and hence $\mathcal{A}=\operatorname{Cyc}(\operatorname{Aut}(\mathcal{A})_{e},G)$ by Lemma 6.3. Assume that $\mathcal{A}$ is not $2$-minimal, i.e. there exists $K<\operatorname{Aut}(\mathcal{A})$ such that $K\geq G_{r}$ and $K\approx_{2}\operatorname{Aut}(\mathcal{A})$. Since $K<\operatorname{Aut}(\mathcal{A})\leq\operatorname{Hol}(G)$, we obtain $K_{e}\leq\operatorname{Aut}(G)$. This implies that $\mathcal{A}=\operatorname{Cyc}(K_{e},G)$. From Lemma 6.4 it follows that $K_{e}=\operatorname{Aut}(\mathcal{A})_{e}$ and consequently $K=G_{r}K_{e}=G_{r}\operatorname{Aut}(\mathcal{A})_{e}=\operatorname{Aut}(\mathcal{A})$, a contradiction to the assumption. ∎
Lemma 6.6.
Under the hypothesis of Lemma 3.5, suppose that $U$ and $G/L$ are cyclic. Then $\mathcal{A}$ is schurian whenever $|S|$ is prime.
Proof.
If $\mathcal{A}_{S}\neq\mathcal{T}_{S}$, then $\mathcal{A}_{S}$ is normal by Lemma 6.2. So $\mathcal{A}_{S}$ is $2$-minimal by Lemma 6.5 and $\mathcal{A}$ is schurian by Lemma 3.6.
Suppose that $\mathcal{A}_{S}=\mathcal{T}_{S}$. Then $\mathcal{A}_{S}=\operatorname{Cyc}(\operatorname{Aut}(S),S)=V(\operatorname{Hol}(S),S)$. Since $\operatorname{Aut}(\mathcal{A}_{U})$ and $\mathcal{A}_{G/L}$ are schurian, [6, Theorem 8.1(2)] implies that there exist groups $K_{1}\leq\operatorname{Aut}(\mathcal{A}_{U})$ and $K_{0}\leq\operatorname{Aut}(\mathcal{A}_{G/L})$ such that
$K_{1}\geq U_{r},~K_{0}\geq(G/L)_{r},~K_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{U}),~K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{G/L}),~K_{1}^{S}=K_{0}^{S}=\operatorname{Hol}(S).$
Thus, all the conditions of Lemma 3.5 hold for $K_{1}$ and $K_{0}$ and hence $\mathcal{A}$ is schurian. ∎
6.2. $S$-rings over cyclic $p$-groups
The lemma below can be found, e.g., in [23, Lemma 2.4].
Lemma 6.7.
Every $S$-ring over a cyclic group of prime order is cyclotomic.
The first statement of the lemma below immediately follows from [5, Lemma 5.1], whereas the second one follows from Lemma 6.2 and the first one.
Lemma 6.8.
Let $p$ be an odd prime and $G$ a cyclic $p$-group.
  1. (1)
    If $K\leq\operatorname{Aut}(G)$ and $|\operatorname{rad}(X)|=1$ for some $X\in\operatorname{Orb}(K,G)$ such that $\langle X\rangle=G$, then $|Y|=|K|\leq p-1$ and all the elements of $Y$ lie in pairwise distinct cosets by every proper subgroup of $\langle Y\rangle$ for every $Y\in\operatorname{Orb}(K,G)\setminus\{\{e\}\}$.
  2. (2)
    If $\mathcal{A}$ is a nontrivial $S$-ring over $G$ such that $|\operatorname{rad}(\mathcal{A})|=1$, then $\mathcal{A}=\operatorname{Cyc}(K,G)$ for some $K\leq\operatorname{Aut}(G)$ with $|K|$ dividing p1p-1 and $|X|=|K|$ for every $X\in\mathcal{S}(\mathcal{A})\setminus\{\{e\}\}$.
Lemma 6.9.
Let $p$ be an odd prime and $\mathcal{A}$ a cyclotomic $S$-ring over a cyclic $p$-group $G$ such that $|\operatorname{rad}(\mathcal{A})|\in\{1,p\}$. Then $|\operatorname{rad}(\mathcal{A}_{S})|=1$ for every $\mathcal{A}$-section $S\neq G$.
Proof.
Let $\mathcal{A}=\operatorname{Cyc}(K,G)$, where $K\leq\operatorname{Aut}(G)$. Since $G$ is a cyclic $p$-group of odd order, $\operatorname{Aut}(G)=A_{1}\times A_{2}$, where
$A_{1}\cong C_{p-1}~\text{and}~A_{2}=\{\sigma_{m}:~m\in\mathbb{Z},~m\equiv 1\mod~p\}\cong C_{|G|/p}.$
So $K=K_{1}\times K_{2}$, where $K_{1}=K\cap A_{1}$ and $K_{2}=K\cap A_{2}$. One can see that
$|K_{2}|=|\operatorname{rad}(\mathcal{A})|\in\{1,p\}.$
Therefore $|K_{2}|=1$ or $K_{2}=\{\sigma_{m}:~m\in\mathbb{Z},~m\equiv 1\mod~|G|/p\}$. This implies that $K^{G/L_{0}}$ and $K^{U_{0}}$, where $L_{0}$ and $U_{0}$ are the subgroups of $G$ of order and index $p$, respectively, are $p^{\prime}$-groups of order at most p1p-1. If $S=U/L$ is an $\mathcal{A}$-section and $S\neq G$, then $U\leq U_{0}$ or $L\geq L_{0}$. This yields that $\mathcal{A}_{S}=\operatorname{Cyc}((K^{G/L_{0}})^{S},S)$ or $\mathcal{A}_{S}=\operatorname{Cyc}((K^{U_{0}})^{S},S)$. As $|(K^{G/L_{0}})^{S}|\leq p-1$ and $|(K^{U_{0}})^{S}|\leq p-1$, we conclude that every basic set of $\mathcal{A}_{S}$ is of size at most p1p-1 and hence has a trivial radical. Thus, $|\operatorname{rad}(\mathcal{A}_{S})|=1$. ∎
Lemma 6.10.
Let $\mathcal{A}$ be a nontrivial generalized wreath product over a cyclic $p$-group $G$. Then $\mathcal{A}$ is the nontrivial $U/L$-wreath product for an $\mathcal{A}$-section $U/L$ of $G$ such that $L$ is the least nontrivial $\mathcal{A}$-subgroup and $|\operatorname{rad}(\mathcal{A}_{U})|=1$.
Proof.
Since $G$ is a cyclic $p$-group, there exists the least $\mathcal{A}$-subgroup $L$. Let $U$ be the $\mathcal{A}$-subgroup generated by all basic set of $\mathcal{A}$ with trivial radical. The statement of the lemma immediately follows from the definitions of $L$ and $U$. ∎
A subset of $G$ is said to be regular if it consists of elements of the same order. The lemma below can be deduced from the description of $S$-rings over cyclic $p$-groups [25]. To make the text self-contained, we provide the proof here.
Lemma 6.11.
Let $\mathcal{A}$ be an $S$-ring over a cyclic $p$-group $G$ and $X\in\mathcal{S}(\mathcal{A})$. Suppose that $X$ is not regular. Then $X=U\setminus L$ for some $\mathcal{A}$-subgroups $U>L$ and $\mathcal{A}=\mathcal{A}_{L}\wr\mathcal{A}_{G/L}=\mathcal{A}_{U}\wr\mathcal{A}_{G/U}$.
Proof.
Let $L=\operatorname{rad}(X)$. The image of $T\subseteq G$ under the canonical epimorphism from $G$ to $G/L$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muT\mkern-1.5mu}\mkern 1.5mu$. The $S$-ring $\mathcal{A}_{\langle\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\rangle}$ has a trivial radical. So $\mathcal{A}_{\langle\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\rangle}$ is trivial or cyclotomic by Lemma 6.2. As $X$ is not regular, $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu$ also is not regular. Therefore $\mathcal{A}_{\langle\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\rangle}$ can not be cyclotomic and hence it is trivial. Thus, $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\cup\{L\}$ is an $\mathcal{A}_{G/L}$-subgroup. Since $L=\operatorname{rad}(X)$, we conclude that $X=U\setminus L$ for some $\mathcal{A}$-subgroup $U$ as required.
Now let us prove the second part of the lemma. If $U=G$, then $\mathcal{A}$ is the $L/L$-wreath product (possibly, trivial) and the trivial $U/U$-wreath product. Further, we assume that $U<G$ and hence $\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu<\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu$. The latter implies that $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ is nontrivial. Note that $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu=(\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu)^{\#}$ is a nonregular basic set of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ with trivial radical. So $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ can not be cyclotomic and hence $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ is a nontrivial generalized wreath product by Lemma 6.2. In view of Lemma 6.10, $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ is the nontrivial $W/V$-wreath product, where $\operatorname{rad}(\mathcal{A}_{W})$ is trivial and $V$ is the least nontrivial $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$-subgroup. Since $\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)$ is trivial, we conclude that $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\subseteq W$. This yields that $\mathcal{A}_{W}$ is not cyclotomic. Due to Lemma 6.2, the $S$-ring $\mathcal{A}_{W}$ is trivial. Therefore $W=V=\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu$. Thus,
$\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}=\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}\wr\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu/\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}.$
In view of the above equality, to prove the lemma it is enough to show that $L\leq\operatorname{rad}(Y)$ for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus U}$. Assume the contrary that $N=\operatorname{rad}(Y)<L$ for some $Y\in\mathcal{S}(\mathcal{A})_{G\setminus U}$. The image of $T\subseteq G$ under the canonical epimorphism from $G$ to $G/N$ is denoted by $\widetilde{T}$. As $Y$ lies outside $U$, we conclude that $X\subseteq U\leq\langle Y\rangle$. So $\widetilde{X}$ is a nonregular basic set of $\mathcal{A}_{\langle\widetilde{Y}\rangle}$. This implies that $\mathcal{A}_{\langle\widetilde{Y}\rangle}$ can not be cyclotomic. Therefore $\mathcal{A}_{\langle\widetilde{Y}\rangle}$ is trivial by Lemma 6.2. On the other hand, $\widetilde{X}$ and $\widetilde{Y}$ are nontrivial basic sets of $\mathcal{A}_{\langle\widetilde{Y}\rangle}$, a contradiction to triviality of $\mathcal{A}_{\langle\widetilde{Y}\rangle}$. ∎
6.3. $S$-rings over some other cyclic groups
Lemma 6.12.
Every nontrivial $S$-ring over a cyclic group of order a product of two distinct primes is cyclotomic or a nontrivial wreath product.
Proof.
An $S$-ring over a group of prime order is cyclotomic by Lemma 6.7 and hence a tensor product of $S$-rings over groups of prime orders is cyclotomic by Eq. (4). Therefore the required follows from Lemma 6.2. ∎
Lemma 6.13.
Let $\mathcal{A}$ be a nontrivial $S$-ring over a cyclic group $G$ of order $2p^{k}$, where $p$ is an odd prime and $k\geq 1$. Suppose that $|\operatorname{rad}(\mathcal{A})|=1$. Then $\mathcal{A}$ is normal unless $\mathcal{A}=\mathbb{Z}L\otimes\mathcal{T}_{U}$, where $L,U\leq G$ are such that $|L|=2$ and $|U|=p^{k}$.
Proof.
Assume that $\mathcal{A}$ is nonnormal. Then $\mathcal{A}=\mathbb{Z}_{L}\otimes\mathcal{A}_{U}$ by Lemma 6.2. If $|\operatorname{rad}(\mathcal{A}_{U})|>1$, then $|\operatorname{rad}(\mathcal{A})|>1$, a contradiction to the assumption of the lemma. So $|\operatorname{rad}(\mathcal{A}_{U})|=1$. If $\mathcal{A}_{U}\neq\mathcal{T}_{U}$, then $\mathcal{A}_{U}$ is normal by Lemma 6.2 and consequently $\mathcal{A}$ so is by Eq. (3), a contradiction to the assumption of the lemma. Thus, $\mathcal{A}_{U}=\mathcal{T}_{U}$ and hence $\mathcal{A}=\mathbb{Z}_{L}\otimes\mathcal{T}_{U}$. ∎
7. $S$-rings over $E_{4}\times C_{n}$
The main goal of this section is to give a description of $S$-rings over the group $E_{4}\times C_{n}$, where $n\in\{p^{k},pq\}$, $p$ and $q$ are odd primes, and $k\geq 1$. We say that an $S$-ring $\mathcal{A}$ over $E_{4}\times C_{n}$ is dense if $E_{4}$ and $C_{n}$ are $\mathcal{A}$-subgroups and nondense otherwise.
Theorem 7.1.
Let $n\in\{p^{k},pq\}$, where $p$ and $q$ are odd primes and $k$ is a positive integer, and $\mathcal{A}$ a nontrivial $S$-ring over the group $G=H\times D$, where $H\cong E_{4}$ and $D\cong C_{n}$. Then $\mathcal{A}$ is cyclotomic, or a nontrivial tensor product, or a nontrivial $S$-wreath product for some $\mathcal{A}$-section $S=U/L$ and one of the following statements holds:
  1. (1)
    $|S|\leq 2$;
  2. (2)
    $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{U}$ or $\mathcal{A}_{S}$ is $\otimes$-complemented in $\mathcal{A}_{G/L}$;
  3. (3)
    $\mathcal{A}$ is nondense and at least one of the $S$-rings $\mathcal{A}_{U}$, $\mathcal{A}_{G/L}$ is a circulant cyclotomic $S$-ring with trivial radical;
  4. (4)
    $\mathcal{A}$ is dense, $U\geq H$, $\mathcal{A}_{U}$ is cyclotomic, $|\operatorname{rad}(\mathcal{A}_{U\cap D})|=1$ unless $n=3^{k}$ and $|\operatorname{rad}(\mathcal{A}_{U\cap D})|=3$, and $L\leq D$.
Observe that $U\cap D$ in Statement $(4)$ of Theorem 7.1 is cyclic and hence $\operatorname{rad}(\mathcal{A}_{U\cap D})$ is defined correctly. Before the proof of Theorem 7.1, we need some preparatory work. In the first subsection, we study a structure of a basic set of $\mathcal{A}$. The second and third subsections deal with the cases when $\mathcal{A}$ is nondense and when $\mathcal{A}$ is dense, respectively. In the fourth subsection, we prove Theorem 7.1. The fifth subsection contains some auxiliary statements which will be used in the proof of Theorem 1.2.
Throughout this section, $p$ and $q$ are odd primes, $k$ is a positive integer, $n\in\{p^{k},pq\}$, $D\cong C_{n}$, $P=D_{p}=\{g\in D:~|g|\in\{1,p\}\}$, $H\cong E_{4}$, $G=H\times D$, and $\mathcal{A}$ is an $S$-ring over $G$.
7.1. Structure of a basic set
Lemma 7.2.
Let $n=p^{k}$ and $X\in\mathcal{S}(\mathcal{A})$. Suppose that $X\nsubseteq D$. Then one of the following statements holds:
  1. (1)
    $P\leq\operatorname{rad}(X)$;
  2. (2)
    $\langle(X^{[p]})_{H}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $H$;
  3. (3)
    $X\cap D\neq\varnothing$ and $X\cup\{e\}\leq G$.
Proof.
Assume that $P\nleq\operatorname{rad}(X)$, i.e. Statement $(1)$ of the lemma does not hold. Then the union $X^{0}$ of all $hX_{h,l}$, $h\in H$, $l\in[n]$, such that $X_{h,l}\neq\varnothing$ and $P\nleq\operatorname{rad}(X_{h,l})$, is nonempty. As $D\cong C_{p^{k}}$, the condition $P\nleq\operatorname{rad}(X_{h,l})$ is equivalent to $|\operatorname{rad}(X_{h,l})|=1$. Lemma 2.5 implies that each $X_{h,l}$ is an orbit of some subgroup $K$ of $\operatorname{Aut}(D_{l})$. From Lemma 6.8(1) it follows that $|K|\leq p-1$ and hence
(5) $|X_{h,l}|\leq p-1$
for every $h\in H$ and $l\in[n]$ with $X_{h,l}\neq\varnothing$ and $P\nleq\operatorname{rad}(X_{h,l})$. Put $H_{0}=(X^{0})_{H}$ and $X^{i}=(X^{i-1})^{[p]}$ for $i\geq 1$. Due to the definition of $X^{0}$ and Eq. (5), we have $X^{1}=X^{[p]}$ and $X^{i}=\{x^{p}:~x\in X^{i-1}\}$. Therefore $(X^{i})_{H}=H_{0}$ for every $i\geq 1$, in particular, $(X^{1})_{H}=(X^{[p]})_{H}=H_{0}$. By Lemma 2.4, the set $X^{i}$ is an $\mathcal{A}$-set for every $i\geq 1$. So $X^{k}=H_{0}$ is an $\mathcal{A}$-set. Thus, $\langle H_{0}\rangle$ is an $\mathcal{A}$-subgroup. If $H_{0}\neq\{e\}$, then Statement $(2)$ of the lemma holds.
In view of the previous paragraph, we may assume that $H_{0}=\{e\}$ and hence
$X^{0}\subseteq X\cap D\neq\varnothing.$
Let $l=p^{j}$ be the greatest integer such that $X_{e,l}\neq\varnothing$ and $P\nleq\operatorname{rad}(X_{e,l})$. By the assumption of the lemma, $X\setminus D\neq\varnothing$. So
(6) $X\setminus D_{l}\neq\varnothing~\text{and}~P\leq\operatorname{rad}(X\setminus D_{l}).$
If $j\neq 1$, then $X^{j-1}$ is a nonemepty $\mathcal{A}$-set inside $P^{\#}$. Therefore $P=\langle X^{j-1}\rangle$ is an $\mathcal{A}$-subgroup. Now Lemma 2.1 and Eq. (6) imply that $P\leq\operatorname{rad}(X)$, a contradiction to the assumption. If $j=1$, then $X_{e,l}\subseteq P^{\#}$. Eq. (6) and Lemma 2.2 applied to $X$ and $P$ yield that $\operatorname{rad}(X)\leq P$ and $X=\langle X\rangle\setminus\operatorname{rad}(X)$. By the assumption, $P\nleq\operatorname{rad}(X)$ and consequently $\operatorname{rad}(X)$ is trivial. Thus, $X=\langle X\rangle\setminus\{e\}$ and Statement $(3)$ of the lemma holds. ∎
7.2. Nondense $S$-rings
Proposition 7.3.
Suppose that $H$ is not an $\mathcal{A}$-subgroup. Then $\mathcal{A}$ is trivial, or a nontrivial tensor product, or a nontrivial $U/L$-wreath product for some $\mathcal{A}$-section $U/L$ satisfying one of the following conditions:
  1. (1)
    $|U/L|\leq 2$;
  2. (2)
    $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{U}$;
  3. (3)
    $4\nmid|U|$ and $|L|\neq 2$.
Proof.
Let $n=pq$. Observe that the subgroup of $G$ of order $4p$ or the subgroup of $G$ of order $4q$ is not an $\mathcal{A}$-subgroup because otherwise $H$ is an $\mathcal{A}$-subgroup as an intersection of these subgroups which contradicts to the assumption of the proposition. Without loss of generality, we may assume that the subgroup of $G$ of order $4q$ is not an $\mathcal{A}$-subgroup. Then
(7) $|H_{1}|\in\{1,2,q,2q\},$
where $H_{1}$ is a maximal $\mathcal{A}$-subgroup whose order is not divisible by $p$.
Let $P_{1}$ be the least $\mathcal{A}$-subgroup containing $P\cong C_{p}$. One of the statements of Lemma 5.3 holds for $\mathcal{A}$, $H_{1}$, and $P_{1}$. If Statement $(1)$ of Lemma 5.3 holds, then $H_{1}P_{1}=G$ and $\mathcal{A}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$. If $H_{1}\cap P_{1}$ is trivial, then $\mathcal{A}=\mathcal{A}_{H_{1}}\otimes\mathcal{A}_{P_{1}}$ is a nontrivial tensor product and we are done. Otherwise $\mathcal{A}$ is the nontrivial $U/L=H_{1}/(H_{1}\cap P_{1})$-wreath product. Together with Eq. (7), this implies that $U/L$ satisfies Condition $(1)$ from Proposition 7.3 as desired or $|H_{1}|=2q$ and $|H_{1}\cap P_{1}|=2$. In the latter case, Lemma 6.12 yields that $\mathcal{A}_{H_{1}}=\mathcal{A}_{H_{1}\cap P_{1}}\otimes\mathcal{A}_{Q}$, where $Q$ is an $\mathcal{A}$-subgroup of order $q$, or $\mathcal{A}_{H_{1}}=\mathcal{A}_{H_{1}\cap P_{1}}\wr\mathcal{A}_{H_{1}/(H_{1}\cap P_{1})}$. In the first case, $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{U}$ and $U/L$ satisfies Condition $(2)$. In the second one, $\mathcal{A}$ is the $L/L=(H_{1}\cap P_{1})/(H_{1}\cap P_{1})$-wreath product and the section $L/L$ satisfies Condition $(1)$ .
If Statement $(2)$ of Lemma 5.3 holds, then $\mathcal{A}$ is trivial whenever $H_{1}$ is trivial and the $U/L=H_{1}/H_{1}$-wreath product satisfying Condition $(1)$ otherwise. If Statement $(3)$ of Lemma 5.3 holds, then $\mathcal{A}$ is the nontrivial $U/L=(H_{1}P_{1})/P_{1}$-wreath product. Due to Eq. (7), the section $U/L$ satisfies Condition $(1)$ or $(3)$.
Now let $n=p^{k}$. As in the previous paragraph, $P_{1}$ denotes the least $\mathcal{A}$-subgroup containing $P$. One of Statements (1)(1)-$(3)$ of Lemma 7.2 holds for every basic set of $\mathcal{A}$ outside $D$. We divide the rest of the proof into two cases depending on existence of a basic set outside $D$ satisfying Statement $(3)$ of Lemma 7.2.
Case 1. Suppose that there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus D}$ satisfying Statement $(3)$ of Lemma 7.2, i.e. $X\cup\{e\}$ is an $\mathcal{A}$-subgroup with $X\cap D\neq\varnothing$. Then $X\cup\{e\}=P_{1}$, $P_{1}\nleq D$, and $\mathcal{A}_{P_{1}}=\mathcal{T}_{P_{1}}$. If $P_{1}=G$, then $\mathcal{A}$ is trivial. Further, we assume that $G\setminus P_{1}\neq\varnothing$.
Observe that $Y\nsubseteq D$ for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$. Indeed, otherwise $\langle Y\rangle$ is an $\mathcal{A}$-subgroup inside $D$. So $\langle Y\rangle\geq P_{1}$ and hence
$D\geq\langle Y\rangle\geq P_{1}\nleq D,$
a contradiction. Thus, Lemma 7.2 holds for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$.
Statement $(3)$ of Lemma 7.2 can not hold for $Y\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$ because $P_{1}$ is the least $\mathcal{A}$-subgroup containing $P$. If Statement $(1)$ of Lemma 7.2 holds for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$, then $P_{1}\leq\operatorname{rad}(Y)$ and hence $\mathcal{A}$ is the $U/L=P_{1}/P_{1}$-wreath product and $U/L$ satisfies Condition $(1)$ from Proposition 7.3.
In view of the above paragraph, we may assume that there is $Y\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$ satisfying Statement $(2)$ of Lemma 7.2 and $P_{1}\nleq\operatorname{rad}(Y)$. Since $H$ is not an $\mathcal{A}$-subgroup, $A=\langle(Y^{[p]})_{H}\rangle$ is an $\mathcal{A}$-subgroup of order $2$. Clearly, $\mathcal{A}_{A}=\mathbb{Z}A$. Note that $A\cap P_{1}$ is trivial because $\mathcal{A}_{P_{1}}=\mathcal{T}_{P_{1}}$. Therefore
(8) $\mathcal{A}_{A\times P_{1}}=\mathbb{Z}A\otimes\mathcal{T}_{P_{1}}.$
If $A\times P_{1}=G$, then we are done by Eq. (8). So we may assume that $A\times P_{1}<G$. If $P\leq\operatorname{rad}(Z)$ for every $Z\in\mathcal{S}(\mathcal{A})_{G\setminus(A\times P_{1})}$, then $\mathcal{A}$ is the nontrivial $U/L=(A\times P_{1})/P_{1}$-wreath product and $U/L$ satisfies Condition $(1)$.
Due to the above discussion, we may assume that there is $Z\in\mathcal{S}(\mathcal{A})_{G\setminus(A\times P_{1})}$ with $P\nleq\operatorname{rad}(Z)$. Then Statement $(2)$ of Lemma 7.2 holds for $Z$, i.e $\langle(Z^{[p]})_{H}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $H$. Eq. (8) and $P_{1}\nleq D$ imply that $A$ is a unique nontrivial $\mathcal{A}$-subgroup of $H$ and hence $\langle(Z^{[p]})_{H}\rangle=A$. The latter equality yields that $Z^{[p]}\subseteq A\times D$. The set $Z^{[p]}$ is an $\mathcal{A}$-set by Lemma 2.4 and consequently $\langle Z^{[p]}\rangle$ is an $\mathcal{A}$-subgroup of $A\times D$. However, $A$ is a unique nontrivial $\mathcal{A}$-subgroup of $A\times D$ by Eq. (8) and $P_{1}\nleq D$. Therefore $\langle Z^{[p]}\rangle=A$. From the latter equality it follows that $Z\cap(A\times P_{1})\neq\varnothing$. As $A\times P_{1}$ is an $\mathcal{A}$-subgroup, we obtain $Z\subseteq A\times P_{1}$, a contradiction to the definition of $Z$.
Case 2. Suppose that Statement $(1)$ or Statement $(2)$ of Lemma 7.2 holds for every basic set of $\mathcal{A}$ outside $D$. Let $N$ be the $\mathcal{A}$-subgroup generated by all basic sets $X$ of $\mathcal{A}$ such that $P\nleq\operatorname{rad}(X)$. Clearly, $N\geq P_{1}$. We are done if $|N|$ is not divisible by $4$. Indeed, then $\mathcal{A}$ is the nontrivial $U/L=N/P_{1}$-wreath product and $U/L$ satisfies Condition $(3)$ from Proposition 7.3. Assume that $|N|$ is divisible by $4$ or, equivalently, $N\geq H$. Then there is $Y\in\mathcal{S}(\mathcal{A})_{G\setminus D}$ such that $P\nleq\operatorname{rad}(Y)$. Lemma 7.2 implies that $A=\langle(Y^{[p]})_{H}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $H$. As $H$ is not an $\mathcal{A}$-subgroup, $|A|=2$.
Since $N\geq H$, there is $Z\in\mathcal{S}(\mathcal{A})_{N}$ such that $P\nleq\operatorname{rad}(Z)$ and $Z_{H}\nsubseteq A$. Note that $Z\nsubseteq H$ and hence
$Z_{D}\setminus\{e\}\neq\varnothing$
because otherwise $H=\langle A,Z\rangle$ is an $\mathcal{A}$-subgroup which contradicts to the assumption of the proposition. The group $\langle(Z^{[p]})_{H}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $H$ by Lemma 7.2. If $\langle(Z^{[p]})_{H}\rangle\neq A$, then $H=\langle A,(Z^{[p]})_{H}\rangle$ is an $\mathcal{A}$-subgroup, a contradiction to the assumption of the proposition. Therefore $\langle(Z^{[p]})_{H}\rangle=A$. This implies that $Z\cap(A^{\#}\times D^{\#})\neq\varnothing$,
(9) $P\nleq\operatorname{rad}(Z\cap(A\times D)),$
and
(10) $P\leq\operatorname{rad}(Z\setminus(A\times D)).$
Let $V=\langle Z\rangle$. The $S$-ring $\mathcal{A}_{V/A}$ is an $S$-ring over a cyclic group $V/A\cong C_{2p^{i}}$, $i\in\{1,\ldots,k\}$. Due to Eq. (9) and the assumption that $H$ is not an $\mathcal{A}$-subgroup, we have $|\operatorname{rad}(Z^{\pi})|=1$, where $\pi$ is the canonical epimorphism from $G$ to $G/A$. So $|\operatorname{rad}(\mathcal{A}_{V/A})|=1$. Note that $\mathcal{A}_{V/A}$ can not be cyclotomic or tensor product of two $S$-rings because $Z$ contains elements of even as well as odd orders. Therefore $\mathcal{A}_{V/A}$ is trivial by Lemma 6.2.
Let us prove that
$A\leq\operatorname{rad}(Z).$
Since $\mathcal{A}_{V/A}$ is trivial, $Z$ contains an element $h_{0}$ of order $2$. Let $z\in Z_{D}\setminus\{e\}\neq\varnothing$ and $h\in H$ such that $hz\in Z$. For every $i\in\mathbb{Z}$ coprime to $p$, there exists $m_{i}\in\mathbb{Z}$ such that $m_{i}\equiv 1\mod~2$ and $m_{i}\equiv i\mod~|z|$. One can see that $h_{0}^{m_{i}}=h_{0}\in X^{(m_{i})}\cap X$ and hence $X^{(m_{i})}=X$ by Lemma 2.3. Therefore $(hz)^{m_{i}}=hz^{i}\in X^{(m_{i})}=X$. This implies that
(11) $hD_{|z|}^{*}\subseteq Z$
for all $z\in Z_{D}\setminus\{e\}\neq\varnothing$ and $h\in H$ such that $hz\in Z$.
Observe that $Z\cap(A\times P^{\#})\neq\varnothing$ because $\mathcal{A}_{V/A}=\mathcal{T}_{V/A}$. Eq. (11) yields that $P^{\#}\subseteq Z$ or $aP^{\#}\subseteq Z$. If the both of these two inclusions hold, then $A\leq\operatorname{rad}(Z)$ by Lemma 2.1. Further, we may assume that exactly one of the above inclusions hold. Then $P^{\#}\subseteq Z$ or $P^{\#}\subseteq aZ$. Let $Z^{\prime}\in\{Z,aZ\}$ such that $P^{\#}\subseteq Z^{\prime}$. From Eqs. (10) and (11) it follows that $Z^{\prime}\setminus P\neq\varnothing$ and $P\leq\operatorname{rad}(Z^{\prime}\setminus P)$. Therefore $|\operatorname{rad}(Z)|=1$ and $Z^{\prime}=\langle Z\rangle^{\#}$ by Lemma 2.2. We conclude that Statement $(3)$ of Lemma 7.2 holds for $Z^{\prime}$, a contradiction to the assumption of this case. Thus, $A\leq\operatorname{rad}(Z)$. Due to triviality of $\mathcal{A}_{V/A}$, we obtain
(12) $Z=V\setminus A~\text{and}~\mathcal{A}_{V}=\mathbb{Z}A\wr\mathcal{T}_{V/A}.$
The latter equality yields that $P_{1}=V$.
If $V=G$, then due to Eq. (12), $\mathcal{A}$ is the $U/L=A/A$-wreath product and $U/L$ satisfies Condition $(1)$. Further, we assume that $V<G$. If there exists $T\in\mathcal{S}(\mathcal{A})_{G\setminus V}$ such that $T\subseteq D$, then $\langle T\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $D$. So $\langle T\rangle\cap V\neq\varnothing$ and hence $V\subseteq\langle T\rangle\leq D$, a contradiction to the definition of $Z$. Therefore every basic set outside $V$ is not a subset of $D$ and hence Statement $(1)$ or $(2)$ of Lemma 7.2 holds for every such basic set. If Statement $(1)$ of Lemma 7.2 holds for every basic set $X$ outside $V$, i.e. $P\leq\operatorname{rad}(X)$, then $V=P_{1}\leq\operatorname{rad}(X)$. This implies that $\mathcal{A}$ is the $U/L=V/V$-wreath product and $U/L$ satisfies Condition $(1)$. If Statement $(2)$ of Lemma 7.2 holds for some $T\in\mathcal{S}(\mathcal{A})_{G\setminus V}$, then $\langle(T^{[p]})_{H}\rangle=A$ because $A$ is a unique nontrivial $\mathcal{A}$-subgroup of $H$. Together with Lemma 2.4, this yields that $W=\langle T^{[p]}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $A\times D$. Since $T$ lies outside $V$, we have $(T^{[p]})_{D}\neq\{e\}$ and hence $|W|$ is divisible by $p$. Therefore $W\cap Z\neq\varnothing$. Thus,
$Z\subseteq W\leq A\times D,$
a contradiction to the definition of $Z$. ∎
The case when $D$ is not an $\mathcal{A}$-subgroup is dual to the case when $H$ is not an $\mathcal{A}$-subgroup in the sense of the duality theory of $S$-rings over an abelian group [31] (see also [7, Section 2.2]). The proposition below immediately follows from Lemma 4.1 and Proposition 7.3.
Proposition 7.4.
Suppose that $D$ is not an $\mathcal{A}$-subgroup. Then $\mathcal{A}$ is trivial, or a nontrivial tensor product, or a nontrivial $U/L$-wreath product for some $\mathcal{A}$-section $U/L$ satisfying one of the following conditions:
  1. (1)
    $|U/L|\leq 2$;
  2. (2)
    $\mathcal{A}_{U/L}$ is $\otimes$-complemented in $\mathcal{A}_{G/L}$;
  3. (3)
    $4\nmid|G/L|$ and $|G/U|\neq 2$.
The next proposition provides a refinement of one of the cases from Proposition 7.3.
Proposition 7.5.
Suppose that $H$ is not an $\mathcal{A}$-subgroup and $\mathcal{A}$ is a nontrivial $U/L$-wreath product for an $\mathcal{A}$-section $U/L$ with $4\nmid|U|$ and $|L|\neq 2$. Then $\mathcal{A}$ is a nontrivial $U_{1}/L_{1}$-wreath product for an $\mathcal{A}$-section $U_{1}/L_{1}$ satisfying one of the following conditions:
  1. (1)
    $|U_{1}/L_{1}|\leq 2$;
  2. (2)
    $\mathcal{A}_{L_{1}}$ is $\otimes$-complemented in $\mathcal{A}_{U_{1}}$;
  3. (3)
    $U_{1}\leq U$ and $|\operatorname{rad}(\mathcal{A}_{U_{1}})|=1$.
Proof.
Note that $U$ is cyclic because $4\nmid|U|$. If $|\operatorname{rad}(\mathcal{A}_{U})|=1$, then $U_{1}/L_{1}=U/L$ satisfies Condition $(3)$ of Proposition 7.5 and we are done. Further, we assume that
$|\operatorname{rad}(\mathcal{A}_{U})|>1.$
Suppose that $|U|$ is odd. Then $|U|=pq$ or $|U|$ is a $p$-power. In the former case, $\mathcal{A}_{U}$ is the $U_{1}/L_{1}=L/L$-wreath product and $|L|$ is prime by Lemma 6.12. So $\mathcal{A}$ is the $U_{1}/L_{1}=L/L$-wreath product and $U_{1}/L_{1}$ satisfies Condition $(1)$ from Proposition 7.5. In the latter one, $\mathcal{A}_{U}$ is the nontrivial $U_{1}/L_{1}$-wreath product for an $\mathcal{A}_{U}$-section $U_{1}/L_{1}$ of $U$ such that $L_{1}$ is the least nontrivial $\mathcal{A}_{U}$-subgroup and $|\operatorname{rad}(\mathcal{A}_{U_{1}})|=1$ by Lemma 6.10. By the definitions of $U_{1}$ and $L_{1}$, we have $U_{1}\leq U$ and $L_{1}\leq L$. This implies that $\mathcal{A}$ is the $U_{1}/L_{1}$-wreath product and $U_{1}/L_{1}$ satisfies Condition $(3)$ as required.
Now suppose that $|U|$ is even. Then $|U|=2m$, where $m$ is odd. Let $A$ and $W$ be the Sylow $2$-subgroup and the Hall $2^{\prime}$-subgroup of $U$, respectively. If $|W|=1$, then $|L|=|U|=2$ and Condition $(1)$ from Proposition 7.5 holds for $U_{1}/L_{1}=U/L$. Further, we assume that $|W|>1$. The least $\mathcal{A}_{U}$-subgroup containing $A$ and a maximal $\mathcal{A}_{U}$-subgroup of $W$ are denoted by $A_{1}$ and $W_{1}$, respectively. If $|W_{1}|=1$, then $\mathcal{A}_{U}$ is trivial or the $A_{1}/A_{1}$-wreath product by Lemma 5.3 applied to $\mathcal{A}_{U}$, $W_{1}$, and $A_{1}$. In the former case, $U=L$ and $U_{1}/L_{1}=U/L$ satisfies Condition $(1)$. In the latter one, $A_{1}$ is the least $\mathcal{A}_{U}$-subgroup because otherwise $|W_{1}|>1$ and hence $A_{1}\leq L$. So $\mathcal{A}$ is the $U_{1}/L_{1}=A_{1}/A_{1}$-wreath product and $U_{1}/L_{1}$ satisfies Condition $(1)$.
Further, we assume that
$|W_{1}|>1.$
Observe that $M\leq L$ for some minimal $\mathcal{A}_{W_{1}}$-subgroup $M$ because $|L|\neq 2$. One can see that $\mathcal{A}_{W_{1}}$ is the $V/M$-wreath product (possibly, trivial), where $V$ is an $\mathcal{A}_{W_{1}}$-subgroup with $|\operatorname{rad}(\mathcal{A}_{V})|=1$. Indeed, this follows from Lemma 6.2 and Lemma 6.12 if $|W_{1}|=pq$ and from Lemma 6.2 and Lemma 6.10 if $|W_{1}|$ is a $p$-power. Lemma 5.4 implies that
(13) $\mathcal{A}_{W_{1}A_{1}}=\mathcal{A}_{W_{1}}\star\mathcal{A}_{A_{1}}=(\mathcal{A}_{V}\wr_{V/M}\mathcal{A}_{W_{1}/M})\star\mathcal{A}_{A_{1}}.$
Let $W_{1}A_{1}=U$. If $W_{1}\cap A_{1}$ is trivial, then $\mathcal{A}_{A_{1}}$ is primitive and hence $\mathcal{A}_{A_{1}}=\mathcal{T}_{A_{1}}$ by Lemma 6.1. Eq. (13) implies that $\mathcal{A}_{U}=\mathcal{A}_{W_{1}}\otimes\mathcal{A}_{A_{1}}$ and consequently $\mathcal{A}_{U}$ is the $U_{1}/L_{1}=(V\times A_{1})/M$-wreath product (possibly, trivial). As $M\leq L$, we conclude that $\mathcal{A}$ is the $U_{1}/L_{1}$-wreath product. Since $|\operatorname{rad}(\mathcal{A}_{V})|=1$ and
$\mathcal{A}_{U_{1}}=\mathcal{A}_{V}\otimes\mathcal{A}_{A_{1}}\cong\mathcal{A}_{V}\otimes\mathcal{T}_{A_{1}},$
we obtain $|\operatorname{rad}(\mathcal{A}_{U_{1}})|=1$. Thus, $U_{1}/L_{1}$ satisfies Condition $(3)$ and we are done.
Further, we assume that $W_{1}A_{1}<U$. This implies that
$W_{1}<W.$
In this case, $|W_{1}|$ is a prime power and hence $M$ is the unique least $\mathcal{A}_{W_{1}}$-subgroup. Lemma 5.3 yields that $\mathcal{A}_{U}=\mathcal{A}_{W_{1}}\wr\mathcal{A}_{U/W_{1}}$ or $\mathcal{A}_{U}=\mathcal{A}_{W_{1}A_{1}}\wr_{(W_{1}A_{1})/A_{1}}\mathcal{A}_{U/A_{1}}$. In the former case, $\mathcal{A}_{U}$ is the $U_{1}/L_{1}=V/M$-wreath product because $\mathcal{A}_{W_{1}}=\mathcal{A}_{V}\wr_{V/M}\mathcal{A}_{W_{1}/M}$ and $M\leq L$. Since $|\operatorname{rad}(\mathcal{A}_{V})|=1$, the section $U_{1}/L_{1}$ satisfies Condition $(3)$ and we are done.
Suppose that
(14) $\mathcal{A}_{U}=\mathcal{A}_{W_{1}A_{1}}\wr_{(W_{1}A_{1})/A_{1}}\mathcal{A}_{U/A_{1}}.$
If $\mathcal{A}_{A_{1}}\neq\mathcal{T}_{A_{1}}$, then $|A_{1}\cap W_{1}|>1$. Due to Eqs. (13) and (14), we conclude that $\mathcal{A}_{U}$ is the nontrivial $W_{1}/(W_{1}\cap A_{1})$-wreath product. Since $M$ is the unique least $\mathcal{A}_{W_{1}}$-subgroup, we have $M\leq A_{1}\cap W_{1}$ and consequently $\mathcal{A}_{U}$ is the $W_{1}/M$-wreath product. As $\mathcal{A}_{W_{1}}=\mathcal{A}_{V}\wr_{V/M}\mathcal{A}_{W_{1}/M}$, we conclude that $\mathcal{A}_{U}$ is the nontrivial $V/M$-wreath product. Observe that $\mathcal{A}$ is the $U_{1}/L_{1}=V/M$-wreath product because $M\leq L$. The section $U_{1}/L_{1}$ satisfies Condition $(3)$ because $|\operatorname{rad}(\mathcal{A}_{V})|=1$ and we are done.
Now let $\mathcal{A}_{A_{1}}=\mathcal{T}_{A_{1}}$. Then $|W_{1}\cap A_{1}|=1$. Eq. (13) implies that
(15) $\mathcal{A}_{W_{1}A_{1}}=\mathcal{A}_{W_{1}}\otimes\mathcal{T}_{A_{1}}.$
If $A_{1}\neq A$, then $|W_{1}A_{1}|=2pq$, $U=W_{1}A_{1}$, and $L\in\{W_{1},A_{1}\}$. So $U_{1}/L_{1}=U/L$ satisfies Condition $(2)$ by Eq. (15) and we are done. Further, we assume that $A_{1}=A$.
Lemma 7.6.
With the above notation, let $X\in\mathcal{S}(\mathcal{A})$ such that $X_{D}\nsubseteq W_{1}$. Then $A\leq\operatorname{rad}(X)$.
Proof.
If $X\subseteq U$, then we are done by Eq. (14). Further, we assume that $X\nsubseteq U$ and hence $L\leq R_{0}=\operatorname{rad}(X)$. Let $R=\langle X\rangle$. One can see that
$R\cap(W\setminus W_{1})\neq\varnothing$
because $X_{D}\nsubseteq W_{1}$. Together with Eq. (14), this implies that $R\geq A$. The image of $Y\subseteq G$ under the canonical epimorphism from $G$ to $G/R_{0}$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu$. Clearly, $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)|=1$ and hence $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is indecomposable.
Assume that $A\nleq R_{0}$. Then $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu\geq\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$. Due to Eq. (14), we have $R_{0}\cap D\leq W_{1}$ and
(16) $\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu\leq\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu)$
for every $Y\in\mathcal{S}(\mathcal{A})_{R}$ inside $R\cap(U\setminus(W_{1}A))$. Observe that $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is nontrivial because $\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$ is a proper nontrivial $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$-subgroup. If $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is cyclic, then $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is cyclotomic or a nontrivial tensor product by Lemma 6.2. However, this contradicts to Eq. (16). Suppose that $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is noncyclic. Then in view of Eq. (16), the Hall $2^{\prime}$-subgroup of $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is not an $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$-subgroup and hence Proposition 7.4 holds for $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$. Since $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is nontrivial and indecomposable, we conclude that
(17) $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}=\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{1}}\otimes\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{2}}$
for some $\mathcal{A}_{R}$-subgroups $R_{1}$ and $R_{2}$. As $R_{0}\geq L$ and $|L|\neq 2$, we may assume without loss of generality that $|\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{1}|$ is a power of $2$. Due to Eqs. (16) and (17), $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{1}\ngeq\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$. Therefore $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{2}$ is a cyclic group containing $\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$. Lemma 6.2 and Eq. (16) yield that $|\operatorname{rad}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{2}})|\neq 1$. Together with Eq. (17), this implies that $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is decomposable, a contradiction. ∎
From Lemma 7.6 it follows that $\mathcal{A}$ is the $(W_{1}A)/A$- or $(W_{1}H)/A$-wreath product. In the first case, $\mathcal{A}_{A}$ is $\otimes$-complemented in $\mathcal{A}_{W_{1}}$ by Eq. (15). So $U_{1}/L_{1}=(AW_{1})/A$ satisfies Condition $(2)$ and we are done.
In the second case, $HW_{1}$ is an $\mathcal{A}$-subgroup. Let $a$ be a nontrivial element of $A$, $b\in H\setminus A$, and $T$ a basic set of $\mathcal{A}$ containing $b$. The $S$-ring $\mathcal{A}_{W_{1}H}$ satisfies [13, Hypothesis $5.1$]. From [13, Lemma $5.5$] and the inclusion $\{a\}=A^{\#}\in\mathcal{S}(\mathcal{A})$ it follows that
$T=W_{0}b~\text{or}~T=W_{0}Ab$
for some $\mathcal{A}_{W_{1}}$-subgroup $W_{0}$. As $H$ is not an $\mathcal{A}$-subgroup, $|W_{0}|>1$. If $T=W_{0}Ab$, then $\mathcal{A}_{W_{1}H}$ is the $(W_{1}A)/(W_{0}A)$-wreath product by [13, Lemma $5.4$] and hence $\mathcal{A}$ is the $(W_{1}A)/A$-wreath product and we are done by the previous paragraph. If $T=W_{0}b$, then $W_{2}=\langle T,W_{1}\rangle$ is an $\mathcal{A}_{W_{1}H}$-subgroup such that $|W_{2}\cap A|=1$ and $W_{1}H=W_{2}\times A$. Lemma 3.2(2) implies that $\mathcal{A}_{W_{1}H}=\mathcal{A}_{W_{2}}\otimes\mathcal{A}_{A}$. Therefore $U_{1}/L_{1}=(W_{1}H)/A$ satisfies Condition $(2)$ and we are done. ∎
The case when $\mathcal{A}$ is a nontrivial $U/L$-wreath product satisfying Condition $(3)$ from Proposition 7.4 is dual to the case from Proposition 7.5. So the next proposition is a corollary of Proposition 7.5 and Lemma 4.1.
Proposition 7.7.
Suppose that $H$ is not an $\mathcal{A}$-subgroup and $\mathcal{A}$ is a nontrivial $U/L$-wreath product for an $\mathcal{A}$-section $U/L$ with $4\nmid|G/L|$ and $|G/U|\neq 2$. Then $\mathcal{A}$ is a nontrivial $U_{1}/L_{1}$-wreath product for an $\mathcal{A}$-section $U_{1}/L_{1}$ satisfying one of the following conditions:
  1. (1)
    $|U_{1}/L_{1}|\leq 2$;
  2. (2)
    $\mathcal{A}_{U_{1}/L_{1}}$ is $\otimes$-complemented in $\mathcal{A}_{G/L_{1}}$
  3. (3)
    $L_{1}\geq L$ and $|\operatorname{rad}(\mathcal{A}_{G/L_{1}})|=1$.
7.3. Dense $S$-rings
The lemma below easily follows from computer calculations using [14].
Lemma 7.8.
Every $S$-ring over $H\cong E_{4}$ is normal and can be presented in the form $\operatorname{Cyc}(K_{H},H)$, where $K_{H}\leq\operatorname{Aut}(H)$ is of order at most $3$.
Remark 7.9.
With the notation of Lemma 7.8, it is easy to see that $\operatorname{Cyc}(K_{H},H)=\mathbb{Z}H$ if $|K_{H}|=1$, $\operatorname{Cyc}(K_{H},H)\cong\mathbb{Z}C_{2}\wr\mathbb{Z}C_{2}$ if $|K_{H}|=2$, and $\operatorname{Cyc}(K_{H},H)=\mathcal{T}_{H}$ if $|K_{H}|=3$.
Further throughout this subsection, we assume that $H$ and $D$ are $\mathcal{A}$-subgroups. Denote two distinct elements from $H$ by $a$ and $b$. Given $X\in\mathcal{S}(\mathcal{A})$, put $\lambda_{X}=|X\cap Hx|$, where $x\in X$. Due to Lemma 2.1, $\lambda_{X}$ does not depend on $x$.
Lemma 7.10.
Let $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$ such that $X\neq X_{H}\times X_{D}$. Then $\lambda_{X}=1$.
Proof.
Clearly, $\lambda_{X}\leq|X_{H}|$ and the equality is attained if and only if $X=X_{H}\times X_{D}$. Therefore the lemma holds if $|X_{H}|\leq 2$. So in view of Lemma 3.2(1) and Remark 7.9, we may assume that $X_{H}=H^{\#}=\{a,b,ab\}$. To finish the proof, it remains to show that $\lambda_{X}\neq 2$. Assume the contrary. Then the set $Y=(X_{H}\times X_{D})\setminus X$ is an $\mathcal{A}$-set containing exactly one element from $Hx$ for every $x\in X_{D}$. Together with Lemma 3.2(1) and the equality $X_{H}=H^{\#}$, this implies that $Y$ is exactly a basic set of $\mathcal{A}$ with
(18) $\lambda_{Y}=1.$
Let $x,x^{\prime}\in X_{D}$ of the same order. Then there is a positive integer $m$ coprime to $2n$ such that $x^{n}=x^{\prime}$. Since $\lambda_{X}=2$, the sets $H_{x}=Xx^{-1}\cap H$ and $H_{x^{\prime}}=X(x^{\prime})^{-1}\cap H$ are subsets of $H^{\#}$ of size $2$. Therefore there is $h\in H_{x}\cap H_{x^{\prime}}$. One can see that $(hx)^{m}=hx^{\prime}\in X^{(m)}\cap X$. So $X^{(m)}=X$ by Lemma 2.3. This yields that
$(H_{x}x)^{m}=H_{x}x^{m}=H_{x}x^{\prime}=H_{x^{\prime}}x^{\prime}$
and hence $H_{x}=H_{x^{\prime}}$.
Let $l$ be the maximum of orders of the elements from $X_{D}$. Due to the above discussion, there is $H_{0}\subseteq H$ such that $|H_{0}|=2$ and $H_{x}=H_{0}$ for every $x\in(X_{D}\cap D_{l}^{*})$. So
(19) $X_{D}\cap D_{l}^{*}\subseteq Y_{h_{0}},$
where $h_{0}$ is a unique nontrivial element from $H\setminus H_{0}$. Let $h\in H_{0}$. From Eq. (18) it follows that $Y_{h_{0}}\cap Y_{h}=\varnothing$. Together with $Y_{h_{0}}\cup Y_{h}\subseteq X_{D}$ (Lemma 3.2(1)) and Eq. (19), this implies that
(20) $Y_{h}\subseteq X_{D}\setminus D_{l}^{*}.$
Observe that $|Y_{h_{0}}|=|Y_{h}|$ by Lemma 2.1. In view of Eqs. (19) and (20), to obtain a contradiction, it is enough to show that
(21) $|X_{D}\cap D_{l}^{*}|>|X_{D}\setminus D_{l}^{*}|.$
The set $X_{D}$ is a basic set of $\mathcal{A}_{D}$ (Lemma 3.2(1)). If $|X_{D}\setminus D_{l}^{*}|=0$, then Eq. (21) is obvious. Suppose that $|X_{D}\setminus D_{l}^{*}|>0$, i.e. $X_{D}$ is nonregular. If $n=pq$, then Lemma 6.12 implies that $\mathcal{A}_{D}$ is trivial or a nontrivial generalized wreath product. So $X_{D}=D^{\#}$ or $X_{D}$ is a union of some $L$-cosets, where $L=\operatorname{rad}(X_{D})$ is a nontrivial $\mathcal{A}_{D}$-subgroup. In the former case, $|X_{D}\cap D_{l}^{*}|=(p-1)(q-1)$ and $|X_{D}\setminus D_{l}^{*}|=p+q-2$, whereas in the latter one $|X_{D}\cap D_{l}^{*}|=r(|L|-1)$ and $|X_{D}\setminus D_{l}^{*}|=r$, where $r=|X_{D}/L|$. If $n=p^{k}$, then due to Lemma 6.11, we obtain $X_{D}=U\setminus L$ for some $\mathcal{A}$-subgroups $U\cong C_{p^{l}}>L$ and consequently $|X_{D}\cap D_{l}^{*}|=p^{l-1}(p-1)$ and $|X_{D}\setminus D_{l}^{*}|=p^{l-1}-|L|$. Since $p$ and $q$ are odd, Eq. (21) holds in all cases and we are done. ∎
Proposition 7.11.
Suppose that $|\operatorname{rad}(\mathcal{A}_{D})|=1$ and $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$. Then the following statements hold:
  1. (1)
    $\mathcal{A}$ is normal, cyclotomic, and $2$-minimal;
  2. (2)
    $|\operatorname{Aut}(\mathcal{A})^{H}|\leq 12$.
Proof.
If $\mathcal{A}_{H}=\mathbb{Z}H$, then $\mathcal{A}=\mathcal{A}_{H}\otimes\mathcal{A}_{D}$ by Lemma 3.2(2), a contradiction to the assumption of the proposition. So we may assume that $\mathcal{A}_{H}\neq\mathbb{Z}H$. Then Lemma 7.8 and Remark 7.9 imply that $\mathcal{A}_{H}=\operatorname{Cyc}(K_{H},H)$, where $K_{H}\leq\operatorname{Aut}(H)$ is of order $2$ or $3$, and a unique nonsingleton basic set of $\mathcal{A}_{H}$ is of size $|K_{H}|$. We divide the rest of the proof into two cases depending on whether $\mathcal{A}_{D}$ is cyclotomic or not.
Case 1: $\mathcal{A}_{D}$ is noncyclotomic. In view of Lemma 6.2, the $S$-ring $\mathcal{A}_{D}$ is trivial and hence $D^{\#}\in\mathcal{S}(\mathcal{A}_{D})$. Moreover, $n$ is not prime by Lemma 6.7. So further, we assume that $n=pq$ or $n=p^{k}$ for $k\geq 2$. If $n=pq$, then $H$ is a maximal $\mathcal{A}$-subgroup whose order is not divisible by $p$ and $D$ is the least $\mathcal{A}$-subgroup whose order is divisible by $p$. By Lemma 5.3 applied to $\mathcal{A}$, $H$, and $D$, we have $\mathcal{A}=\mathcal{A}_{H}\star\mathcal{A}_{D}$. Since $H\cap D$ is trivial, we obtain $\mathcal{A}=\mathcal{A}_{H}\otimes\mathcal{A}_{D}$, a contradiction to the assumption of the proposition.
Further, we assume that $n=p^{k}$, $k\geq 2$. Since $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$, there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$ such that $X\neq X_{H}\times X_{D}$. Note that $X_{H}\in\mathcal{S}(\mathcal{A}_{H})$ and $X_{D}=D^{\#}$ by Lemma 3.2(1). Since $X\neq X_{H}\times X_{D}$, we have $|X_{H}|>1$. So $|X_{H}|\in\{2,3\}$ by Lemma 7.8. Lemma 7.10 implies that $\lambda_{X}=1$. If $P\nleq\operatorname{rad}(X_{h,p^{l}})$ for some $h\in X_{H}$ and $l\in\{2,\ldots,k\}$, then the set $Y=X^{[p]}$, which is an $\mathcal{A}$-set by Lemma 2.4, lies outside $H\cup D$ and $Y_{D}\subsetneq D^{\#}$, a contradiction to Lemma 3.2(1). Therefore
(22) $P\leq\operatorname{rad}(X_{h}\setminus X_{h,p})$
for every $h\in X_{H}$. Observe that
$\bigcup\limits_{h\in X_{H}}X_{h,p}=P^{\#}$
because $X_{D}=D^{\#}$. Together with Lemma 2.5, this yields $\mathcal{B}=\operatorname{Cyc}(K,P)$, where $\mathcal{B}=\operatorname{Span}_{\mathbb{Z}}\{\underline{e},\underline{X}_{h,p}:~h\in X_{H}\}$ and $K$ is a subgroup of $\operatorname{Aut}(P)\cong C_{p-1}$ of index $|X_{H}|$, and hence
(23) $|X_{h,p}|=(p-1)/|X_{H}|.$
If $|X_{H}|=2$ and (p1)/2(p-1)/2 is odd, then $\mathcal{B}$ is an antisymmetric $S$-ring of rank $3$. Using the formulas for the intersection numbers of an antisymmetric association scheme of rank $3$ (see, e.g. [3, Exercise 2.7.57]), one can compute that
(24) $\sum\limits_{h\in X_{H}}\underline{X_{h,p}}^{2}=\frac{p-1}{2}\underline{P}^{\#}.$
If $|X_{H}|=2$ and (p1)/2(p-1)/2 is even or $|X_{H}|=3$, then from [29, Lemma 5.2] it follows that
(25) $\sum\limits_{h\in X_{H}}\underline{X_{h,p}}^{2}=(p-1)e+(\frac{p-1}{|X_{H}|}-1)\underline{P}^{\#}.$
One can see that
$\underline{X}^{2}=\sum\limits_{h\in X_{h}}\underline{X_{h}}^{2}+\sum\limits_{h,h^{\prime}\in X_{H}}hh^{\prime}\underline{X_{h}}\cdot\underline{X_{h^{\prime}}}\in\mathcal{A}.$
Since $D$ is an $\mathcal{A}$-subgroup, we conclude that
$\xi=\sum\limits_{h\in X_{h}}\underline{X_{h}}^{2}\in\mathcal{A}.$
Due to Eqs. (1), (22), and (23),
$\xi=\sum\limits_{h\in X_{h}}(\underline{X_{h}\setminus X_{h,p}}+\underline{X_{h,p}})^{2}=p\eta+2(p-1)/|X_{H}|\sum\limits_{h\in X_{H}}(\underline{X_{h}}-\underline{X_{h,p}})+\sum\limits_{h\in X_{H}}\underline{X_{h,p}}^{2}$
for some $\eta\in\mathbb{Z}D$. The above equality implies that every element from $D\setminus P$ enters $\xi$ with the coefficient equal to $2(p-1)/|X_{H}|$ modulo $p$. On the other hand, the above expression for $\xi$ and Eqs. (24) and (25) imply that every element from $P^{\#}$ enters $\xi$ with coefficient equal to $(p-1)/|X_{H}|$ or $(p-1)/|X_{H}|-1$ modulo $p$. However, each element from $D^{\#}$ must enter $\xi$ with the same coefficient because $D^{\#}\in\mathcal{S}(\mathcal{A})$, a contradiction.
Case 2: $\mathcal{A}_{D}$ is cyclotomic. In this case, all basic sets of $\mathcal{A}_{D}$ are regular. Let $K_{D}\leq\operatorname{Aut}(D)$ such that $\mathcal{A}_{D}=\operatorname{Cyc}(K_{D},D)$. We start the proof in this case with the following claim.
Claim 1.
Let $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$. Then $X^{(r)}\in\mathcal{S}(\mathcal{A})$ for every $r\in\mathbb{Z}$.
Proof.
It is enough to prove the claim for every prime $r$. Indeed, if $r$ is composite, then successively applying the claim for each prime divisor of $r$, we obtain the required. If $r$ is coprime to $2n$, then the claim follows from Lemma 2.3. Lemma 3.2(1) implies that $X_{D}\in\mathcal{S}(\mathcal{A}_{D})$. If $r=2$, then $X^{(r)}=X^{(2)}=X_{D}^{(2)}\in\mathcal{S}(\mathcal{A}_{D})$ by Lemma 2.3. Further, we assume that $r$ is an odd prime divisor of $n$.
Put $Y=X^{(r)}$. Let $R$ be the (unique) subgroup of $D$ of order $R$. Since $\mathcal{A}_{D}$ is cyclotomic, $R$ is an $\mathcal{A}$-subgroup. Observe that $R\nleq\operatorname{rad}(X_{D})$. Indeed, if $n=p^{k}$, then this follows from Lemma 6.8(2), whereas if $n=pq$, then this immediately follows from the the fact that $\mathcal{A}$ is cyclotomic with trivial radical. Therefore $R\nleq\operatorname{rad}(X)$ and hence $Y=X^{[r]}$. Lemma 2.4 yields that $Y$ is an $\mathcal{A}$-set.
If $Y_{D}$ is trivial, then $Y=X_{H}$ and we are done by Lemma 3.2(1). Further, we assume that $Y_{D}$ is nontrivial. Let us show that
(26) $Y_{D}\in\mathcal{S}(\mathcal{A}_{D})$
Due to Lemma 3.2(1), the set $Y_{D}$ is an $\mathcal{A}_{D}$-set. One can see that $Y_{D}=X_{D}^{(r)}$. If $n=p^{k}$, then $|Y_{D}|=|X_{D}|$ because all the elements of $X_{D}$ lie in pairwise distinct $R$-cosets (Lemma 6.8(1)). From Lemma 6.8(2) it follows that all nontrivial basic set of $\mathcal{A}_{D}$ has the same size and hence Eq. (26) holds. If $n=pq$, then $R$ and the Hall $r^{\prime}$-subgroup $U$ of $D$ are $\mathcal{A}$-subgroups because $\mathcal{A}_{D}$ is cyclotomic. So $H$, $U$, and $R$ are $\mathcal{A}$-subgroups such that $G=H\times U\times R$. As $R\nleq\operatorname{rad}(X)$, one can see that $Y=X_{H\times U}$ and consequently $Y_{D}=(X_{H\times U})_{D}=X_{U}$. Therefore $Y_{D}\in\mathcal{S}(\mathcal{A}_{U})$ by Lemma 3.2(1) applied to $G_{1}=H\times R$ and $G_{2}=U$. Thus, Eq. (26) holds.
Assume that $Y\notin\mathcal{S}(\mathcal{A})$. Then there exists $Z\in\mathcal{S}(\mathcal{A})$ such that $Z\subsetneq Y$. Suppose that $\lambda_{X}=1$. Then $\lambda_{Y}=1$ and hence $|Y|=|Y_{D}|$. Together with $Z\subsetneq Y$, this implies that $Z_{D}\subsetneq Y_{D}$. Due to Eq. (26), we conclude that $Z_{D}\notin\mathcal{S}(\mathcal{A}_{D})$, a contradiction to Lemma 3.2(1).
Now suppose that $\lambda_{X}\neq 1$. Then $X=X_{H}\times X_{D}$ by Lemma 7.10 and $\lambda_{X}=|X_{H}|\geq 2$. So $Y=Y_{H}\times Y_{D}=X_{H}\times Y_{D}$. As $Z_{H},Y_{H}\in\mathcal{S}(\mathcal{A}_{H})$ and $Z_{H}\subseteq Y_{H}$, we have $Z_{H}=Y_{H}$. Besides, $Z_{D}=Y_{D}$ by Eq. (26). Since $Z\subsetneq Y$, we have $\lambda_{Z}=1$ by Lemma 7.10.
If $n=p^{k}$, then given $x\in X_{D}$, all the elements from $Y_{D}x$ lie in the same $\langle Y_{D}\rangle$-coset. So by Lemma 6.8(1), they lie in pairwise distinct basic sets of $\mathcal{A}_{D}$. Together with $|X_{D}|=|Y_{D}|$, this implies that there are at least $|X_{D}|$ pairwise distinct $T_{D}\in\mathcal{S}(\mathcal{A}_{D})$ with $c_{X_{D}Y_{D}}^{T_{D}}\geq 1$. Therefore there are at least $|X_{D}|$ pairwise distinct $T\in\mathcal{S}(\mathcal{A})$ with $c_{X_{D}Z}^{T}\geq 1$. Every such $T$ is rationally conjugate to $X$ (Lemma 2.3) and hence $|T|=|X|=|X_{H}||X_{D}|$. Thus, at least $|X_{H}||X_{D}|^{2}$ elements enter the element $\underline{X_{D}}\cdot\underline{Z}$. On the other hand, $|Z|=|Z_{D}|=|Y_{D}|=|X_{D}|$ because $\lambda_{Z}=1$ and consequently exactly $|X_{D}|^{2}$ elements enter the element $\underline{X}_{D}\cdot\underline{Z}$, a contradiction to $|X_{H}|=|K_{H}|\geq 2$.
If $n=pq$, then $X^{\pi}\cap Z^{\pi}\neq\varnothing$, where $\pi$ is the canonical epimorphism from $G$ to $G/R$. Since $X^{\pi},Z^{\pi}\in\mathcal{S}(\mathcal{A}_{G/R})$, we conclude that $X^{\pi}=Z^{\pi}$. On the other hand, $|X^{\pi}\cap H^{\pi}x|=|X_{H}|\geq 2$ for every $x\in X^{\pi}$, whereas $|Z^{\pi}\cap H^{\pi}z|=1$ for every $z\in Z^{\pi}$ by Lemma 2.1, a contradiction. ∎
In view of Lemma 3.2(1) and the assumptions on $\mathcal{A}_{H}$ and $\mathcal{A}_{D}$, one can choose a basic set $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$ such that
$|X_{H}|=|K_{H}|>1$
and $X_{D}$ consists of generators of $D$. Clearly, $X_{H}\in\operatorname{Orb}(K_{H},H)$ and $X_{D}\in\operatorname{Orb}(K_{D},D)$. Due to Lemma 7.10, $\lambda_{X}\in\{1,|X_{H}|\}$. Let $\mathcal{S}_{0}=\{X^{(r)}:~r\in\mathbb{Z}\}$. Claim $1$ implies that $\mathcal{S}_{0}\subseteq\mathcal{S}(\mathcal{A})$.
Claim 2.
If $Y\in\mathcal{S}(\mathcal{A})\setminus(\mathcal{S}(\mathcal{A}_{H})\cup\mathcal{S}(\mathcal{A}_{D})\cup\mathcal{S}_{0})$, then $|Y_{H}|=1$ and, in particular, $Y=Y_{H}\times Y_{D}$.
Proof.
If $Y_{H}=X_{H}$, then $Y\subseteq H$ or $Y\in\mathcal{S}_{0}$, a contradiction to the choice of $Y$. So $Y_{H}\neq X_{H}$. Since $X_{H},Y_{H}\in\mathcal{S}(\mathcal{A}_{H})$ (Lemma 3.2(1)) and $|X_{H}|>1$, we obtain $|Y_{H}|=1$ as required. ∎
If $\lambda_{X}=|X_{H}|$, then $X=X_{H}\times X_{D}$. Let $Y\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$. Then $Y=Y_{H}\times Y_{D}$. Indeed, this is clear if $Y\in\mathcal{S}_{0}$ and follows from Claim $2$ otherwise. Therefore $\mathcal{A}=\mathcal{A}_{H}\otimes\mathcal{A}_{D}$, a contradiction to the assumption of the proposition. Thus,
$\lambda_{X}=1.$
Lemma 2.5 implies that $X_{h}$ is an orbit of some $K^{0}_{D}\leq\operatorname{Aut}(D)$ for every $h\in H$. Since both $K_{D}$ and $K^{0}_{D}$ are semiregular on $X_{D}$, we conclude that $K_{0}\leq K_{D}$ and $|K_{D}:K^{0}_{D}|=|X_{H}|$. As $|K_{H}|=|X_{H}|=|K_{D}/K^{0}_{D}|=|K_{H}|\leq 3$, there is an isomorphism $\psi$ from $K_{H}$ to $K_{D}/K^{0}_{D}$ such that
$(X_{h})^{\sigma^{\psi}}=X_{h^{\sigma}}.$
Put
$K=\{(\sigma,\tau)\in K_{H}\times K_{D}:~(\sigma)^{\psi}=K^{0}_{D}\tau\}.$
By the definition, $K$ is a subdirect product of $K_{H}$ and $K_{D}$.
Put $\mathcal{B}=\operatorname{Cyc}(K,G)$. To prove that $\mathcal{A}$ is cyclotomic, it is enough to show that $\mathcal{A}=\mathcal{B}$. One can see that $\mathcal{A}_{H}=\mathcal{B}_{H}$ ($\mathcal{A}_{D}=\mathcal{B}_{D}$, respectively) because $K^{H}=K_{H}$ ($K^{D}=K_{D}$, respectively). As $X_{H}$ and $X_{D}$ are orbits of $K_{H}$ and $K_{D}$, respectively, we conclude that $X\in\operatorname{Orb}(K,G)$ and hence $X\in\mathcal{S}(\mathcal{B})$. Due to Claim $1$ applied to $\mathcal{A}$ and to $\mathcal{B}$, we have
$\mathcal{S}_{0}\subseteq\mathcal{S}(\mathcal{A})\cap\mathcal{S}(\mathcal{B}).$
We are done if $\mathcal{S}(\mathcal{A})=\mathcal{S}(\mathcal{A}_{H})\cup\mathcal{S}(\mathcal{A}_{D})\cup\mathcal{S}_{0}$. Otherwise, $Y=Y_{H}\times Y_{D}$ for every $Y\in\mathcal{S}(\mathcal{A})\setminus(\mathcal{S}(\mathcal{A}_{H})\cup\mathcal{S}(\mathcal{A}_{D})\cup\mathcal{S}_{0})$ by Claim $2$ and hence $Y\in\operatorname{Orb}(K,G)=\mathcal{S}(\mathcal{B})$ by the definition of $K$. Thus, $\mathcal{A}=\mathcal{B}$ and we are done.
Now let us prove that $\mathcal{A}$ is $2$-minimal. Let $x\in X$. It is enough to show that the pointwise stabilizer $\operatorname{Aut}(\mathcal{A})_{e,x}$ is trivial. Indeed, then $X$ is a faithful regular orbit of $\operatorname{Aut}(\mathcal{A})_{e}$ and $\mathcal{A}$ is $2$-minimal by [21, Lemma 8.2]. It suffices to verify that every $f\in\operatorname{Aut}(\mathcal{A})_{e,x}$ is trivial.
Claim 3.
If $y^{f}=y$ for some $y\in G\setminus H$, then $(Hy)^{f}=Hy$ and $f^{Hy}=\operatorname{id}_{Hy}$.
Proof.
As $H$ is an $\mathcal{A}$-subgroup, each $H$-coset is a block of $\operatorname{Aut}(\mathcal{A})$. One can see that $y=y^{f}\in Hy\cap(Hy)^{f}$. So $(Hy)^{f}=Hy$. Since $\mathcal{A}=\operatorname{Cyc}(K,G)$, we have $\lambda_{Y}=1$ for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$. Therefore all the elements of $Hy$ lie in pairwise distinct basic sets. Together with $(Hy)^{f}=Hy$, this implies that $f^{Hy}=\operatorname{id}_{Hy}$. ∎
The element $x$ can be uniquely presented in the form $x=hx_{0}$, where $h\in H^{\#}$ and $x_{0}\in D$ is a generator of $D$. From Claim $3$ it follows that $x_{0}^{f}=x_{0}$. Observe that $D^{f}=D$ because $D$ is an $\mathcal{A}$-subgroup and hence $f^{D}\in\operatorname{Aut}(\mathcal{A}_{D})$. Lemma 6.5 yields that $\operatorname{Aut}(\mathcal{A}_{D})=G_{r}\rtimes K_{D}\leq\operatorname{Hol}(G)$. So $f^{D}\in K_{D}\leq\operatorname{Aut}(D)$. Together with $x_{0}^{f}=x_{0}$, this implies that $f^{D}=\operatorname{id}_{D}$. Therefore $f$ fixes an element from every $H$-coset. Thus, $f$ is trivial by Claim $3$.
Since $\mathcal{A}=\operatorname{Cyc}(K,G)$ and $\mathcal{A}$ is $2$-minimal, we conclude that $\operatorname{Aut}(\mathcal{A})=G_{r}\rtimes K$. This implies that $\mathcal{A}$ is normal and $|\operatorname{Aut}(\mathcal{A})^{H}|=|H||K^{H}|\leq 12$ which finishes the proof of the proposition. ∎
Proposition 7.12.
Suppose that $\mathcal{A}_{D}$ is the nontrivial $U/L$-wreath product, where $L$ is a minimal nontrivial $\mathcal{A}_{D}$-subgroup and $U$ is an $\mathcal{A}_{D}$-subgroup with $|\operatorname{rad}(\mathcal{A}_{U})|=1$. Then one of the following statements holds:
  1. (1)
    $\mathcal{A}$ is the nontrivial $(H\times U)/L$-wreath product;
  2. (2)
    $n=3^{k}$, $\mathcal{A}$ is the $(H\times W)/L$-wreath product (possibly, trivial) for an $\mathcal{A}_{D}$-subgroup $W>U$ such that $|\operatorname{rad}(\mathcal{A}_{W})|=3$ and $\mathcal{A}_{H\times W}$ is cyclotomic.
Proof.
At first, suppose that $n=pq$. Without loss of generality, we may assume that $|U|=|L|=p$. Then $H$ is a maximal $\mathcal{A}$-subgroup whose order is not divisible by $p$. So $\mathcal{A}$ is the nontrivial $(H\times U)/L$-wreath product by Lemma 5.3 applied to $H_{1}=H$ and $P_{1}=L$ and Statement $(1)$ of the proposition holds.
Now suppose that $n=p^{k}$.
Claim 4.
If $L\nleq\operatorname{rad}(X)$ for some $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times U)}$, then $n=3^{k}$, $\mathcal{A}_{H}=\mathcal{T}_{H}$, $L=P$, and $X$ is one of the following forms
(27) $h_{1}\{x\}\cup h_{2}\{xx_{0}\}\cup h_{3}\{xx_{0}^{2}\},~h_{1}\{x,x^{-1}\}\cup h_{2}\{xx_{0},x^{-1}x_{0}^{2}\}\cup h_{3}\{xx_{0}^{2},x^{-1}x_{0}\},$
where $h_{1},h_{2},h_{3}$ are pairwise distinct nontrivial elements of $H$, $x_{0}$ is a generator of $P$, and $x\in D$. In particular, $|X|=|X_{D}|\leq 6$.
Proof.
Note that $X_{D}\in\mathcal{S}(\mathcal{A}_{D})$ by Lemma 3.2(1) and hence $L\leq\operatorname{rad}(X_{D})$. If $X=X_{H}\times X_{D}$, then $L\leq\operatorname{rad}(X)$, a contradiction to the assumption of the claim. Otherwise, $|X_{H}|\in\{2,3\}$ by Lemma 7.8 and $\lambda_{X}=1$ by Lemma 7.10. The number $\mu_{X}=|X\cap Lx|$ does not depend on $x\in X$ by Lemma 2.1. Since $\lambda_{X}=1$, we conclude that $\mu_{X}=|L|$ or $\mu_{X}=|L|/|X_{H}|$. In the former case, $L\leq\operatorname{rad}(X)$, a contradiction to the assumption of the claim. In the latter one, since $|X_{H}|\in\{2,3\}$ and $|L|$ is a $p$-power for an odd prime $p$, we obtain $|X_{H}|=p=3$. This implies that $n=3^{k}$, $\mathcal{A}_{H}=\mathcal{T}_{H}$, and $X_{H}=H^{\#}$.
Let $X_{D}$ be regular. Then all $X_{h}$’s, $h\in H^{\#}$, are orbits of some $K\leq\operatorname{Aut}(\langle X_{D}\rangle)$ by Lemma 2.5. Since $L\nleq\operatorname{rad}(X)$, we obtain $P\nleq\operatorname{rad}(X_{h})$ and hence $|\operatorname{rad}(X_{h})|=1$ for every $h\in H^{\#}$. Lemma 6.8(1) implies that $|K|=|X_{h}|\leq 2$ for every $h\in H^{\#}$. Therefore each $X_{h}$ is singleton or each $X_{h}$ is of the form $\{x,x^{-1}\}$, $x\in D$. Due to $\lambda_{X}=1$, all $X_{h}$’s are pairwise disjoint and consequently
$\sum\limits_{h\in H^{\#}}|X_{h}|=|X_{D}|\in\{3,6\}.$
As $L\leq\operatorname{rad}(X_{D})$, we have $L=\operatorname{rad}(X_{D})=P$ and hence $X_{D}=Px$ or $X_{D}=Px\cup Px^{-1}$. Together with the previous discussion, this yields that $X$ is one of the required forms.
Now let $X_{D}$ be nonregular. Then $X_{D}=V\setminus N$ for some $\mathcal{A}_{D}$-subgroups $V>N\geq U$ by Lemma 6.11. Let $m$ be the minimum of orders of the elements from $X_{D}$. If $P\nleq\operatorname{rad}(X_{h,l})$ for some $h\in H^{\#}$ and $l>m$, then
$\varnothing\neq X^{[p]}_{D}\subsetneq X_{D}.$
However, $X^{[p]}$ is an $\mathcal{A}$-set by Lemma 2.4 and hence $X^{[p]}_{D}$ is a union of some basic sets of $\mathcal{A}_{D}$ by Lemma 3.2(1), a contradiction to the above inclusion. Therefore
(28) $P\leq\operatorname{rad}(X_{h,l})$
for all $h\in H^{\#}$ and $l>m$. Due to Lemma 2.5, all $X_{h,m}$, $h\in H^{\#}$, are orbits of some $K\leq\operatorname{Aut}(D_{m})$. As $L\nleq\operatorname{rad}(X)$, Eq. (28) implies that $P\nleq\operatorname{rad}(X_{h,m})$ and hence $|\operatorname{rad}(X_{h,m})|=1$ for every $h\in H^{\#}$. From Lemma 6.8(1) it follows that $|X_{h,m}|\leq 2$ for every $h\in H^{\#}$. Together with $X_{D}=V\setminus N$, this yields that
$3^{m}-3^{m-1}=|D_{m}^{*}|=|\bigcup\limits_{h\in H^{\#}}X_{h,m}|\leq 6.$
Therefore $m\leq 2$. If $m=1$, then $U$ is trivial, a contradiction to the assumption of the proposition. If $m=2$, then $U=L=P$ is an $\mathcal{A}$-subgroup. By Lemma 2.1 and Eq. (28), we obtain $P=L\leq\operatorname{rad}(X)$, a contradiction to the assumption of the claim. ∎
If $L\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times U)}$, then $\mathcal{A}$ is the nontrivial $(H\times U)/L$-wreath product and Statement $(1)$ of the proposition holds.
Suppose that $L\nleq\operatorname{rad}(X)$ for some $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times U)}$. By Claim $4$, we have $n=3^{k}$, $\mathcal{A}_{H}=\mathcal{T}_{H}$, and $L=P$. Let $W$ be the $\mathcal{A}_{D}$-subgroup generated by $X_{D}$, where $X$ runs over all basic sets of $\mathcal{A}$ outside $H\times U$ such that $P\nleq\operatorname{rad}(X)$. Note that $W$ is indeed an $\mathcal{A}_{D}$-subgroup by Lemma 3.2(1). By the definition, $W>U$. Clearly, $\mathcal{A}$ is the $(H\times W)/P$-wreath product. One can see that $\operatorname{rad}(\mathcal{A}_{W})=P$ by Claim $4$.
In view of the above paragraph, to prove Statement $(2)$ of the proposition, it is enough to show that $\mathcal{A}_{H\times W}$ is cyclotomic. Let $X\in\mathcal{S}(\mathcal{A}_{H\times W})_{G\setminus(H\times U)}$ such that $P\nleq\operatorname{rad}(X)$ and $X_{D}$ contains a generator of $W$. By Claim $4$, the set $X$ is one of the forms from Eq. (27), where $x\in W$ is a generator of $W$. Let $\sigma\in\operatorname{Aut}(H)$ such that $\sigma=(h_{1}h_{2}h_{3})$ and $\tau_{1},\tau_{2}\in\operatorname{Aut}(W)$ such that $x^{\tau_{1}}=xx_{0}$ and $x^{\tau_{2}}=x^{-1}$. Put
$K_{1}=\langle\varphi\rangle\leq\operatorname{Aut}(H\times W)~\text{and}~K_{2}=K_{1}\times\langle\psi\rangle\leq\operatorname{Aut}(H\times W),$
where $\varphi,\psi\in\operatorname{Aut}(H\times W)$ are such that $\varphi^{H}=\sigma$, $\varphi^{D}=\tau_{1}$, $\psi^{H}=\operatorname{id}_{H}$, $\psi^{D}=\tau_{2}$. Clearly, $K_{1}\cong C_{3}$ and $K_{2}\cong C_{6}$. Due to the definitions of $K_{1}$ and $K_{2}$, we conclude that $X\in\operatorname{Orb}(K,H\times W)$ and $|X|=|K|$, where $K=K_{1}$ if $X$ is of the first form from Eq. (27) and $K=K_{2}$ otherwise. It is easy to see that $H^{\#},X_{D}\in\operatorname{Orb}(K,H\times W)$. Put $X^{0}=X$ and $X^{i}=(X^{i-1})^{[p]}$ for $i\in\{1,\ldots,l-1\}$, where $l$ is such that $|W|=3^{l}$. Eq. (27) implies that $X^{i}=H^{\#}x^{p^{i}}$ or $X^{i}=H^{\#}x^{p^{i}}\cup H^{\#}x^{-p^{i}}$ for $i\in\{1,\ldots,l-1\}$. Each $X^{i}$ is an $\mathcal{A}$-set by Lemma 2.4. Moreover, using Lemma 2.1 and Lemma 3.2(1), it is easy to verify that each $X^{i}$ and $X^{i}_{D}$ are exactly basic sets of $\mathcal{A}_{H\times W}$ and $\mathcal{A}_{W}$, respectively. Due to the definition of $K$, each $X^{i}$ and $X^{i}_{D}$ are orbits of $K$. Every $Y\in\mathcal{S}(\mathcal{A}_{H\times W})_{G\setminus H}$ is rationally conjugate to $X^{i}$ or $X^{i}_{D}$ for some $i\in\{0,\ldots,l-1\}$ by Lemma 2.3 and hence $Y\in\operatorname{Orb}(K,H\times W)$. Thus, every basic set of $\mathcal{A}_{H\times W}$ is an orbit of $K$, i.e. $\mathcal{A}_{H\times W}=\operatorname{Cyc}(K,G)$ as required. ∎
Lemma 7.13.
Suppose that $\mathcal{A}$ is schurian, $\mathcal{A}_{H}=\mathcal{T}_{H}$, and $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$. Then $|\operatorname{Aut}(\mathcal{A})^{H}|=12$.
Proof.
One can see that $\operatorname{Aut}(\mathcal{A})^{H}\approx_{2}\operatorname{Aut}(\mathcal{A}_{H})\cong\operatorname{Sym}(4)$ because $\mathcal{A}$ is schurian. Due to Remark 2.7, we conclude that $|\operatorname{Aut}(\mathcal{A})^{H}|\in\{12,24\}$.
Since $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$, there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus(U\cup D)}$ such that $X\neq X_{H}\times X_{D}$. Lemma 7.10 implies that $\lambda_{X}=1$. From Lemma 3.2(1) it follows that $X_{D}\in\mathcal{S}(\mathcal{A}_{D})$ and $X_{H}=H^{\#}$. Let $L=\operatorname{rad}(X)$, $U=\langle X\rangle$, and $W=\langle X_{D}\rangle$. The assumptions of the lemma yield that $L\leq D$. The image of $T\subseteq G$ under the canonical epimorphism from $G$ to $G/L$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muT\mkern-1.5mu}\mkern 1.5mu$. The argument from the previous paragraph applied to $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$ implies that $|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|\in\{12,24\}$. To prove the lemma, it is enough to show that
$|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=12.$
Indeed, then
$|\operatorname{Aut}(\mathcal{A})^{H}|=|\operatorname{Aut}(\mathcal{A})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|\leq|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=12$
as required, where the first equality holds because $L\leq D$.
By the definition of $L$, we have
(29) $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)|=1.$
As $\lambda_{X}=1$ and $L\leq D$, we obtain $\lambda_{\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu}=|\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\cap\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu|=1$ for every $\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu\in\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu$. So $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}\neq\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}\otimes\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muW\mkern-1.5mu}\mkern 1.5mu}$. The set $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D}$ is a basic set of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muD\mkern-1.5mu}\mkern 1.5mu}$. If $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D})|=1$, then $|\operatorname{rad}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muW\mkern-1.5mu}\mkern 1.5mu})|=1$ and Proposition 7.11(2) applied to $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$ yields that $|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=12$ as desired.
Suppose that $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D})|>1$. In view of Lemma 6.12 if $|W|=pq$ and Lemma 6.10 if $|W|$ is a $p$-power, we conclude that Proposition 7.12 holds for $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$. If Statement $(1)$ of Proposition 7.12 holds for $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$, then $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)|>1$, a contradiction to Eq. (29). Therefore Statement $(2)$ of Proposition 7.12 holds for $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$. Due to Claim $4$ and Eq. (29), the basic sets of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$ inside $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{H}\times\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D}$ are of the form
$\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{1}\cup\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{2}\cup\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{3},~\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{2}\cup\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{3}\cup\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{1},~\mkern 1.5mu\overline{\mkern-1.5muZ\mkern-1.5mu}\mkern 1.5mu=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{3}\cup\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{1}\cup\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{2},$
where the $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{i}$’s are pairwise distinct elements of $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{H}$ and the $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{i}$’s are pairwise disjoint subsets of $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D}$ of size at most $2$. If $|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=24$, then there is $f\in\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})$ such that $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}^{f}=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}$, $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}^{f}=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}$, and $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}^{f}=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}$. Eq. (2) implies that
$\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{1}^{f}=(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}\cap\mkern 1.5mu\overline{\mkern-1.5muD\mkern-1.5mu}\mkern 1.5mu)^{f}=\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}^{f}\cap\mkern 1.5mu\overline{\mkern-1.5muD\mkern-1.5mu}\mkern 1.5mu=\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{1}$
and
$\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{1}^{f}=(\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}\cap\mkern 1.5mu\overline{\mkern-1.5muD\mkern-1.5mu}\mkern 1.5mu)^{f}=\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}^{f}\cap\mkern 1.5mu\overline{\mkern-1.5muD\mkern-1.5mu}\mkern 1.5mu=\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}\cap\mkern 1.5mu\overline{\mkern-1.5muD\mkern-1.5mu}\mkern 1.5mu=\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{3},$
a contradiction. Thus, $|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=12$ and we are done. ∎
7.4. Proof of Theorem 7.1
If $H$ ($D$, respectively) is not an $\mathcal{A}$-subgroup, then Proposition 7.3 and Proposition 7.5 (Proposition 7.4 and Proposition 7.7, respectively) imply that Theorem 7.1 holds for $\mathcal{A}$ or $\mathcal{A}$ is a nontrivial $S$-wreath product for an $\mathcal{A}$-section $S=U/L$ such that $4\nmid|W|$ and $|\operatorname{rad}(\mathcal{A}_{W})|=1$, where $W=U$ if $H$ is not an $\mathcal{A}$-subgroup and $W=G/L$ if $D$ is not an $\mathcal{A}$-subgroup. Note that $|W_{H}|\leq 2$, where $W_{H}$ is the Sylow $2$-subgroup of $W$, because $|W|$ is not divisible by $4$ and consequently $\mathcal{A}_{W_{H}}=\mathbb{Z}W_{H}$. So Lemma 6.2 and Lemma 3.2(2) yield that
(30) $\mathcal{A}_{W}=\mathcal{A}_{W_{H}}\otimes\mathcal{A}_{W_{D}},$
where $W_{D}$ is the Hall $2^{\prime}$-subgroup of $W$ (here $W_{H}$ and $W_{D}$ can be trivial). Since $|\operatorname{rad}(\mathcal{A}_{W})|=1$, we conclude that $|\operatorname{rad}(\mathcal{A}_{W_{D}})|=1$. If $\mathcal{A}_{W_{D}}=\mathcal{T}_{W_{D}}$, then due to Eq. (30), we obtain $L\in\{W_{H},W_{D},W\}$ whenever $W=U$ and $S\in\{W_{H},W_{D},\{L\}\}$ whenever $W=G/L$. This yields that Statement $(1)$ or $(2)$ from Theorem 7.1 holds. Suppose that $\mathcal{A}_{W_{D}}\neq\mathcal{T}_{W_{D}}$. Then $\mathcal{A}_{W_{D}}=\operatorname{Cyc}(K_{D},W_{D})$ for some $K_{D}\leq\operatorname{Aut}(W_{D})$ by Lemma 6.2. As $|W_{H}|\leq 2$, we have $\mathcal{A}_{W_{H}}=\operatorname{Cyc}(K_{H},W_{H})$, where $K_{H}$ is trivial. Therefore $\mathcal{A}_{W}=\operatorname{Cyc}(K_{H}\times K_{D},W)$ by Eqs. (4) and (30) and Statement $(3)$ from Theorem 7.1 holds.
Suppose that $H$ and $D$ are $\mathcal{A}$-subgroups. If $|\operatorname{rad}(\mathcal{A}_{D})|=1$, then $\mathcal{A}=\mathcal{A}_{H}\otimes\mathcal{A}_{D}$ or $\mathcal{A}$ is cyclotomic by Proposition 7.11(1) and we are done. If $|\operatorname{rad}(\mathcal{A}_{D})|>1$, then $\mathcal{A}_{D}$ is a nontrivial $U/L$-wreath product for some $\mathcal{A}_{D}$-section $U/L$ such that $L$ is a minimal nontrivial $\mathcal{A}_{D}$-subgroup and $|\operatorname{rad}(\mathcal{A}_{U})|=1$. Indeed, this follows from Lemma 6.12 if $n=pq$ and from Lemma 6.10 if $n=p^{k}$. So Proposition 7.12 holds for $\mathcal{A}$. If Statement $(2)$ of Proposition 7.12 holds for $\mathcal{A}$, then $n=3^{k}$, $\mathcal{A}$ is the $S=(H\times W)/L$-wreath product (possibly, trivial) for an $\mathcal{A}_{D}$-subgroup $W>U$ such that $|\operatorname{rad}(\mathcal{A}_{W})|=3$, and $\mathcal{A}_{H\times W}$ is cyclotomic. In this case, $H\times W=G$ and $\mathcal{A}$ is cyclotomic or $H\times W<G$, $\mathcal{A}$ is the nontrivial $S$-wreath product, and Statement $(4)$ from Theorem 7.1 holds as required.
Suppose that Statement $(1)$ of Proposition 7.12 holds for $\mathcal{A}$, i.e. $\mathcal{A}$ is the nontrivial $S=V/L$-wreath product, where $V=H\times U$. If $\mathcal{A}_{V}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{U}$, then $\mathcal{A}_{V}$ is cyclotomic by Proposition 7.11(1). So Statement $(4)$ from Theorem 7.1 holds for $\mathcal{A}$ and $S$. Further, we assume that
(31) $\mathcal{A}_{V}=\mathcal{A}_{H}\otimes\mathcal{A}_{U}.$
One can see that $\mathcal{A}_{H}=\operatorname{Cyc}(K_{H},H)$ for some $K_{H}\leq\operatorname{Aut}(H)$ by Lemma 7.8. If $\mathcal{A}_{U}=\operatorname{Cyc}(K_{U},U)$ for some $K_{U}\leq\operatorname{Aut}(U)$, then $\mathcal{A}_{V}=\operatorname{Cyc}(K_{H}\times K_{U},V)$ by Eqs. (4) and (31). Therefore Statement $(4)$ from Theorem 7.1 holds for $\mathcal{A}$ and $S$. Otherwise, $\mathcal{A}_{U}=\mathcal{T}_{U}$ by Lemma 6.2. This yields that $L=U$. Together with Eq. (31), this implies that $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{V}$. Thus, Statement $(2)$ from Theorem 7.1 holds for $S$ and we are done.
7.5. Auxiliary statements
Lemma 7.14.
Let $n=p^{k}$, $p\geq 5$, and $S$ an $\mathcal{A}$-section such that $p^{2}~|~|S|$, $\mathcal{A}_{S}$ is indecomposable, and there is an $\mathcal{A}_{S}$-subgroup of order or index $p$. Then $\mathcal{A}_{S}$ is normal.
Proof.
Clearly, $\mathcal{A}_{S}$ is nontrivial. Firstly, suppose that $S$ is cyclic. If $|S|$ is a power of $p$, then $\mathcal{A}_{S}$ is normal by Lemma 6.2. If $|S|=2p^{l}$ for some $l\geq 2$ and $\mathcal{A}_{S}$ is not normal, then $\mathcal{A}_{S}=\mathbb{Z}S_{1}\otimes\mathcal{T}_{S_{2}}$, where $S_{1},S_{2}\leq S$ with $|S_{1}|=2$ and $|S_{2}|=p^{l}$, by Lemma 6.13. However, the latter equality contradicts to the condition of the lemma that there is an $\mathcal{A}_{S}$-subgroup of order or index $p$.
Now suppose that $S$ is noncyclic, i.e. 4||S|4~|~|S|. Let
$\mathcal{A}_{S}=\mathcal{A}_{S_{1}}\otimes\mathcal{A}_{S_{2}}$
for some proper nontrivial $\mathcal{A}_{S}$-subgroups $S_{1}$ and $S_{2}$ such that $S=S_{1}\times S_{2}$. Without loss of generality, we may assume that $|S_{1}|=4$ and $|S_{2}|=p^{l}$ or $|S_{1}|=2$ and $|S_{2}|=2p^{l}$ for some $l\geq 2$. In particular, $S_{2}$ is cyclic. Note that $|\operatorname{rad}(\mathcal{A}_{S_{2}})|=1$ because otherwise $\mathcal{A}_{S_{2}}$ is decomposable by Lemma 6.2 and hence $\mathcal{A}_{S}$ so is. If $|S_{1}|=4$ ($|S_{1}|=2$, respectively), then the last condition of Lemma 7.14 and Lemma 6.2 (Lemma 6.13, respectively) imply that $\mathcal{A}_{S_{2}}$ is normal. One can see that $\mathcal{A}_{S_{1}}$ is normal. Indeed, this is clear if $|S_{2}|=2$ and follows from Lemma 7.8 if $|S_{2}|=4$. Therefore $\mathcal{A}_{S}$ is normal by Eq. (3).
In view of the above paragraph, we may assume that $\mathcal{A}_{S}$ is not a nontrivial tensor product. Since $\mathcal{A}_{S}$ is also nontrivial and indecomposable, Propositions 7.3 and 7.4 can not hold for $\mathcal{A}$. So $\mathcal{A}_{S}$ is dense. In particular, the Sylow $p$-subgroup $S_{2}$ of $S$ is an $\mathcal{A}_{S}$-subgroup. If $|\operatorname{rad}(\mathcal{A}_{S_{2}})|>1$, then $\mathcal{A}_{S}$ is decomposable or $p=3$ by Proposition 7.12, a contradiction to the assumption of the lemma. Therefore $|\operatorname{rad}(\mathcal{A}_{S_{2}})|=1$. Thus, $\mathcal{A}_{S}$ is normal by Proposition 7.11(1). ∎
Lemma 7.15.
Let $n=p^{k}$ and $\mathcal{A}$ an $S$-ring over $G$. The following statements hold.
  1. (1)
    If $P\cong C_{p}\leq G$ is an $\mathcal{A}$-subgroup, then $\operatorname{Aut}(\mathcal{A})^{P}\leq\operatorname{Hol}(P)$ or $\mathcal{A}$ is the $U/P$-wreath product (possibly, trivial) for an $\mathcal{A}$-subgroup $U$ such that $|U/P|\leq 4$.
  2. (2)
    If $U\cong E_{4}\times C_{p^{k-1}}\leq G$ is an $\mathcal{A}$-subgroup, then $\operatorname{Aut}(\mathcal{A})^{G/U}\leq\operatorname{Hol}(G/U)$ or $\mathcal{A}$ is the $U/L$-wreath product (possibly, trivial) for an $\mathcal{A}$-subgroup $L$ such that $|U/L|\leq 4$.
Proof.
If $p=3$, then $\operatorname{Hol}(C_{3})\cong S_{3}$ and the lemma is clear. Let $p\geq 5$. Let us prove Statement $(1)$. If $P\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times P)}$, then $\mathcal{A}$ is the $U/P$-wreath product for an $\mathcal{A}$-subgroup $U\leq H\times P$ and we are done. Further, we assume that there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times P)}$ such that
$\operatorname{rad}(X)\ngeq P.$
Let $R=\operatorname{rad}(X)$ and $V=\langle X\rangle$. Clearly, $R\leq H$. As $|R\cap P|=1$, the kernel of the action of $\operatorname{Aut}(\mathcal{A})^{P}$ on $P/R$ is trivial and hence it is enough to show that
$\operatorname{Aut}(\mathcal{A})^{P/R}\leq\operatorname{Hol}(P/R).$
Due to the definitions of $R$ and $V$, we conclude that $\mathcal{A}_{V/R}$ is indecomposable. Clearly, $P/R$ is an $\mathcal{A}_{V/R}$-subgroup of order $p$. One can see that $p^{2}~|~|V/R|$ because $X\nsubseteq H\times P$ and $R\leq H$. So $V/R$ satisfies all the conditions of Lemma 7.14 and consequently $\mathcal{A}_{V/R}$ is normal. Therefore
$\operatorname{Aut}(\mathcal{A})^{P/R}\leq\operatorname{Aut}(\mathcal{A}^{V/R})^{P/R}\leq\operatorname{Aut}(\mathcal{A}_{V/R})^{P/R}\leq(\operatorname{Hol}(V/R))^{P/R}\leq\operatorname{Hol}(P/R),$
where the third inclusion follows from the normality of $\mathcal{A}_{V/R}$, and we are done.
Now let us prove Statement $(2)$. If $U\cap D\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus U}$, then $\mathcal{A}$ is the $U/L$-wreath product for an $\mathcal{A}$-subgroup $L\geq(U\cap D)$ and we are done. Further, we assume that there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus U}$ such that
$\operatorname{rad}(X)\ngeq(U\cap D).$
Let $V=\langle X\rangle$. Clearly, $|G:V|\in\{1,2,4\}$ and $|V:(V\cap U)|=|G:U|=p$. Observe that $\operatorname{Aut}(\mathcal{A})^{G/U}$ and $\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}$ are isomorphic as permutation groups. Indeed, the kernel of the action of $\operatorname{Aut}(\mathcal{A})$ on $G/U$ contains $H_{r}$ and hence for every $f\in\operatorname{Aut}(\mathcal{A})$, there is $f^{\prime}\in\operatorname{Aut}(\mathcal{A})$ such that $f^{G/U}=(f^{\prime})^{G/U}$ and $V^{f^{\prime}}=V$. It is easy to see that for every $\Delta\in G/U$, there is a unique $\Delta^{\prime}\in V/(V\cap U)$ such that $\Delta^{\prime}\subseteq\Delta$. Therefore the mappings $f^{G/U}\mapsto(f^{\prime})^{V/(V\cap U)}$ from $\operatorname{Aut}(\mathcal{A})^{G/U}$ to $\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}$ and $\Delta\mapsto\Delta^{\prime}$ provide an isomorphism of permutation groups $\operatorname{Aut}(\mathcal{A})^{G/U}$ and $\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}$. Thus, it is enough to show that
$\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}\leq\operatorname{Hol}(V/(V\cap U)).$
Let $R=\operatorname{rad}(X)$. Clearly, the $S$-ring $\mathcal{A}_{V/R}$ is indecomposable and $(V\cap U)/R$ is an $\mathcal{A}_{V/R}$-subgroup of index $p$. Since $R\ngeq(U\cap D)$, we conclude that $p^{2}~|~|V/R|$. So $V/R$ satisfies all the conditions of Lemma 7.14 and hence $\mathcal{A}_{V/R}$ is normal. Therefore
$\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}\leq\operatorname{Aut}(\mathcal{A}^{V/R})^{V/(V\cap U)}\leq\operatorname{Aut}(\mathcal{A}_{V/R})^{V/(V\cap U)}\leq$
$\leq(\operatorname{Hol}(V/R))^{V/(V\cap U)}\leq\operatorname{Hol}(V/(V\cap U)),$
where the third inclusion follows from the normality of $\mathcal{A}_{V/R}$, and we are done. ∎
It should be mentioned that if $\mathcal{A}_{P}$ ($\mathcal{A}_{G/U}$, respectively) in Statement $(1)$ (Statement $(2)$, respectively) of Lemma 7.15 is nontrivial, then $\mathcal{A}_{P}$ ($\mathcal{A}_{G/U}$, respectively) is normal by Lemma 6.2 and the claim of Lemma 7.15 easily follows from the fact that an automorphism group of a normal $S$-ring is a subgroup of a holomorph of the underlying group. We also note that if the generalized wreath product in Statement $(1)$ or $(2)$ of Lemma 7.15 is trivial, then $|G|=4p$.
8. Proof of Theorem 1.2
Throughout this section, we keep the notation from the previous one. We start with a lemma concerned with a schurity of proper subgroups of $G$.
Lemma 8.1.
Every proper subgroup of $G$ is Schur.
Proof.
Let $N<G$. If $N$ is cyclic, then $N$ is Schur by [9, Theorem 1.1], whereas if $N\cong E_{4}$, then $N$ is Schur by [10, Theorem 1.2]. Otherwise, $N\cong E_{4}\times C_{r^{m}}$, where $r\in\{p,q\}$ and $1\geq m<k$, is Schur by induction on $m$ whose base follows from [10, Theorem 1.5]. ∎
Let $\mathcal{A}$ be a nontrivial $S$-ring over $G$. Let us prove that $\mathcal{A}$ is schurian. Theorem 7.1 holds for $\mathcal{A}$. We are done if $\mathcal{A}$ is cyclotomic. If $\mathcal{A}$ is a nontrivial tensor product, then $\mathcal{A}$ is schurian by Lemma 3.1 and Lemma 8.1. So we may assume that $\mathcal{A}$ is a nontrivial $S$-wreath product for some $\mathcal{A}$-section $S=U/L$ and one of Statements (1)(1)-$(4)$ from Theorem 7.1 holds. The $S$-rings $\mathcal{A}_{U}$ and $\mathcal{A}_{G/L}$ are schurian by Lemma 8.1. If Statement $(1)$ holds, i.e. $|S|\leq 2$, then $\mathcal{A}_{S}$ is $2$-minimal by Lemma 2.6 and hence $\mathcal{A}$ is schurian by Lemma 3.6. If Statement $(2)$ holds, i.e. $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{U}$ or $\mathcal{A}_{S}$ is $\otimes$-complemented in $\mathcal{A}_{G/L}$, then $\mathcal{A}$ is schurian by Lemma 3.7.
Now suppose that Statement $(3)$ or $(4)$ holds. Then one can choose $W\in\{U,G/L\}$ such that $\mathcal{A}_{W}$ is cyclotomic. This yields that $\mathcal{A}_{S}$ is also cyclotomic. Let $S_{H}$ and $S_{D}$ be the Sylow $2$-subgroup and the Hall $2^{\prime}$-subgroup of $S$, respectively. Note that $S_{H}$ and $S_{D}$ are $\mathcal{A}_{S}$-subgroups because $\mathcal{A}_{S}$ is cyclotomic. If $n=pq$ and $|S_{D}|=pq$, then $S=S_{D}$, $|U|=|G/L|=2pq$, and $|L|=|G/U|=2$. So $|W_{H}|=2$, where $W_{H}$ is the Sylow $2$-subgroup of $W$. Since $W$ is cyclotomic, $W_{H}$ is an $\mathcal{A}_{W}$-subgroup. Clearly, $\mathcal{A}_{W_{H}}=\mathbb{Z}W_{H}$. Together with Lemma 3.2(2), this implies that $\mathcal{A}_{W}\cong\mathbb{Z}W_{H}\otimes\mathcal{A}_{S}$. Therefore $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{U}$ or $\mathcal{A}_{S}$ is $\otimes$-complemented in $\mathcal{A}_{G/L}$ and hence $\mathcal{A}$ is schurian by Lemma 3.7.
In view of the above paragraph, we may assume further that $|S_{D}|$ is a $p$-power. The Hall $2^{\prime}$-subgroup of $W$ is denoted by $W_{D}$. Note that $W_{D}$ is an $\mathcal{A}_{W}$-subgroup because $\mathcal{A}_{W}$ is cyclotomic.
Lemma 8.2.
With the above notation, $\mathcal{A}_{S_{D}}$ is a cyclotomic $S$-ring with trivial radical.
Proof.
One can see that $\mathcal{A}_{S_{D}}$ is cyclotomic because $\mathcal{A}_{W}$ so is. Suppose that Statement $(3)$ of Theorem 7.1 holds. Then $|\operatorname{rad}(\mathcal{A}_{W})|=1$ and $|W_{H}|\leq 2$. Observe that $|\operatorname{rad}(\mathcal{A}_{W_{D}})|=1$. Indeed, Lemma 3.2(2) and Lemma 6.2 imply that $\mathcal{A}=\mathbb{Z}W_{H}\otimes\mathcal{A}_{W_{D}}$ and hence $|\operatorname{rad}(\mathcal{A}_{W_{D}})|=|\operatorname{rad}(\mathcal{A}_{W})|=1$. If $S_{D}=W_{D}$, then obviously $|\operatorname{rad}(\mathcal{A}_{S_{D}})|=1$. Otherwise, $|\operatorname{rad}(\mathcal{A}_{S_{D}})|=1$ by Lemma 6.9.
Now suppose that Statement $(4)$ of Theorem 7.1 holds. Then $W=U$, $\{e\}<L\leq D$, and $|\operatorname{rad}(\mathcal{A}_{W_{D}})|\in\{1,3\}$. As $\{e\}<L\leq D$, we conclude that $S_{D}\neq W_{D}$. Therefore $|\operatorname{rad}(\mathcal{A}_{S_{D}})|=1$ by Lemma 6.9. ∎
Lemma 8.3.
With the above notation, $\mathcal{A}_{S}$ is $2$-minimal unless $\mathcal{A}_{S}=\mathcal{A}_{S_{H}}\otimes\mathcal{A}_{S_{D}}$ and at least one of the following statements holds:
  1. (1)
    $|S_{H}|=4$ and $\mathcal{A}_{S_{H}}=\mathcal{T}_{S_{H}}$;
  2. (2)
    $|S_{D}|=p$ and $\mathcal{A}_{S_{D}}=\mathcal{T}_{S_{D}}$.
Proof.
Assume that $\mathcal{A}_{S}$ is not $2$-minimal. Clearly, $|S_{H}|\in\{1,2,4\}$. Let us prove that
(32) $\mathcal{A}_{S}=\mathcal{A}_{S_{H}}\otimes\mathcal{A}_{S_{D}}$
If $|S_{H}|\leq 2$, then $\mathcal{A}_{S_{H}}=\mathbb{Z}S_{H}$ and Eq. (32) follows from Lemma 3.2(2). If $|S_{H}|=4$ and Eq. (32) does not hold, then $\mathcal{A}_{S}$ is $2$-minimal by Proposition 7.11(1) and Lemma 8.2, a contradiction to the assumption.
As $\mathcal{A}_{S}$ is not $2$-minimal, Corollary 3.4 yields that at least one of the $S$-rings $\mathcal{A}_{S_{H}}$, $\mathcal{A}_{S_{D}}$ is not $2$-minimal. If $\mathcal{A}_{S_{H}}$ is not $2$-minimal, then Statement $(1)$ of the lemma holds by Lemma 2.6. If $\mathcal{A}_{S_{D}}$ is not $2$-minimal, then $\mathcal{A}_{S_{D}}$ is not normal by Lemma 6.5. Note that $|\operatorname{rad}(\mathcal{A}_{S_{D}})|=1$ by Lemma 8.2 and hence $\mathcal{A}_{S_{D}}=\mathcal{T}_{S_{D}}$ by Lemma 6.2. Since $\mathcal{A}_{S_{D}}$ is cyclotomic, we conclude that $|S_{D}|=p$. Thus, Statement $(2)$ of the lemma holds. ∎
It is worth to mention that the tensor product from Lemma 8.3 can be trivial, i.e. $S_{H}$ or $S_{D}$ can be trivial. In view of Lemma 3.6 and Lemma 8.3, we may assume further that
$\mathcal{A}_{S}=\mathcal{A}_{S_{H}}\otimes\mathcal{A}_{S_{D}}$
and Statement $(1)$ or $(2)$ of Lemma 8.3 holds. The canonical epimorphism from $G$ to $G/L$ is denoted by $\pi$ and the image of $X\subseteq G$ under $\pi$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu$.
Lemma 8.4.
If Statement $(3)$ of Theorem 7.1 holds, then $\mathcal{A}$ is schurian.
Proof.
By the condition of the lemma, $\mathcal{A}_{W}$ is a circulant cyclotomic $S$-ring with trivial radical. Clearly, $|W_{H}|\leq 2$ in this case. So $|S_{H}|\leq 2$. This implies that Statement $(2)$ of Lemma 8.3 holds for $\mathcal{A}_{S}$, i.e. $|S_{D}|=p$ and $\mathcal{A}_{S_{D}}=\mathcal{T}_{S_{D}}$. Let $V\in\{U,G/L\}\setminus\{W\}$. If $4\nmid|V|$, then $V$ is cyclic and 2||L|2~|~|L|. The latter yields that $2\nmid|S|$ and hence $|S|=|S_{D}|=p$. Therefore $\mathcal{A}$ is schurian by Lemma 6.6.
Now suppose that 4||V|4~|~|V|. Note that
$|V|=4p^{l}$
for some $l\geq 1$. This is obvious if $n=p^{k}$ and follows from $|S_{D}|=p$ and $V<G$ if $n=pq$. One can see that
$|S_{H}|=2$
because $4\nmid|W|$ and 4||V|4~|~|V|. Clearly, $\mathcal{A}_{S_{H}}=\operatorname{Cyc}(\operatorname{Aut}(S_{H}),S_{H})$ and $\mathcal{A}_{S_{D}}=\mathcal{T}_{S_{D}}=\operatorname{Cyc}(\operatorname{Aut}(S_{D}),S_{D})$. So
$\mathcal{A}_{S}=\mathcal{A}_{S_{H}}\otimes\mathcal{A}_{S_{D}}=\operatorname{Cyc}(\operatorname{Aut}(S_{H})\times\operatorname{Aut}(S_{D}),S)=\operatorname{Cyc}(\operatorname{Aut}(S),S)$
by Eq. (4). The latter is equivalent to
$\mathcal{A}_{S}=V(\operatorname{Hol}(S),S).$
As $\mathcal{A}_{W}$ is cyclotomic, there exists $K_{1}\leq\operatorname{Hol}(W)$ such that $\mathcal{A}_{W}=V(K_{1},W)$. It is easy to see that $K_{1}^{S}\leq\operatorname{Hol}(S)$ and $\mathcal{A}_{S}=V(K_{1}^{S},S)$. Since $\mathcal{A}_{S}$ is Cayley minimal (Lemma 6.4) and $\mathcal{A}_{S}=V(\operatorname{Hol}(S),S)$, we conclude that
(33) $K_{1}^{S}=\operatorname{Hol}(S).$
Let $\operatorname{Aut}(\mathcal{A}_{V})^{S_{D}}\leq\operatorname{Hol}(S_{D})$. Then $\operatorname{Aut}(\mathcal{A}_{V})^{S_{D}}=\operatorname{Hol}(S_{D})$ because $\operatorname{Hol}(S_{D})\in\mathcal{K}^{\min}(\mathcal{A}_{S_{D}})$ and $\mathcal{A}_{S_{D}}=V(\operatorname{Aut}(\mathcal{A}_{V})^{S_{D}},S_{D})$. Note that $\operatorname{Aut}(\mathcal{A}_{V})^{S_{H}}=\operatorname{Hol}(S_{H})$ because $|S_{H}|=2$. Therefore
$\operatorname{Aut}(\mathcal{A}_{V})^{S}=\operatorname{Aut}(\mathcal{A}_{V})^{S_{H}}\times\operatorname{Aut}(\mathcal{A}_{V})^{S_{D}}=\operatorname{Hol}(S_{H})\times\operatorname{Hol}(S_{D})=\operatorname{Hol}(S).$
Together with Eq. (33), this yields that $K_{1}^{S}=\operatorname{Aut}(\mathcal{A}_{V})^{S}$. Thus, $\mathcal{A}$ is schurian by Lemma 3.5.
In view of the above paragraph, we may assume that
$\operatorname{Aut}(\mathcal{A}_{V})^{S_{D}}\nleq\operatorname{Hol}(S_{D}).$
If $V=G/L$, then $\mathcal{A}_{V}$ is the $\mkern 1.5mu\overline{\mkern-1.5muU_{1}\mkern-1.5mu}\mkern 1.5mu/S_{D}$-wreath product for an $\mathcal{A}_{G/L}$-subgroup $\mkern 1.5mu\overline{\mkern-1.5muU_{1}\mkern-1.5mu}\mkern 1.5mu\geq S$ with $|\mkern 1.5mu\overline{\mkern-1.5muU_{1}\mkern-1.5mu}\mkern 1.5mu/S_{D}|\leq 4$ by Lemma 7.15(1) applied to $\mathcal{A}_{G/L}$, whereas if $V=U$, then $\mathcal{A}_{V}$ is the $U_{2}/L_{2}$-wreath product, where $U_{2}=S_{H}^{\pi^{-1}}$ and $L_{2}\leq L$ is an $\mathcal{A}_{U}$-subgroup with $|S_{H}^{\pi^{-1}}/L_{2}|\leq 4$, by Lemma 7.15(2) applied to $\mathcal{A}_{U}$.
Suppose that $|V|\neq 4p$. Then $\mathcal{A}$ is the nontrivial $S_{1}=U_{1}/L_{1}$-wreath product, where $U_{1}=\mkern 1.5mu\overline{\mkern-1.5muU_{1}\mkern-1.5mu}\mkern 1.5mu^{\pi^{-1}}$ and $L_{1}=(S_{D})^{\pi^{-1}}$, whenever $V=G/L$, and $\mathcal{A}$ is the nontrivial $S_{2}=U_{2}/L_{2}$-wreath product whenever $V=U$. As $|\mkern 1.5mu\overline{\mkern-1.5muU_{1}\mkern-1.5mu}\mkern 1.5mu/S_{D}|\leq 4$ ($|S_{H}^{\pi^{-1}}/L_{2}|\leq 4$, respectively), we obtain $|S_{1}|\leq 4$ ($|S_{2}|\leq 4$, respectively). If $|S_{i}|=4$, $i\in\{1,2\}$, then $\mathcal{A}_{S_{i}}\neq\mathcal{T}_{S_{i}}$ because $S_{D}$ is an $\mathcal{A}_{G/L}$-subgroup of order $2$. Therefore $\mathcal{A}_{S_{i}}$, $i\in\{1,2\}$, is $2$-minimal by Lemma 2.6. Thus, $\mathcal{A}$ is schurian by Lemma 3.6.
Now suppose that $|V|=4p$. If $\mathcal{A}_{V}$ is dense, then $\mathcal{A}$ is cyclotomic by [10, Case 33, p. 115]. Therefore $\mathcal{A}_{V}=V(K_{0},V)$ for some $K_{0}\leq\operatorname{Hol}(V)$. Clearly, $K_{0}^{S}\leq\operatorname{Hol}(S)$ and $\mathcal{A}_{S}=V(K_{0}^{S},S)$. Since $\mathcal{A}_{S}$ is Cayley minimal (Lemma 6.4) and $\mathcal{A}_{S}=V(\operatorname{Hol}(S),S)$, we conclude that $K_{0}^{S}=\operatorname{Hol}(S)$. Together with Eq. (33) and Lemma 3.5, this implies that $\mathcal{A}$ is schurian.
Let $\mathcal{A}$ be nondense. If $V=G/L$, then $\mathcal{A}_{V}=\mathcal{A}_{G/L}$ is the $S/S_{D}$-wreath product by [10, Case 11, p. 114] and hence $\mathcal{A}$ is the $S_{1}=U/L_{1}$-wreath product, where $L_{1}=(S_{D})^{\pi^{-1}}$. If $V=U$, then $\mathcal{A}_{V}=\mathcal{A}_{U}$ is the $S_{2}=H/L$-wreath product by [10, Case 22, p. 114] (in this case, $|L|=2$) and hence $\mathcal{A}$ is the $S_{2}$-wreath product. One can see that $|S_{1}|=|U/L_{1}|=|S/S_{D}|=|S_{H}|=2$ and $|S_{2}|=|H/L|=2$. Thus, $\mathcal{A}$ is schurian by Lemma 3.6. ∎
In view of Lemma 8.4, we may assume that Statement $(4)$ of Theorem 7.1 holds, i.e. $\mathcal{A}$ is dense, $U\geq H$, $\mathcal{A}_{U}$ is cyclotomic, $|\operatorname{rad}(\mathcal{A}_{U\cap D})|=1$ unless $n=3^{k}$ and $|\operatorname{rad}(\mathcal{A}_{U\cap D})|=3$, and $L\leq D$. In this case, $|S_{H}|=4$.
Lemma 8.5.
There exist $K_{1}\leq\operatorname{Aut}(\mathcal{A}_{U})$ and $K_{0}\leq\operatorname{Aut}(\mathcal{A}_{G/L})$ such that $K_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{U})$, $K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{G/L})$, and $K_{1}^{S_{H}}=K_{0}^{S_{H}}\in\mathcal{K}^{\min}(\mathcal{A}_{S_{H}})$.
Proof.
If $\mathcal{A}_{S_{H}}\neq\mathcal{T}_{S_{H}}$, then $\mathcal{A}_{S_{H}}$ is $2$-minimal by Lemma 2.6. So $\operatorname{Aut}(\mathcal{A}_{U})^{S_{H}}=\operatorname{Aut}(\mathcal{A}_{G/L})^{S_{H}}=\operatorname{Aut}(\mathcal{A}_{S_{H}})$ and the lemma holds for $K_{1}=\operatorname{Aut}(\mathcal{A}_{U})$ and $K_{0}=\operatorname{Aut}(\mathcal{A}_{G/L})$.
Further, we assume that $|S_{H}|=4$ and $\mathcal{A}_{S_{H}}=\mathcal{T}_{S_{H}}$. Due to Remark 2.7, there is a unique proper $2$-equivalent subgroup $K_{\min}$ of $\operatorname{Aut}(\mathcal{A}_{S_{H}})$ of order $12$ such that $K_{\min}\in\mathcal{K}^{\min}(\mathcal{A}_{S_{H}})$. If $\mathcal{A}_{U}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{U\cap D}$, then put $K_{1}=\operatorname{Aut}(\mathcal{A}_{U})$. In this case, $K_{1}^{S_{H}}=K_{\min}$ by Lemma 7.13. Otherwise, $\operatorname{Aut}(\mathcal{A}_{U})=\operatorname{Aut}(\mathcal{A}_{H})\times\operatorname{Aut}(\mathcal{A}_{U\cap D})$. Let $K_{H}\leq\operatorname{Sym}(H)$ such that $K_{H}\cong K_{\min}$. The group $K_{H}$ is $2$-equivalent to $\operatorname{Aut}(\mathcal{T}_{H})$ (see Remark 2.7) and hence $K_{1}=K_{H}\times\operatorname{Aut}(\mathcal{A}_{U\cap D})$ is $2$-equivalent to $\operatorname{Aut}(\mathcal{A}_{U})$. By the definition of $K_{1}$,
$K_{1}^{S_{H}}=(K_{H}\times\operatorname{Aut}(\mathcal{A}_{U\cap D}))^{S_{H}}=(K_{H})^{S_{H}}=K_{\min}.$
The groups $S_{H}$ and $D/L$ are $\mathcal{A}_{G/L}$-subgroups. If $\mathcal{A}_{G/L}\neq\mathcal{A}_{S_{H}}\otimes\mathcal{A}_{D/L}$, then put $K_{0}=\operatorname{Aut}(\mathcal{A}_{G/L})$; otherwise put $K_{0}=K_{\min}\times\operatorname{Aut}(\mathcal{A}_{D/L})$. The argument which is the same to the argument from the previous paragraph implies that $K_{0}$ is $2$-equivalent to $\operatorname{Aut}(\mathcal{A}_{G/L})$ and $K_{0}^{S_{H}}=K_{\min}=K_{1}^{S_{H}}$ as required. ∎
Let $K_{1}$ and $K_{0}$ are defined as in Lemma 8.5.
Lemma 8.6.
With the above notation, $K_{1}^{S_{D}}=K_{0}^{S_{D}}\in\mathcal{K}^{\min}(\mathcal{A}_{S_{D}})$ unless $\mathcal{A}$ is the $(H\times L)/L$- or $U/(U\cap D)$-wreath product.
Proof.
Due to Lemma 6.2 and Lemma 8.2, the $S$-ring $\mathcal{A}_{S_{D}}$ is normal or trivial over a group of order $p$. If $\mathcal{A}_{S_{D}}$ is normal or $|S_{D}|=3$, then $\mathcal{A}_{S_{D}}$ is $2$-minimal by Lemma 6.5 or Lemma 2.6, respectively. So $K_{1}^{S_{D}}=K_{0}^{S_{D}}=\operatorname{Aut}(\mathcal{A}_{S_{D}})\in\mathcal{K}^{\min}(\mathcal{A}_{S_{D}})$ as required. Further, we assume that
$|S_{D}|=p\geq 5~\text{and}~\mathcal{A}_{S_{D}}=\mathcal{T}_{S_{D}}.$
This assumption implies that $|(U\cap D):L|=p$.
If $\operatorname{Aut}(\mathcal{A}_{U})^{S_{D}}\leq\operatorname{Hol}(S_{D})$ and $\operatorname{Aut}(\mathcal{A}_{G/L})^{S_{D}}\leq\operatorname{Hol}(S_{D})$, then $K_{1}^{S_{D}}\leq\operatorname{Hol}(S_{D})$ and $K_{0}^{S_{D}}\leq\operatorname{Hol}(S_{D})$. Since $\operatorname{Hol}(S_{D})\in\mathcal{K}^{\min}(\mathcal{A}_{S_{D}})$, we obtain $K_{1}^{S_{D}}=K_{0}^{S_{D}}=\operatorname{Hol}(S_{D})$ as desired.
If $\operatorname{Aut}(\mathcal{A}_{U})^{S_{D}}\nleq\operatorname{Hol}(S_{D})$ or $\operatorname{Aut}(\mathcal{A}_{G/L})^{S_{D}}\nleq\operatorname{Hol}(S_{D})$, then $\mathcal{A}_{U}$ is the $(H\times L)/L$-wreath product or $\mathcal{A}_{G/L}$ is the $S/S_{D}$-wreath product, respectively, by Lemma 7.15. In the former case, $\mathcal{A}$ is the $(H\times L)/L$-wreath product, whereas in the latter one, $\mathcal{A}$ is the $U/(U\cap D)$-wreath product and we are done. ∎
By Lemma 8.5, we have $K_{1}^{S_{H}}=K_{0}^{S_{H}}\in\mathcal{K}^{\min}(\mathcal{A}_{S_{H}})$. If $\mathcal{A}$ is not the $T$-wreath product, where $T=U_{1}/L_{1}\in\{(H\times L)/L,U/(U\cap D)\}$, then Lemma 8.6 implies that $K_{1}^{S_{D}}=K_{0}^{S_{D}}\in\mathcal{K}^{\min}(\mathcal{A}_{S_{D}})$. Therefore $K_{1}^{S}=K_{0}^{S}$ by Lemma 3.3. Thus, $\mathcal{A}$ is schurian by Lemma 3.5.
Suppose that $\mathcal{A}$ is the $T$-wreath product. Clearly, $|T|=|H|=4$ and Statement $(4)$ of Theorem 7.1 holds for $\mathcal{A}$ and $T$. Lemma 8.5 yields that there are $M_{1}\leq\operatorname{Aut}(\mathcal{A}_{U_{1}})$ and $M_{0}\leq\operatorname{Aut}(\mathcal{A}_{G/L_{1}})$ such that $M_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{U_{1}})$, $M_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{G/L_{1}})$, and $M_{1}^{T}=M_{0}^{T}$. Again, $\mathcal{A}$ is schurian by Lemma 3.5.
9. Proof of Theorem 1.3
Since a section of a Schur group is a Schur group, to prove Theorem 1.3, it suffices to prove that the group $C_{8}\times C_{2}\times C_{p}$, where $p$ is an odd prime, is not a Schur group. Let $A\cong C_{8}$, $B\cong C_{2}$, $P\cong C_{p}$, and $G=A\times B\times P\cong C_{8}\times C_{2}\times C_{p}$. Denote generators of $A$ and $B$ by $a$ and $b$, respectively. Put $A_{1}=\langle a^{2}\rangle\cong C_{4}$, $A_{2}=\langle a^{4}\rangle\cong C_{2}$, $U_{1}=A_{1}\times P\cong C_{4p}$, and $U=A_{1}\times B\times P=B\times U_{1}\cong C_{2}\times C_{4p}$.
Let us construct a nonschurian $S$-ring over $G$. The group $M=\operatorname{Aut}(P)$ is a cyclic group of order p1p-1. Let $M_{0}$ be the subgroup of $M$ of index $2$. The canonical epimorphism from $M$ to $M/M_{0}$ is denoted by $\pi$. The group $N=\operatorname{Aut}(A_{1})$ is a cyclic group of order $2$. There exists a unique isomorphism $\varphi$ from $M/M_{0}\cong C_{2}$ to $N\cong C_{2}$. Put
$F=\{(\sigma,\tau)\in M\times N:~(\sigma^{\pi})^{\varphi}=\tau\}\leq\operatorname{Aut}(U_{1})$
and $\mathcal{A}_{1}=\operatorname{Cyc}(F,U_{1})$. Let $\mathcal{A}_{2}$ be an $S$-ring over $B\times P$ such that $\mathcal{A}_{2}=\mathcal{T}_{B}\otimes\mathcal{T}_{P}$.
One can see that $P$ is an $\mathcal{A}_{1}$- and $\mathcal{A}_{2}$-section and $(\mathcal{A}_{1})_{P}=(\mathcal{A}_{2})_{P}=\mathcal{T}_{P}$. Therefore one can form the $S$-ring
$\mathcal{A}_{12}=\mathcal{A}_{1}\wr_{U_{1}/A_{1}}\mathcal{A}_{2}$
over $U$.
Let $\mathcal{A}_{3}=\operatorname{Cyc}(\langle\sigma_{0}\rangle,A)$, where $\sigma_{0}\in A$ is such that $a^{\sigma_{0}}=a^{-1}$. It is easy to check in a straightforward way that the partition of $(A/A_{2})\times B\cong C_{4}\times C_{2}$ into the sets
$\{A_{2}\},~\{A_{2}a^{2}\},~\{A_{2}ab\},~\{A_{2}a^{3}b\},~\{A_{2}a,A_{2}a^{3}\},~\{A_{2}b,A_{2}a^{2}b\}$
defines the $S$-ring $\mathcal{A}_{4}$ over $(A/A_{2})\times B$. In fact,
$\mathcal{A}_{4}\cong\mathbb{Z}C_{4}\wr_{C_{2}}\mathbb{Z}E_{4}.$
One can see that $A/A_{2}$ is an $\mathcal{A}_{3}$- and $\mathcal{A}_{4}$-section and $(\mathcal{A}_{3})_{A/A_{2}}=(\mathcal{A}_{4})_{A/A_{2}}\cong\mathbb{Z}C_{2}\wr\mathbb{Z}C_{2}$. Therefore one can form the $S$-ring
$\mathcal{A}_{34}=\mathcal{A}_{3}\wr_{A/A_{2}}\mathcal{A}_{4}$
over $A\times B$.
The section $U/P=A_{1}\times B$ is an $\mathcal{A}_{12}$- and $\mathcal{A}_{34}$-section and
$(\mathcal{A}_{12})_{U/P}=(\mathcal{A}_{34})_{U/P}=(\mathbb{Z}A_{2}\wr\mathbb{Z}(A_{1}/A_{2}))\wr\mathbb{Z}B\cong(\mathbb{Z}C_{2}\wr\mathbb{Z}C_{2})\wr\mathbb{Z}C_{2}.$
Therefore one can form the $S$-ring
$\mathcal{A}=\mathcal{A}_{12}\wr_{U/P}\mathcal{A}_{34}$
over $G$. The basic sets of $\mathcal{A}$ are the following:
$X_{0}=\{e\},~X_{1}=\{a^{4}\},~X_{2}=\{a^{2},a^{6}\},~X_{3}=P^{\#},~X_{4}=a^{4}P^{\#},$
$Y_{1}=a^{2}P_{1}\cup a^{6}P_{2},~Y_{2}=a^{2}P_{2}\cup a^{6}P_{1},$
$Z_{1}=A_{1}b,~Z_{2}=A_{1}bP^{\#},$
$T_{1}=\{a,a^{7}\}P,~T_{2}=\{a^{3},a^{5}\}P,~T_{3}=\{ab,a^{5}b\}P,~T_{4}=\{a^{3}b,a^{7}b\}P,$
where $P_{1}$ and $P_{2}$ are the nontrivial orbits of $M_{0}$ on $P$.
Proposition 9.1.
The $S$-ring $\mathcal{A}$ is nonschurian.
Proof.
Assume the contrary. Then
(34) $\mathcal{S}(\mathcal{A})=\operatorname{Orb}(K,G),$
where $K=\operatorname{Aut}(\mathcal{A})_{e}$. Let $x_{1}\in P_{1}$ and $x_{2}\in P_{2}$. Since the elements $bx_{1}$ and $bx_{2}$ lie in the same basic set $Z_{2}$ of $\mathcal{A}$, there exists $f\in K$ such that $(bx_{1})^{f}=bx_{2}$. From Eqs. (2) and (34) it follows that
$(P^{\#}\setminus\{x_{1}\})^{f}=(A_{1}P^{\#}x_{1}\cap P^{\#})^{f}=(Z_{2}bx_{1}\cap X_{3})^{f}=Z_{2}bx_{2}\cap X_{3}=A_{1}P^{\#}x_{2}\cap P^{\#}=P^{\#}\setminus\{x_{2}\}.$
Since $X_{3}=P^{\#}$ is an orbit of $K$, we conclude that
$x_{1}^{f}=x_{2}.$
The above equality together with Eqs. (2) and (34) imply that
$((a^{2}P_{1}x_{1}\cup a^{6}P_{2}x_{1})\cap\{a^{2},a^{6}\})^{f}=(Y_{1}x_{1}\cap X_{2})^{f}=Y_{1}x_{2}\cap X_{2}=(a^{2}P_{1}x_{2}\cup a^{6}P_{2}x_{2})\cap\{a^{2},a^{6}\}.$
If $|M_{0}|$ is even, then $P_{1}=P_{1}^{-1}$ and $P_{2}=P_{2}^{-1}$ and hence the left-hand side and the right-hand side of the latter equality are equal to $\{a^{2}\}$ and $\{a^{6}\}$, respectively. If $|M_{0}|$ is odd, then $P_{2}=P_{1}^{-1}$ and hence the left-hand side and the right-hand side of the latter equality are equal to $\{a^{6}\}$ and $\{a^{2}\}$, respectively. Since $X_{2}=\{a^{2},a^{6}\}\in\operatorname{Orb}(K,G)$, we have
(35) $(a^{2})^{f}=a^{6}$
in both cases.
Now due to Eqs. (2), (34), and (35), we obtain
(36) $(Pa)^{f}=(T_{1}a^{2}\cap T_{1})^{f}=T_{1}a^{6}\cap T_{1}=Pa^{7}.$
On the other hand, since $(bx_{1})^{f}=bx_{2}$, we have
$(Pa)^{f}=(T_{3}bx_{1}\cap T_{1})^{f}=T_{3}bx_{2}\cap T_{1}=Pa,$
a contradiction to Eq. (36). ∎
Theorem 1.3 follows from Proposition 9.1.
10. Proof of Theorem 1.4
Let $p$ be an odd prime, $P\cong C_{p}$, $H\cong E_{16}$, and $G=H\times P$. If $p=3$, then $G$ is a Schur group by the computational results [33]. Further, we assume that $p\geq 5$. In this case, we are going to construct a nonschurian $S$-ring over $G$ and thereby to prove Theorem 1.4.
Let $a,b,c,d\in H$ be such that $\langle a,b,c,d\rangle=H$, $A=\langle a\rangle$, $B=\langle b\rangle$, $C=\langle c\rangle$, and $\sigma_{0}\in\operatorname{Aut}(H)$ such that
$\sigma_{0}:(a,b,c,d)\mapsto(a,ab,bc,cd),$
$M_{0}=\langle\sigma_{0}\rangle$, and $\mathcal{A}_{0}=\operatorname{Cyc}(M_{0},H)$. It is easy to verify in a straightforward way that $|\sigma_{0}|=4$ and the basic sets of $\mathcal{A}_{0}$ are the following:
$\{e\},~\{a\},~Ab,~(A\times B)c,~\{e,b,c,abc\}d,~\{a,ab,ac,bc\}d.$
One can see that $\mathcal{A}_{0}$ is a 22-$S$-ring, i.e. a size of every basic set is a $2$-power, and $\mathcal{A}_{0}$ is indecomposable. Therefore $\mathcal{A}_{0}$ is $2$-minimal by [17, Lemma 5.6]. Thus,
(37) $K_{0}=H_{r}\rtimes M_{0}$
for every $K_{0}\leq\operatorname{Sym}(H)$ such that $K_{0}\geq H_{r}$ and $K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{0})$. In particular, $\operatorname{Aut}(\mathcal{A}_{0})=H_{r}\rtimes M_{0}$.
Let $V=A\times B\times C$ and $U=V\times P$. As $p\geq 5$, the group $M_{P}=\operatorname{Aut}(P)\cong C_{p-1}$ has a nontrivial subgroup $M_{P}^{0}$ of index $2$. Let $\sigma_{1},\sigma_{2}\in\operatorname{Aut}(V)$ be such that
$\sigma_{1}:(a,b,c)\mapsto(a,ab,bc),~\sigma_{2}:(a,b,c)\mapsto(a,b,bc).$
Put $M_{V}=\langle\sigma_{1},\sigma_{2}\rangle$. It can be verified straightforwardly that $|\sigma_{1}|=4$, $|\sigma_{2}|=2$, and $\sigma_{2}\sigma_{1}\sigma_{2}=\sigma_{1}^{-1}$. So $M_{V}\cong D_{8}$.
Let $M^{0}_{V}=\langle\sigma_{1}^{2},\sigma_{2}\sigma_{1}\rangle$. Clearly, $M_{V}^{0}\cong E_{4}$ and hence $M_{V}/M_{V}^{0}\cong C_{2}$. Put
$M_{1}=\{(\sigma,\tau)\in M_{V}\times M_{P}:~(\sigma^{\pi_{1}})^{\psi}=\tau^{\pi_{2}}\}\leq\operatorname{Aut}(U),$
where $\pi_{1}$ and $\pi_{2}$ are the canonical epimorphisms from $M_{V}$ to $M_{V}/M_{V}^{0}$ and from $M_{P}$ to $M_{P}/M_{P}^{0}$, respectively, and $\psi$ is a unique isomorphism from $M_{V}/M_{V}^{0}\cong C_{2}$ to $M_{P}/M_{P}^{0}\cong C_{2}$, and $\mathcal{A}_{1}=\operatorname{Cyc}(M_{1},U)$. Due to the definition of $M_{1}$, we conclude that $V$ and $P$ are $\mathcal{A}_{1}$-subgroups and the basic sets of $\mathcal{A}_{1}$ are the following:
$\{e\},~\{a\},~Ab,~(A\times B)c,~P^{\#},~aP^{\#},~AbP^{\#},$
$AcP_{1}\cup AbcP_{2},~AcP_{2}\cup AbcP_{1},$
where $P_{1}$ and $P_{2}$ are the nontrivial orbits of $M_{P}^{0}$ on $P$.
One can see that $V$ is an $\mathcal{A}_{0}$- and $\mathcal{A}_{1}$-subgroup and $(\mathcal{A}_{0})_{V}=(\mathcal{A}_{1})_{V}\cong(\mathbb{Z}C_{2}\wr\mathbb{Z}C_{2})\wr\mathbb{Z}C_{2}$. Therefore one can form the $S$-ring
$\mathcal{A}=\mathcal{A}_{1}\wr_{S}\mathcal{A}_{0}$
over $G$, where $S=U/P$ and $\mathcal{A}_{0}$ is considered as an $S$-ring over $G/P\cong H$.
Proposition 10.1.
The $S$-ring $\mathcal{A}$ is nonschurian.
Proof.
Assume the contrary that $\mathcal{A}$ is schurian. Then there exist two groups $K_{1}\leq\operatorname{Sym}(U)$ and $K_{0}\leq\operatorname{Sym}(G/P)$ such that
$K_{1}\geq U_{r},~K_{0}\geq(G/L)_{r},~K_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{1}),~K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{0}),~\text{and}~K_{0}^{S}=K_{1}^{S}$
by Lemma 3.5. From Eq. (37) it follows that
(38) $K_{0}^{S}=S_{r}\rtimes M_{0}^{S}.$
Let $X\in\mathcal{S}(\mathcal{A}_{U})=\mathcal{S}(\mathcal{A}_{1})$ be such that
$X=AcP_{1}\cup AbcP_{2},$
$x\in P_{1}$, and $f\in M_{x}$, where $M=(K_{1})_{e}$. Due to $K_{0}^{S}=K_{1}^{S}$ and Eq. (38), we have $f^{S}\in M_{0}^{S}$. Eq. (2) and inclusions $(A\times B)c,X^{-1}\in\mathcal{S}(\mathcal{A})$ imply that
$(Ac)^{f}=(X^{-1}x\cap(A\times B)c)^{f}=X^{-1}x\cap(A\times B)c=Ac.$
Together with $f^{S}\in M_{0}^{S}$ and the definition of $M_{0}$, this yields that $f^{S}\in\langle\sigma_{0}^{2}\rangle$ and hence $b^{f}=b$. Since $H$ and $P$ are $\mathcal{A}$-subgroups, $Hx$ and $Pb$ are blocks of $\operatorname{Aut}(\mathcal{A})$. As $x^{f}=x$ and $b^{f}=b$, we conclude that $(Hx)^{f}=Hx$ and $(Pb)^{f}=Pb$. Therefore
$(bx)^{f}=(Hx\cap Pb)^{f}=Hx\cap Pb=bx.$
Thus, $f\in M_{bx}$ and consequently
(39) $M_{x}\leq M_{bx}.$
The definition of $M_{1}$ implies that the basic set of $\mathcal{A}$ containing $x$ is $P^{\#}$ and the basic set of $\mathcal{A}$ containing $bx$ is $Y=AbP^{\#}$. Clearly, |Y|=2(p1)|Y|=2(p-1). Due to the assumption, we have $P^{\#},Y\in\operatorname{Orb}(M,U)$. Therefore
$|M_{x}|=\frac{|M|}{p-1}>\frac{|M|}{2(p-1)}=\frac{|M|}{|Y|}=|M_{bx}|,$
a contradiction to Eq. (39). ∎
Now Theorem 1.4 follows from Proposition 10.1.
Acknowledgment: The author is very grateful to Dr. Matan Ziv-Av for the help with computer calculations.