### v3

# Classification of abelian Schur groups I

Abstract.
A finite group $G$ is called a Schur group if every Schur ring over $G$ is schurian, i.e. associated in a natural way with a subgroup of the symmetric group $\operatorname{Sym}(G)$ that contains all right translations of $G$. The list of all possible abelian Schur groups was obtained by Evdokimov, Kovács, and Ponomarenko in 2016. In two papers, we complete a classification of abelian Schur groups. In the present paper, we study schurity of several groups from the list. First, we prove that a direct product of the elementary abelian group of order 4 and a cyclic group, whose order is an odd prime power or a product of two distinct odd primes, is a Schur group. Second, we establish nonschurity of some other groups from the list.
Keywords: $S$-rings, Schur groups, permutation groups.
MSC: 05E30, 20B25.
The author was supported by the state contract of the Sobolev Institute of Mathematics (project number FWNF-2026-0011).
1. Introduction
Let $G$ be a finite group. A subring of the group ring $\mathbb{Z}G$ is called an $S$-ring (a Schur ring) over $G$ if it is a free $\mathbb{Z}$-module spanned by a partition of $G$, closed under taking inverse and containing the identity element $e$ of $G$ as a class. The notion of $S$-ring was introduced by Schur [32]. Using the $S$-ring approach, Schur proved that every primitive permutation group containing a regular cyclic subgroup of composite order is $2$-transitive which generalizes the Burnside theorem [2, p. 339]. The general theory of $S$-rings was developed by Wielandt (see [34, Chapter IV]). To date, $S$-rings are realized as a powerful tool for studying permutation groups containing regular subgroups, combinatorial Cayley objects, especially, isomorphisms of them, and representations of finite groups. For more details on $S$-rings and their applications, we refer the reader to [3, Section 2.4] and [22].
To study permutation groups having regular subgroups, Schur used $S$-rings whose partition of the underlying group consists of the orbits of a one-point stabilizer of an appropriate permutation group (see Section 2.3 for the exact definition). As was shown by Wielandt [34], not all $S$-rings can be constructed in this way. An $S$-ring is said to be schurian if it can [27]. The schurity (the property to be schurian) is one of the crucial properties of $S$-rings, closely related to the isomorphism problem for Cayley graphs (see [22, Section 5.3] for details).
A finite group $G$ is defined to be a Schur group if every $S$-ring over $G$ is schurian [27]. The main problem considered in the present paper is the following one.
Problem (Pöschel, 1974).
Determine all Schur groups.
In general, the above problem seems to be hard, in particular, because the number of $S$-rings over a given group can be exponential in the order of the group. There are several results on schurity of nonabelian groups [23, 26, 28, 30, 31]. However, most of the results on the schurity problem are concerned with abelian groups. The first result on schurity of abelian groups was obtained in paper [27]. Namely, it was proved that cyclic $p$-groups of odd order are Schur. This result was extended to the case $p=2$ in paper [12]. In paper [18], the schurity of cyclic groups of order $pq$, where $p$ and $q$ are distinct primes, was proved. The above results were used later for solving the isomorphism problem for Cayley graphs over the corresponding groups [17, 18].
A complete classification of cyclic Schur groups was obtained in paper [10]. Namely, it was proved that a cyclic group of order $n$ is Schur if and only if $n$ belongs to one of the following families of integers:
$p^{k},~pq^{k},~2pq^{k},~pqr,~2pqr,$
where $p$, $q$, and $r$ are distinct primes and $k\geq 0$ is an integer.
Investigations of schurity of abelian groups was continued in paper [11]. At first, necessary and sufficient condition of schurity for elementary abelian groups was obtained there. Namely, an elementary abelian noncyclic group of order $n$ is Schur if and only if
$n\in\{4,8,9,16,27,32\}.$
Further, it was shown that an abelian Schur group which is neither cyclic nor elementary abelian belongs to one of the explicitly given infinite families of groups. Given a positive integer $n$, the cyclic and elementary abelian groups of order $n$ are denoted by $C_{n}$ and $E_{n}$, respectively.
Theorem 1.1.
[11, Theorem 1.3] An abelian Schur group, which is neither cyclic nor elementary abelian, is isomorphic to a group from one of the following nine families:
  1. (1)
    $C_{2}\times C_{2^{k}}$, $C_{2p}\times C_{2^{k}}$, $E_{4}\times C_{p^{k}}$, $E_{4}\times C_{pq}$, $E_{16}\times C_{p}$,
  2. (2)
    $C_{3}\times C_{3^{k}}$, $C_{6}\times C_{3^{k}}$, $E_{9}\times C_{q}$, $E_{9}\times C_{2q}$,
where $p$ and $q$ are distinct primes, $p\neq 2$, and $k\geq 1$ is an integer.
In the same paper, it was proved that the groups $E_{4}\times C_{p}$, where $p$ is an odd prime, are Schur groups. A schurity property of the groups $C_{2}\times C_{2^{k}}$ and $C_{3}\times C_{3^{k}}$ was established in papers [23] and [29], respectively. In paper [25], it was checked that $E_{9}\times C_{q}$ is a Schur group for every prime $q$.
In two papers, we study the shurity problem for all groups from Theorem 1.1 whose schurity is unknown. In the present paper, we deal with the most laborious case, namely, with the groups $E_{4}\times C_{p^{k}}$ and $E_{4}\times C_{pq}$. We prove the following theorem.
Theorem 1.2.
The groups $E_{4}\times C_{p^{k}}$ and $E_{4}\times C_{pq}$, where $p$ and $q$ are odd primes and $k\geq 1$, are Schur groups.
To prove Theorem 1.2, we characterize all $S$-rings over the groups from this theorem.
Theorem 1.3.
Every nontrivial $S$-ring over $E_{4}\times C_{p^{k}}$ or $E_{4}\times C_{pq}$, where $p$ and $q$ are odd primes and $k\geq 1$, is cyclotomic or a nontrivial tensor or generalized wreath product.
A similar statement was proved for an arbitrary cyclic group in papers [20, 21] and for some other abelian groups close to cyclic groups in papers [11, 22, 25, 29]. On the other hand, the above statement does not hold, e.g., for elementary abelian groups (see [34]). It seems interesting to characterize all groups whose all nontrivial $S$-rings are cyclotomic or nontrivial tensor or generalized wreath products (see Subsection 2.3 and Section 3 for the definitions).
Perhaps, the most surprising fact concerning the groups from Theorem 1.1 is that not all of them are Schur groups. This is demonstrated by two theorems below.
Theorem 1.4.
The group $C_{2p}\times C_{2^{k}}$, where $p$ is an odd prime and $k\geq 3$, is not a Schur group.
Since the class of Schur groups is closed under taking subgroups, it suffices to prove Theorem 1.4 only for $k=3$. It should be noted that if $k=1,2$, then the group from Theorem 1.4 is a Schur group. Indeed, the schurity of the group $E_{4}\times C_{p}$ (case $k=1$) follows from [11, Theorem 1.4] as was mentioned before, whereas the schurity of the group $C_{2p}\times C_{4}$ (case $k=2$) will be verified in the subsequent paper together with the schurity of other abelian groups of order $8p$, where $p$ is an odd prime.
One more family of non-Schur groups from Theorem 1.1 is presented by the theorem below.
Theorem 1.5.
Let $p$ be an odd prime. The group $E_{16}\times C_{p}$ is a Schur group if and only if $p=3$.
In the proofs of Theorems 1.4 and 1.5, we construct new nonschurian $S$-rings over the groups $C_{2p}\times C_{8}$ and $E_{16}\times C_{p}$, respectively. These nonschurian $S$-rings are generalized wreath products of $S$-rings over proper subgroups of the above groups (see Sections 9 and 10). The nonschurian $S$-ring over $C_{2p}\times C_{8}$ from Section 9 in case $p=3$ was found by computer calculations using the package [16] and appeared in paper [35].
We finish the introduction with a brief outline of the paper. Section 2 contains a necessary background of $S$-rings, especially, of the multiplier theorems and isomorphisms. In Section 3, we provide an information on three constructions of $S$-rings, namely, tensor, generalized wreath, and star products. Several required facts on dual $S$-rings over the group of all irreducible complex characters of an abelian group are given in Section 4. Recent results on $S$-rings over an abelian group having a Sylow subgroup of prime order are provided in Section 5. In Section 6, we give a background of $S$-rings over cyclic groups, especially, over cyclic $p$-groups. In Section 7, we obtain a description of $S$-rings over the groups $E_{4}\times C_{p^{k}}$ and $E_{4}\times C_{pq}$ (Theorem 7.1) that implies Theorem 1.3. In Sections 8, 9, and 10, we prove Theorems 1.21.4 and 1.5, respectively.
Notation.
The identity element and the set of all nonidentity elements of a group $G$ are denoted by $e$ and $G^{\#}$, respectively.
The projections of $X\subseteq G_{1}\times G_{2}$ to $G_{1}$ and $G_{2}$ are denoted by $X_{G_{1}}$ and $X_{G_{2}}$, respectively.
If $X\subseteq G$, then the element $\sum_{x\in X}{x}$ of the group ring $\mathbb{Z}G$ is denoted by $\underline{X}$.
The set $\{x^{-1}:~x\in X\}$ is denoted by $X^{-1}$.
The subgroup of $G$ generated by $X$ is denoted by $\langle X\rangle$; we also set $\operatorname{rad}(X)=\{g\in G:\ gX=Xg=X\}$.
The set $\{(g,xg):~x\in X,~g\in G\}\subseteq G^{2}$ is denoted by $r(X)$.
If $m\in\mathbb{Z}$, then the set $\{x^{m}:~x\in X\}$ is denoted by $X^{(m)}$.
If $G$ is abelian and $m\in\mathbb{Z}$ coprime to $|G|$, then the automorphism of $G$ which maps every $g\in G$ to $g^{m}$ is denoted by $\sigma_{m}$.
The group of all permutations of a set $\Omega$ is denoted by $\operatorname{Sym}(\Omega)$.
The subgroup of $\operatorname{Sym}(G)$ induced by all right multiplications of $G$ is denoted by $G_{r}$.
If $f\in\operatorname{Sym}(\Omega)$ and $\Delta\subseteq\Omega$ is such that $\Delta^{f}=\Delta$, then the permutation of $\Delta$ induced by $f$ is denoted $f^{\Delta}$.
For a set $\Delta\subseteq\operatorname{Sym}(G)$ and a section $S=U/L$ of $G$ we set
$\Delta^{S}=\{f^{S}:~f\in\Delta,~S^{f}=S\},$
where $S^{f}=S$ means that $f$ permutes the $L$-cosets in $U$ and $f^{S}$ denotes the bijection of $S$ induced by $f$.
If $K\leq\operatorname{Sym}(\Omega)$ and $\alpha\in\Omega$, then the set of all orbits of $K$ on $\Omega$ and the stabilizer of $\alpha$ in $K$ are denoted by $\operatorname{Orb}(K,\Omega)$ and $K_{\alpha}$, respectively.
The set of all positive divisors of $n\in\mathbb{Z}$ is denoted by $\operatorname{Div}(n)$.
Given a cyclic group $D$ of order $n$ and $l\in\operatorname{Div}(n)$, the subgroup of $D$ of order $l$ is denoted by $D_{l}$ and the set of all generators of $D_{l}$ is denoted by $D_{l}^{*}$.
If $X\subseteq H\times D$, where $D$ is a cyclic group of order $n$, $h\in H$, and $l\in\operatorname{Div}(n)$, then put $X_{h}=h^{-1}X\cap D$ and $X_{h,l}=h^{-1}X\cap D_{l}^{*}$.
2. $S$-rings
In this section, we provide a necessary background of $S$-rings. In general, we follow [11, 23, 29], where most of the material is contained.
2.1. Definitions and basic facts
A subring $\mathcal{A}\subseteq\mathbb{Z}G$ is called an $S$-ring (a Schur ring) over $G$ if there exists a partition $\mathcal{S}=\mathcal{S}(\mathcal{A})$ of $G$ such that:
$(1)$ $\{e\}\in\mathcal{S}$;
$(2)$ if $X\in\mathcal{S}$, then $X^{-1}\in\mathcal{S}$;
$(3)$ $\mathcal{A}=\operatorname{Span}_{\mathbb{Z}}\{\underline{X}:\ X\in\mathcal{S}\}$.
The elements of $\mathcal{S}$ are called the basic sets of $\mathcal{A}$. The number of basic sets is called the rank of $\mathcal{A}$ and denoted by $\operatorname{rk}(\mathcal{A})$. Clearly, the group ring $\mathbb{Z}G$ is an $S$-ring. If $|G|\neq 1$, then the partition $\{\{e\},G^{\#}\}$ defines the $S$-ring $\mathcal{T}_{G}$ of rank $2$ over $G$. It is easy to check that if $X,Y\in\mathcal{S}(\mathcal{A})$, then $XY\in\mathcal{S}(\mathcal{A})$ whenever $|X|=1$ or $|Y|=1$. The $S$-ring $\mathcal{A}$ is called symmetric if $X=X^{-1}$ for every $X\in\mathcal{S}(\mathcal{A})$ and antisymmetric if for every $X\in\mathcal{S}(\mathcal{A})$ the equality $X=X^{-1}$ implies that $X=\{e\}$.
Let $X,Y\in\mathcal{S}$. If $Z\in\mathcal{S}$, then by Condition $(3)$, the number of distinct representations of $z\in Z$ in the form $z=xy$ with $x\in X$ and $y\in Y$ does not depend on the choice of $z\in Z$. Denote this number by $c^{Z}_{XY}$. One can see that
$\underline{X}\cdot\underline{Y}=\sum_{Z\in\mathcal{S}(\mathcal{A})}c^{Z}_{XY}\underline{Z}.$
Therefore the numbers $c^{Z}_{XY}$ are the structure constants of $\mathcal{A}$ with respect to the basis $\{\underline{X}:\ X\in\mathcal{S}\}$.
A set $T\subseteq G$ is called an $\mathcal{A}$-set if $\underline{T}\in\mathcal{A}$. If $T$ is an $\mathcal{A}$-set, then put
$\mathcal{S}(\mathcal{A})_{T}=\{X\in\mathcal{S}(\mathcal{A}):~X\subseteq T\}.$
A subgroup $H\leq G$ is called an $\mathcal{A}$-subgroup if $H$ is an $\mathcal{A}$-set. The $S$-ring $\mathcal{A}$ is called primitive if there is no a nontrivial proper $\mathcal{A}$-subgroup of $G$ and imprimitive otherwise. One can verify that for every $\mathcal{A}$-set $X$, the groups $\langle X\rangle$ and $\operatorname{rad}(X)$ are $\mathcal{A}$-subgroups. The set of all $\mathcal{A}$-subgroups of $G$ is denoted by $\mathcal{H}(\mathcal{A})$. It is easy to see that if $X\subseteq\operatorname{rad}(Y)$, then
(1) $\underline{X}\cdot\underline{Y}=\underline{Y}\cdot\underline{X}=|X|\underline{Y}.$
Lemma 2.1.
[11, Lemma 2.1] Let $\mathcal{A}$ be an $S$-ring over a group $G$, $H$ an $\mathcal{A}$-subgroup of $G$, and $X\in\mathcal{S}(\mathcal{A})$. Then the number $|X\cap Hx|$ does not depend on $x\in X$.
Lemma 2.2.
[23, Theorem 2.6] Let $X$ be a basic set of an $S$-ring $\mathcal{A}$ over a group $G$. Suppose that $H\leq\operatorname{rad}(X\setminus H)$ for some subgroup $H$ of $G$ such that $X\cap H\neq\varnothing$ and $X\setminus H\neq\varnothing$. Then $\operatorname{rad}(X)\leq H$ and $X=\langle X\rangle\setminus\operatorname{rad}(X)$.
Let $\{e\}\leq L\unlhd U\leq G$. A section $U/L$ is called an $\mathcal{A}$-section if $U$ and $L$ are $\mathcal{A}$-subgroups. If $S=U/L$ is an $\mathcal{A}$-section, then the module
$\mathcal{A}_{S}=\operatorname{Span}_{\mathbb{Z}}\left\{\underline{X}^{\pi}:~X\in\mathcal{S}(\mathcal{A}),~X\subseteq U\right\},$
where $\pi:U\rightarrow U/L$ is the canonical epimorphism, is an $S$-ring over $S$.
2.2. Multiplier theorems
Let $G$ be an abelian group. If $X,Y\subseteq G$ are such that $Y^{(m)}=X$ for some $m\in\mathbb{Z}$ coprime to $|G|$, then we say that $X$ and $Y$ are rationally conjugate. In this case, if $X\in\operatorname{Orb}(K,G)$ for some $K\leq\operatorname{Aut}(G)$, then $Y\in\operatorname{Orb}(K,G)$. The following two statements are known as the first and second Schur theorems on multipliers (see [34, Theorem 23.9, (a)-(b)]).
Lemma 2.3.
Let $\mathcal{A}$ be an $S$-ring over an abelian group $G$. Then $X^{(m)}\in\mathcal{S}(\mathcal{A})$ for every $X\in\mathcal{S}(\mathcal{A})$ and every $m\in\mathbb{Z}$ coprime to $|G|$.
Lemma 2.4.
Let $\mathcal{A}$ be an $S$-ring over an abelian group $G$, $p$ a prime divisor of $|G|$, and $H=\{g\in G:g^{p}=e\}$. Then for every $\mathcal{A}$-set $X$, the set
$X^{[p]}=\{x^{p}:x\in~X,~|X\cap Hx|\not\equiv 0\mod p\}$
is an $\mathcal{A}$-set.
Below, we deduce from Lemma 2.3 one more lemma which will be used further in the proof of Theorem 1.2.
Lemma 2.5.
Let $G=H\times D$, where $H$ is an abelian group and $D\cong C_{n}$ with $n$ coprime to $|H|$, $\mathcal{A}$ an $S$-ring over $G$, and $X\in\mathcal{S}(\mathcal{A})$. Then there exists $K^{D}\leq\operatorname{Aut}(D)$ satisfying the following: $X_{h,l}=h^{-1}X\cap D_{l}^{*}\in\operatorname{Orb}(K^{D},D)$ for all $h\in H$ and $l\in\operatorname{Div}(n)$ such that $X_{h,l}\neq\varnothing$.
Proof.
Since $|H|$ and $n$ are coprime, $\operatorname{Aut}(G)=\operatorname{Aut}(H)\times\operatorname{Aut}(D)$. Let $K$ be the setwise stabilizer of $X$ in the group $\{\operatorname{id}_{H}\}\times\operatorname{Aut}(D)\leq\operatorname{Aut}(G)$ and $h\in H$ and $l\in\operatorname{Div}(n)$ be such that $X_{h,l}\neq\varnothing$. By the definition of $K$, the set $X_{h,l}$ is $K$-invariant. On the other hand, for all $x,y\in X_{h,l}$, there is an integer $m$ coprime to $n$ such that $x^{\sigma_{m}}=x^{m}=y$ and $m\equiv 1\mod|H|$ and hence $(hx)^{m}=hy$. The latter implies that $X^{(m)}\cap X\neq\varnothing$. So $X^{(m)}=X$ by Lemma 2.3. Therefore $\sigma_{m}\in K$ and consequently $K$ is transitive on $X_{h,l}$. Thus, $X_{h,l}$ is an orbit of $K$ and hence of the group $K^{D}$ induced by $K$ on $D$. ∎
2.3. Isomorphisms and schurity
Put $\mathcal{R}(\mathcal{A})=\{r(X):~X\in\mathcal{S}(\mathcal{A})\}$. Let $\mathcal{A}^{\prime}$ be an $S$-ring over a group $G^{\prime}$. A bijection $f$ from $G$ to $G^{\prime}$ is called a (combinatorial) isomorphism from $\mathcal{A}$ to $\mathcal{A}^{\prime}$ if $\mathcal{R}(\mathcal{A})^{f}=\mathcal{R}(\mathcal{A}^{\prime})$, where $\mathcal{R}(\mathcal{A})^{f}=\{r(X)^{f}:~X\in\mathcal{S}(\mathcal{A})\}$ and $r(X)^{f}=\{(g^{f},h^{f}):~(g,h)\in r(X)\}$. If there exists an isomorphism from $\mathcal{A}$ to $\mathcal{A}^{\prime}$, then $\mathcal{A}$ and $\mathcal{A}^{\prime}$ are said to be isomorphic.
A bijection $f\in\operatorname{Sym}(G)$ is defined to be a (combinatorial) automorphism of $\mathcal{A}$ if $r(X)^{f}=r(X)$ for every $X\in\mathcal{S}(\mathcal{A})$. The set of all automorphisms of $\mathcal{A}$ forms the group called the automorphism group of $\mathcal{A}$ and denoted by $\operatorname{Aut}(\mathcal{A})$. One can see that $\operatorname{Aut}(\mathcal{A})\geq G_{r}$. If $f\in\operatorname{Aut}(\mathcal{A})$, then
(2) $(Xy)^{f}=Xy^{f}$
for every $X\in\mathcal{S}(\mathcal{A})$ and $y\in G$. If $H$ is an $\mathcal{A}$-subgroup of $G$, then the set of all right $H$-cosets is an imprimitivity system of $\operatorname{Aut}(\mathcal{A})$.
Let $K$ be a subgroup of $\operatorname{Sym}(G)$ containing $G_{r}$. Schur proved in paper [32] that the $\mathbb{Z}$-submodule
$V(K,G)=\operatorname{Span}_{\mathbb{Z}}\{\underline{X}:~X\in\operatorname{Orb}(K_{e},G)\}$
is an $S$-ring over $G$. An $S$-ring $\mathcal{A}$ over $G$ is called schurian if $\mathcal{A}=V(K,G)$ for some $K\leq\operatorname{Sym}(G)$ with $K\geq G_{r}$. One can see that $\mathcal{T}_{G}=V(K,G)$ for every $2$-transitive group $K\leq\operatorname{Sym}(G)$ containing $G_{r}$ and hence $\mathcal{T}_{G}$ is schurian. The group $G$ is called a Schur group if every $S$-ring over $G$ is schurian.
Let $\mathcal{A}=V(K,G)$ for some $K\leq\operatorname{Sym}(G)$ such that $K\geq G_{r}$. One can see that $\operatorname{Aut}(\mathcal{A})\geq K$. If $S$ is an $\mathcal{A}$-section, then $\mathcal{A}_{S}=V(K^{S},S)$. This yields that $\mathcal{A}_{S}$ is also schurian. Therefore a section (in particular, a subgroup) of a Schur group is also a Schur group.
One can verify that $\mathcal{A}$ is schurian if and only if
$\mathcal{A}=V(\operatorname{Aut}(\mathcal{A}),G)$
or, equivalently, $\mathcal{S}(\mathcal{A})=\operatorname{Orb}(\operatorname{Aut}(\mathcal{A})_{e},G)$.
The $S$-ring $\mathcal{A}$ is said to be normal if $G_{r}$ is normal in $\operatorname{Aut}(\mathcal{A})$ or, equivalently,
$\operatorname{Aut}(\mathcal{A})\leq\operatorname{Hol}(G)=G_{r}\rtimes\operatorname{Aut}(G).$
Let $K\leq\operatorname{Aut}(G)$. Then $\operatorname{Orb}(K,G)$ forms a partition of $G$ that defines an $S$-ring $\mathcal{A}$ over $G$. In this case, $\mathcal{A}$ is called cyclotomic and denoted by $\operatorname{Cyc}(K,G)$. One can see that $\mathcal{A}=V(G_{r}K,G)$. So every cyclotomic $S$-ring is schurian. If $S$ is an $\mathcal{A}$-section, then $\mathcal{A}_{S}=\operatorname{Cyc}(K^{S},G)$, i.e. $\mathcal{A}_{S}$ is also cyclotomic. Note that a normal $S$-ring can be nonschurian and hence noncyclotomic as well as a cyclotomic $S$-ring can be nonnormal. However, if a normal $S$-ring is schurian, then it is also cyclotomic.
2.4. Minimal $S$-rings
Two permutation groups $K_{1}$ and $K_{2}$ on a set $\Omega$ are called $2$-equivalent if $\operatorname{Orb}(K_{1},\Omega^{2})=\operatorname{Orb}(K_{2},\Omega^{2})$ (here we assume that $K_{1}$ and $K_{2}$ act on $\Omega^{2}$ componentwise). In this case, we write $K_{1}\approx_{2}K_{2}$. The relation $\approx_{2}$ is an equivalence relation on the set of all subgroups of $\operatorname{Sym}(\Omega)$. Every $\approx_{2}$-equivalence class has the unique maximal element and may have more than one minimal elements with respect to inclusion (for the latter, see Example $1$). Given $K\leq\operatorname{Sym}(\Omega)$, the unique maximal element from the class containing $K$ is called the $2$-closure of $K$ and denoted by $K^{(2)}$.
Let $K\leq\operatorname{Sym}(G)$ with $K\geq G_{r}$ and $\mathcal{A}=V(K,G)$. Then $K^{(2)}=\operatorname{Aut}(\mathcal{A})$ and $V(K^{\prime},G)=\mathcal{A}$ if and only if $K^{\prime}\approx_{2}K$ for every $K^{\prime}\leq\operatorname{Sym}(G)$ with $K^{\prime}\geq G_{r}$. A schurian $S$-ring $\mathcal{A}$ over $G$ is said to be $2$-minimal if
$\{K\leq\operatorname{Sym}(G):~K\geq G_{r},~K\approx_{2}\operatorname{Aut}(\mathcal{A})\}=\{\operatorname{Aut}(\mathcal{A})\}.$
The set of all minimal with respect to inclusion permutation groups $K^{\prime}\leq\operatorname{Sym}(G)$ such that $K^{\prime}\geq G_{r}$ and $V(K^{\prime},G)=\mathcal{A}$ is denoted by $\mathcal{K}^{\min}(\mathcal{A})$. Note that $\mathcal{A}$ is $2$-minimal if and only if $\operatorname{Aut}(\mathcal{A})$ is $2$-isolated in the sense of [15].
Example 1. This example demonstrates that $\mathcal{K}^{\min}(\mathcal{A})$ may contain more than one element. Let $p\geq 5$ be an odd prime, $G\cong C_{p}$, and $\mathcal{A}=\mathcal{T}_{G}$. Let $K_{1}=\operatorname{Hol}(G)$. Clearly, $K_{1}\geq G_{r}$ and $\mathcal{A}=V(K_{1},G)$. One can see that the one-point stabilizer $(K_{1})_{e}=\operatorname{Aut}(G)$ acts regularly on $G^{\#}$. So $K_{1}\in\mathcal{K}^{\min}(\mathcal{A})$.
Let $K_{2}=\operatorname{Alt}(G)$. Since $p$ is odd, every element from $G_{r}$ is an even permutation of $G$ and hence $K_{2}\geq G_{r}$. The group $K_{2}$ is (p2)(p-2)-transitive. Together with $p\geq 5$, this implies that $K_{2}$ is $2$-transitive and consequently $\mathcal{A}=V(K_{2},G)$. As $p$ is odd, a generator of $\operatorname{Aut}(G)$ is an odd permutation of $G$. So $K_{1}\nleq K_{2}$. Therefore $K_{2}^{\prime}\neq K_{1}$, where $K_{2}^{\prime}\in\mathcal{K}^{\min}(\mathcal{A})$ with $K_{2}^{\prime}\leq K_{2}$. Thus, $|\mathcal{K}^{\min}(\mathcal{A})|\geq 2$.
The next lemma and remark after it easily follow from computer calculations using the package [16].
Lemma 2.6.
Let $\mathcal{A}$ be an $S$-ring over a group $G$. If $|G|\leq 3$ or $|G|=4$ and $\mathcal{A}\neq\mathcal{T}_{G}$, then $\mathcal{A}$ is $2$-minimal.
Remark 2.7.
Let $|G|=4$. Then $\operatorname{Aut}(\mathcal{T}_{G})=\operatorname{Sym}(G)\cong\operatorname{Sym}(4)$ and $\mathcal{T}_{G}=V(\operatorname{Sym}(G),G)=V(\operatorname{Alt}(G),G)$. This implies that $\mathcal{T}_{G}$ is not $2$-minimal. The group $\operatorname{Alt}(G)\cong\operatorname{Alt}(4)$ of order $12$ is a unique proper subgroup of $\operatorname{Aut}(\mathcal{T}_{G})=\operatorname{Sym}(G)$ which is $2$-equivalent to $\operatorname{Aut}(\mathcal{T}_{G})$ and hence $\mathcal{K}^{\min}(\mathcal{T}_{G})=\{\operatorname{Alt}(G)\}$.
Two groups $K_{1},K_{2}\leq\operatorname{Aut}(G)$ are said to be Cayley equivalent if $\operatorname{Orb}(K_{1},G)=\operatorname{Orb}(K_{2},G)$. In this case, we write $K_{1}\approx_{\operatorname{Cay}}K_{2}$. If $\mathcal{A}=\operatorname{Cyc}(K,G)$ for some $K\leq\operatorname{Aut}(G)$, then $\operatorname{Aut}_{G}(\mathcal{A})=\operatorname{Aut}(\mathcal{A})\cap\operatorname{Aut}(G)$ is the largest group which is Cayley equivalent to $K$. A cyclotomic $S$-ring $\mathcal{A}$ over $G$ is called Cayley minimal if
$\{K\leq\operatorname{Aut}(G):~K\approx_{\operatorname{Cay}}\operatorname{Aut}_{G}(\mathcal{A})\}=\{\operatorname{Aut}_{G}(\mathcal{A})\}.$
It is easy to see that $\mathbb{Z}G$ is $2$-minimal and Cayley minimal.
3. Tensor, generalized wreath, and star products
This section contains a necessary information on three main constructions for producing $S$-rings.
3.1. Tensor product
Let $\mathcal{A}$ be an $S$-ring over a group $G$. Suppose that $G_{1}$ and $G_{2}$ are $\mathcal{A}$-subgroups such that $G=G_{1}\times G_{2}$. The $S$-ring $\mathcal{A}$ is defined to be a tensor product of $\mathcal{A}_{1}=\mathcal{A}_{G_{1}}$ and $\mathcal{A}_{2}=\mathcal{A}_{G_{2}}$ if
$\mathcal{S}(\mathcal{A})=\mathcal{S}(\mathcal{A}_{1})\otimes\mathcal{S}(\mathcal{A}_{2})=\{X_{1}\times X_{2}:~X_{1}\in\mathcal{S}(\mathcal{A}_{1}),~X_{2}\in\mathcal{S}(\mathcal{A}_{2})\}.$
In this case, we write $\mathcal{A}=\mathcal{A}_{1}\otimes\mathcal{A}_{2}$. The tensor product is called nontrivial if $\{e\}<G_{1}<G$ and $\{e\}<G_{2}<G$, and trivial otherwise. On the other hand, if we are given $S$-rings $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ over $G_{1}$ and $G_{2}$, respectively, then the above partition defines the $S$-ring $\mathcal{A}=\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ over $G=G_{1}\times G_{2}$.
Let $\mathcal{A}$ be an $S$-ring over a group $G$ and $L$ an $\mathcal{A}$-subgroup of $G$. We say that $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}$ if there exists an $\mathcal{A}$-subgroup $U$ such that $G=L\times U$ and $\mathcal{A}=\mathcal{A}_{L}\otimes\mathcal{A}_{U}$.
One can check that
(3) $\operatorname{Aut}(\mathcal{A}_{1}\otimes\mathcal{A}_{2})=\operatorname{Aut}(\mathcal{A}_{1})\otimes\operatorname{Aut}(\mathcal{A}_{2}).$
The lemma below immediately follows from Eq. (3).
Lemma 3.1.
A tensor product of two $S$-rings is schurian if and only if each of them is schurian.
If $\mathcal{A}_{G_{1}}=\operatorname{Cyc}(K_{1},G_{1})$ and $\mathcal{A}_{G_{2}}=\operatorname{Cyc}(K_{2},G_{2})$, then
(4) $\mathcal{A}_{G_{1}}\otimes\mathcal{A}_{G_{2}}=\operatorname{Cyc}(K_{1}\times K_{2},G_{1}\times G_{2}).$
Lemma 3.2.
[11, Lemma 2.3] Let $\mathcal{A}$ be an $S$-ring over an abelian group $G=G_{1}\times G_{2}$. Suppose that $G_{1}$ and $G_{2}$ are $\mathcal{A}$-subgroups. Then
  1. (1)
    $X_{G_{i}}\in\mathcal{S}(\mathcal{A})$ for all $X\in\mathcal{S}(\mathcal{A})$ and $i\in\{1,2\}$;
  2. (2)
    $\mathcal{A}\geq\mathcal{A}_{G_{1}}\otimes\mathcal{A}_{G_{2}}$, and the equality is attained whenever $\mathcal{A}_{G_{i}}=\mathbb{Z}G_{i}$ for some $i\in\{1,2\}$.
Lemma 3.3.
Let $K\leq\operatorname{Sym}(G)$ such that $K\geq G_{r}$ and $\mathcal{A}=V(K,G)$. Suppose that $\mathcal{A}=\mathcal{A}_{G_{1}}\otimes\mathcal{A}_{G_{2}}$ for some $\mathcal{A}$-subgroups $G_{1}$ and $G_{2}$ such that $G=G_{1}\times G_{2}$ and $K^{G_{i}}\in\mathcal{K}^{\min}(\mathcal{A}_{G_{i}})$ for some $i\in\{1,2\}$. Then $K=K^{G_{1}}\times K^{G_{2}}$.
Proof.
One can see that
$K\leq K^{G_{1}}\times K^{G_{2}}\leq\operatorname{Aut}(\mathcal{A}_{1})\times\operatorname{Aut}(\mathcal{A}_{2})\leq\operatorname{Aut}(\mathcal{A})$
and hence $K$ is a subdirect product of $K^{G_{1}}$ and $K^{G_{2}}$. This implies that there exist groups $K_{1}\trianglelefteq K^{G_{1}}$, $K_{2}\trianglelefteq K^{G_{2}}$, and an isomorphism $\psi$ from $K^{G_{1}}/K_{1}$ to $K^{G_{2}}/K_{2}$ such that
$K=\{(f_{1},f_{2})\in K^{G_{1}}\times K^{G_{2}}:~(K_{1}f_{1})^{\psi}=K_{2}f_{2}\}.$
Since $V(K^{G_{1}},G_{1})=\mathcal{A}_{G_{1}}$, $V(K^{G_{2}},G_{2})=\mathcal{A}_{G_{2}}$, and $V(K,G)=\mathcal{A}_{G_{1}}\otimes\mathcal{A}_{G_{2}}$, we conclude that $\mathcal{A}_{G_{1}}=V(K_{1},G_{1})$ and $\mathcal{A}_{G_{2}}=V(K_{2},G_{2})$. By the condition of the lemma, $K^{G_{i}}\in\mathcal{K}^{\min}(\mathcal{A}_{G_{i}})$ which yields that $K_{i}=K^{G_{i}}$. Thus, $K=K^{G_{1}}\times K^{G_{2}}$ as required. ∎
Example 2. The following example demonstrates that the condition $K^{G_{i}}\in\mathcal{K}^{\min}(\mathcal{A}_{G_{i}})$ for some $i\in\{1,2\}$ in Lemma 3.3 is essential. Let $\mathcal{A}_{G_{1}}=\mathcal{T}_{G_{1}}$ and $\mathcal{A}_{G_{2}}=\mathcal{T}_{G_{2}}$, where $|G_{i}|=p$ is an odd prime for each $i\in\{1,2\}$, and
$K=\operatorname{Alt}(G_{1})\times\operatorname{Alt}(G_{2})\cup(\operatorname{Sym}(G_{1})\setminus\operatorname{Alt}(G_{1}))\times(\operatorname{Sym}(G_{2})\setminus\operatorname{Alt}(G_{2})).$
Then $K\geq\operatorname{Alt}(G_{1})\times\operatorname{Alt}(G_{2})\geq(G_{1})_{r}\times(G_{2})_{r}=G_{r}$ because $p$ is an odd prime, $\mathcal{A}=V(K,G)$, $K^{G_{1}}=\operatorname{Sym}(G_{1})$, $K^{G_{2}}=\operatorname{Sym}(G_{2})$, and $K\neq K^{G_{1}}\times K^{G_{2}}$. In this case, $K^{G_{i}}\notin\mathcal{K}^{\min}(\mathcal{A}_{G_{i}})$ for each $i\in\{1,2\}$ because $\operatorname{Alt}(G_{i})<K^{G_{i}}$ and $\operatorname{Alt}(G_{i})\approx_{2}K^{G_{i}}$.
The corollary below immediately follows from Lemma 3.3 and Eq. (3).
Corollary 3.4.
A tensor product of two $2$-minimal $S$-rings is $2$-minimal.
3.2. Generalized wreath product
Let $S=U/L$ be an $\mathcal{A}$-section of $G$. The $S$-ring $\mathcal{A}$ is called the $S$-wreath product or generalized wreath product of $\mathcal{A}_{U}$ and $\mathcal{A}_{G/L}$ if $L\trianglelefteq G$ and every basic set $X$ of $\mathcal{A}$ outside $U$ is a union of some $L$-cosets or, equivalently, $L\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus U}$. In this case, we write $\mathcal{A}=\mathcal{A}_{U}\wr_{S}\mathcal{A}_{G/L}$. The $S$-wreath product is said to be nontrivial if $L\neq\{e\}$ and $U\neq G$, and trivial otherwise. The construction of a generalized wreath product of $S$-rings was introduced in paper [5]. If $L=U$, then the $S$-wreath product coincides with the wreath product $\mathcal{A}_{L}\wr\mathcal{A}_{G/L}$ of $\mathcal{A}_{L}$ and $\mathcal{A}_{G/L}$. The $S$-ring $\mathcal{A}$ is said to be decomposable if $\mathcal{A}$ is a nontrivial $S$-wreath product for some $\mathcal{A}$-section $S$ and indecomposable otherwise.
Given a section $S=U/L$ of a group $G$ such that $L\trianglelefteq G$ and $S$-rings $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ over $U$ and $G/L$, respectively, such that $S$ is both an $\mathcal{A}_{1}$- and an $\mathcal{A}_{2}$-section, and $(\mathcal{A}_{1})_{S}=(\mathcal{A}_{2})_{S}$, there is a unique $S$-ring $\mathcal{A}$ over $G$ that is the $S$-wreath product with $\mathcal{A}_{U}=\mathcal{A}_{1}$ and $\mathcal{A}_{G/L}=\mathcal{A}_{2}$ (see [5]).
Lemma 3.5.
[7, Corollary 5.7] Let $\mathcal{A}$ be an $S$-ring over an abelian group $G$ and $S=U/L$ an $\mathcal{A}$-section. Suppose that $\mathcal{A}$ is the $S$-wreath product and $\mathcal{A}_{U}$ and $\mathcal{A}_{G/L}$ are schurian. Then $\mathcal{A}$ is schurian if and only if there exist two groups $K_{1}\leq\operatorname{Sym}(U)$ and $K_{0}\leq\operatorname{Sym}(G/L)$ such that
$K_{1}\geq U_{r},~K_{0}\geq(G/L)_{r},~K_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{U}),~K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{G/L}),~\text{and}~K_{0}^{S}=K_{1}^{S}.$
Lemma 3.6.
[23, Corollary 10.3] Under the hypothesis of Lemma 3.5, the $S$-ring $\mathcal{A}$ is schurian if $\mathcal{A}_{S}$ is $2$-minimal. In particular, $\mathcal{A}$ is schurian if $U=L$.
The lemma below for cyclic groups is [10, Theorem 7.5]. In fact, the proof for abelian groups is the same. To make the text self-contained, we provide a proof here.
Lemma 3.7.
Under the hypothesis of Lemma 3.5, the $S$-ring $\mathcal{A}$ is schurian whenever $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{U}$ or $\mathcal{A}_{S}$ is $\otimes$-complemented in $\mathcal{A}_{G/L}$.
Proof.
Suppose that the first condition of the lemma holds, i.e. $\mathcal{A}_{U}=\mathcal{A}_{L}\otimes\mathcal{A}_{H}$ for some $\mathcal{A}_{U}$-subgroup $H$. Clearly, $\mathcal{A}_{S}\cong\mathcal{A}_{H}$. Eq. (3) implies that $\operatorname{Aut}(\mathcal{A}_{U})=\operatorname{Aut}(\mathcal{A}_{L})\times\operatorname{Aut}(\mathcal{A}_{H})$. Let $K_{0}=\operatorname{Aut}(\mathcal{A}_{G/L})$. Observe that $K_{0}^{S}\approx_{2}\operatorname{Aut}(\mathcal{A}_{S})$ because $\mathcal{A}_{G/L}$ is schurian. Since $\mathcal{A}_{S}\cong\mathcal{A}_{H}$, there is $K_{2}\leq\operatorname{Sym}(H)$ such that $K_{2}\geq H_{r}$ and $K_{2}^{S}=K_{0}^{S}$. Put $K_{1}=\operatorname{Aut}(\mathcal{A}_{L})\times K_{2}$. One can see that:
  1. (1)
    $K_{1}\geq U_{r}$ because $K_{2}\geq H_{r}$;
  2. (2)
    $K_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{U})$ because $K_{2}^{S}\approx_{2}\operatorname{Aut}(\mathcal{A}_{S})$ and $\operatorname{Aut}(\mathcal{A}_{U})=\operatorname{Aut}(\mathcal{A}_{L})\times\operatorname{Aut}(\mathcal{A}_{H})$;
  3. (3)
    $K_{1}^{S}=K_{2}^{S}=K_{0}^{S}$.
Therefore all the conditions of Lemma 3.5 hold for $K_{1}$ and $K_{0}$. Thus, $\mathcal{A}$ is schurian.
Now suppose that the second condition of the lemma holds, i.e. $\mathcal{A}_{G/L}=\mathcal{A}_{S}\otimes\mathcal{A}_{H}$ for some $\mathcal{A}_{G/L}$-subgroup $H$. Eq. (3) yields that $\operatorname{Aut}(\mathcal{A}_{G/L})=\operatorname{Aut}(\mathcal{A}_{S})\times\operatorname{Aut}(\mathcal{A}_{H})$. Let $K_{1}=\operatorname{Aut}(\mathcal{A}_{U})$. Note that $K_{1}^{S}\approx_{2}\operatorname{Aut}(\mathcal{A}_{S})$ because $\mathcal{A}_{U}$ is schurian and $K_{1}^{S}\geq S_{r}$ because $K_{1}\geq U_{r}$. Put $K_{0}=K_{1}^{S}\times\operatorname{Aut}(\mathcal{A}_{H})$. Then:
  1. (1)
    $K_{0}\geq(G/L)_{r}$ because $K_{1}^{S}\geq S_{r}$;
  2. (2)
    $K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{G/L})$ because $K_{1}^{S}\approx_{2}\operatorname{Aut}(\mathcal{A}_{S})$ and $\operatorname{Aut}(\mathcal{A}_{G/L})=\operatorname{Aut}(\mathcal{A}_{S})\times\operatorname{Aut}(\mathcal{A}_{H})$;
  3. (3)
    $K_{0}^{S}=K_{1}^{S}$.
Thus, all the conditions of Lemma 3.5 hold for $K_{1}$ and $K_{0}$ and consequently $\mathcal{A}$ is schurian. ∎
3.3. Star product
Let $L$ and $U$ be normal $\mathcal{A}$-subgroups. The $S$-ring $\mathcal{A}$ is called the star product of $\mathcal{A}_{L}$ and $\mathcal{A}_{U}$ if the following conditions hold:
$(1)$ each $X\in\mathcal{S}(\mathcal{A})$ with $X\subseteq(U\setminus L)$ is a union of some $(L\cap U)$-cosets;
$(2)$ for each $X\in\mathcal{S}(\mathcal{A})$ with $X\subseteq G\setminus(L\cup U)$ there exist $Y\in\mathcal{S}(\mathcal{A}_{L})$ and $Z\in\mathcal{S}(\mathcal{A}_{U})$ such that $X=YZ$.
In this case, we write $\mathcal{A}=\mathcal{A}_{L}\star\mathcal{A}_{U}$. The construction of a star product of $S$-rings was introduced in paper [13] and extended to association schemes under the name “crested product” in paper [1]. A star product is called nontrivial if $L\neq\{e\}$ and $U\neq G$ and trivial otherwise.
Let $L$ and $U$ be proper nontrivial normal $\mathcal{A}$-subgroups of $G$ and $\mathcal{A}=\mathcal{A}_{L}\star\mathcal{A}_{U}$. It follows immediately from the definitions that:
  1. (1)
    if $|L\cap U|=1$, then $\mathcal{A}=\mathcal{A}_{L}\otimes\mathcal{A}_{U}$;
  2. (2)
    if $|L\cap U|>1$, then $\mathcal{A}$ is the nontrivial $L/(L\cap U)$-wreath product.
4. Duality
Let $G$ be an abelian group. Denote by $\widehat{G}$ the group dual to $G$, i.e., the group of all irreducible complex characters of $G$. It is well known that $\widehat{G}\cong G$ and there is a uniquely determined lattice antiisomorphism between the subgroups of $G$ and $\widehat{G}$. The image of a subgroup $H$ of $G$ with respect to this antiisomorphism is denoted by $H^{\bot}$.
For any $S$-ring $\mathcal{A}$ over the group $G$, one can define the dual $S$-ring $\widehat{\mathcal{A}}$ over $\widehat{G}$ as follows: two irreducible characters of $G$ belong to the same basic set of $\widehat{\mathcal{A}}$ if they have the same value on each basic set of $\mathcal{A}$ (for the exact definition, we refer the reader to [8, 33]). One can verify that
$\operatorname{rk}(\widehat{\mathcal{A}})=\operatorname{rk}(\mathcal{A})$
and the $S$-ring dual to $\widehat{\mathcal{A}}$ is equal to $\mathcal{A}$. The following lemma collects some facts on the dual $S$-ring taken from [9, Section 2.3].
Lemma 4.1.
Let $G$ be an abelian group and $\mathcal{A}$ an $S$-ring over $G$. Then
  1. (1)
[…]
  2. (2)
    $\widehat{\mathcal{A}_{H}}=\widehat{\mathcal{A}}_{\widehat{G}/H^{\bot}}$ and $\widehat{\mathcal{A}_{G/H}}=\widehat{\mathcal{A}}_{H^{\bot}}$ for every $H\in\mathcal{H}(\mathcal{A})$;
  3. (3)
    $\mathcal{A}=\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ if and only if $\widehat{\mathcal{A}}=\widehat{\mathcal{A}_{1}}\otimes\widehat{\mathcal{A}_{2}}$;
  4. (4)
    $\mathcal{A}$ is the $U/L$-wreath product for some $\mathcal{A}$-section $U/L$ if and only if $\widehat{\mathcal{A}}$ is the $L^{\bot}/U^{\bot}$-wreath product.
  5. (5)
    $\mathcal{A}$ is cyclotomic if and only if so is $\widehat{\mathcal{A}}$.
5. $S$-rings over $H\times C_{p}$
Throughout this section, $G=H\times P$, where $H$ is an abelian group and $P\cong C_{p}$ with prime $p$ coprime to $|H|$, and $\mathcal{A}$ is an $S$-ring over $G$. Let $H_{1}$ be a maximal $\mathcal{A}$-subgroup contained in $H$ and $P_{1}$ the least $\mathcal{A}$-subgroup containing $P$. Note that $H_{1}P_{1}$ is an $\mathcal{A}$-subgroup.
Lemma 5.1.
[19, Lemma 6.3] If $H_{1}\neq(H_{1}P_{1})_{p^{\prime}}$, the Hall $p^{\prime}$-subgroup of $H_{1}P_{1}$, then $\mathcal{A}_{H_{1}P_{1}}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$.
Lemma 5.2.
[24, Proposition 15] If $\mathcal{A}_{(H_{1}P_{1})/H_{1}}\cong\mathbb{Z}C_{p}$, then $\mathcal{A}_{H_{1}P_{1}}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$.
Lemma 5.3.
Suppose that $H_{1}<H$. Then one of the following statements holds:
  1. (1)
    $H_{1}P_{1}=G$, $P_{1}\lneq G$, and $\mathcal{A}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$;
  2. (2)
    $\mathcal{A}=\mathcal{A}_{H_{1}}\wr\mathcal{A}_{G/H_{1}}$ with $\mathcal{A}_{G/H_{1}}=\mathcal{T}_{G/H_{1}}$;
  3. (3)
    $\mathcal{A}$ is the nontrivial $(H_{1}P_{1})/P_{1}$-wreath product.
Proof.
From [11, Lemma 6.2] it follows that Statement $(2)$ of the lemma holds or $\mathcal{A}$ is the $(H_{1}P_{1})/P_{1}$-wreath product with $P_{1}<G$. If the above wreath product is nontrivial, then Statement $(3)$ of the lemma holds, whereas if it is trivial, i.e. if $H_{1}P_{1}=G$, then $H_{1}\neq(H_{1}P_{1})_{p^{\prime}}=G_{p^{\prime}}=H$. So $\mathcal{A}=\mathcal{A}_{H_{1}P_{1}}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$ by Lemma 5.1 and Statement $(1)$ of the lemma holds. ∎
Lemma 5.4.
If $p=2$, then $\mathcal{A}_{H_{1}P_{1}}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$.
Proof.
If $H_{1}\neq(H_{1}P_{1})_{p^{\prime}}$, then the lemma follows from Lemma 5.1. If $H_{1}=(H_{1}P_{1})_{p^{\prime}}$, then $(H_{1}P_{1})/H_{1}\cong C_{2}$ and hence $\mathcal{A}_{(H_{1}P_{1})/H_{1}}\cong\mathbb{Z}C_{2}$. So in this case, the lemma follows from Lemma 5.2. ∎
6. $S$-rings over cyclic groups
In this section, we provide several results on $S$-rings over cyclic groups, required for the proofs of the main results. An $S$-ring over a cyclic group is said to be circulant.
6.1. General results
Lemma 6.1.
[34, Theorem 25.4] If $G$ is an abelian group of composite order having a cyclic Sylow subgroup, then $\mathcal{T}_{G}$ is the only primitive $S$-ring over $G$.
The description of $S$-rings over a cyclic group was obtained in papers [20, 21]. Below, we provide this description in a convenient for us form taken from [7, Theorem 4.1]. If $\mathcal{A}$ is an $S$-ring over a cyclic group $G$, then put $\operatorname{rad}(\mathcal{A})=\operatorname{rad}(X)$, where $X$ is a basic set of $\mathcal{A}$ containing a generator of $G$. Due to Lemma 2.3, all basic sets of $\mathcal{A}$, containing a generator of $G$, are rationally conjugate and hence have the same radical. Therefore $\operatorname{rad}(\mathcal{A})$ does not depend on the choice of $X$.
Lemma 6.2.
Let $\mathcal{A}$ be an $S$-ring over a cyclic group. Then the following statements hold:
  1. (1)
    $|\operatorname{rad}(\mathcal{A})|=1$ if and only if $\mathcal{A}$ is a tensor product (possibly, trivial) of a normal cyclotomic $S$-ring with trivial radical and trivial $S$-rings;
  2. (2)
    $|\operatorname{rad}(\mathcal{A})|>1$ if and only if $\mathcal{A}$ is a nontrivial generalized wreath product.
Lemma 6.3.
[7, Theorem 4.2(1)] Every normal $S$-ring over a cyclic group is cyclotomic and, in particular, schurian.
Lemma 6.4.
Every cyclotomic $S$-ring over a cyclic group is Cayley minimal.
Proof.
Let $G$ be a cyclic group and $\mathcal{A}=\operatorname{Cyc}(K,G)$ for some $K\leq\operatorname{Aut}(G)$. Assume that there exists $K^{\prime}<K$ such that $\mathcal{A}=\operatorname{Cyc}(K^{\prime},G)$. Let $X$ be a basic set of $\mathcal{A}$ containing a generator $x$ of $G$. Clearly, the one-point stabilizer of $x$ in $\operatorname{Aut}(G)$ is trivial. Since $X\in\operatorname{Orb}(K,G)\cap\operatorname{Orb}(K^{\prime},G)$, we conclude that $|K|=|K_{x}||X|=|X|=|K^{\prime}_{x}||X|=|K^{\prime}|$, a contradiction to $K^{\prime}<K$. ∎
Lemma 6.5.
Every normal $S$-ring over a cyclic group is $2$-minimal.
Proof.
Let $\mathcal{A}$ be a normal $S$-ring over a cyclic group $G$. Clearly, $\operatorname{Aut}(\mathcal{A})\leq\operatorname{Hol}(G)$. As $\mathcal{A}$ is normal, $\mathcal{A}$ is schurian and hence $\mathcal{A}=\operatorname{Cyc}(\operatorname{Aut}(\mathcal{A})_{e},G)$ by Lemma 6.3. Assume that $\mathcal{A}$ is not $2$-minimal, i.e. there exists $K<\operatorname{Aut}(\mathcal{A})$ such that $K\geq G_{r}$ and $K\approx_{2}\operatorname{Aut}(\mathcal{A})$. Since $K<\operatorname{Aut}(\mathcal{A})\leq\operatorname{Hol}(G)$, we obtain $K_{e}\leq\operatorname{Aut}(G)$. This implies that $\mathcal{A}=\operatorname{Cyc}(K_{e},G)$. From Lemma 6.4 it follows that $K_{e}=\operatorname{Aut}(\mathcal{A})_{e}$ and consequently $K=G_{r}K_{e}=G_{r}\operatorname{Aut}(\mathcal{A})_{e}=\operatorname{Aut}(\mathcal{A})$, a contradiction to the assumption. ∎
Lemma 6.6.
Under the hypothesis of Lemma 3.5, suppose that $U$ and $G/L$ are cyclic. Then $\mathcal{A}$ is schurian whenever $|S|$ is prime.
Proof.
If $\mathcal{A}_{S}\neq\mathcal{T}_{S}$, then $\mathcal{A}_{S}$ is normal by Lemma 6.2. So $\mathcal{A}_{S}$ is $2$-minimal by Lemma 6.5 and $\mathcal{A}$ is schurian by Lemma 3.6.
Suppose that $\mathcal{A}_{S}=\mathcal{T}_{S}$. Then $\mathcal{A}_{S}=\operatorname{Cyc}(\operatorname{Aut}(S),S)=V(\operatorname{Hol}(S),S)$. Since $\operatorname{Aut}(\mathcal{A}_{U})$ and $\mathcal{A}_{G/L}$ are schurian, [7, Theorem 8.1(2)] implies that there exist groups $K_{1}\leq\operatorname{Aut}(\mathcal{A}_{U})$ and $K_{0}\leq\operatorname{Aut}(\mathcal{A}_{G/L})$ such that
$K_{1}\geq U_{r},~K_{0}\geq(G/L)_{r},~K_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{U}),~K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{G/L}),~K_{1}^{S}=K_{0}^{S}=\operatorname{Hol}(S).$
Thus, all the conditions of Lemma 3.5 hold for $K_{1}$ and $K_{0}$ and hence $\mathcal{A}$ is schurian. ∎
6.2. $S$-rings over cyclic $p$-groups
For the lemma below see, e.g., [25, Lemma 2.4].
Lemma 6.7.
Every $S$-ring over a cyclic group of prime order is cyclotomic.
The first statement of the lemma below immediately follows from [6, Lemma 5.1], whereas the second one follows from Lemma 6.2 and the first one.
Lemma 6.8.
Let $p$ be an odd prime, $G$ a cyclic $p$-group, $\mathcal{A}$ an $S$-ring over $G$, and $X\in\mathcal{S}(\mathcal{A})\setminus\{\{e\}\}$.
  1. (1)
    If $\mathcal{A}=\operatorname{Cyc}(K,G)$ for some $K\leq\operatorname{Aut}(G)$ and $|\operatorname{rad}(\mathcal{A})|=1$, then $|X|=|K|\leq p-1$ and all the elements of $X$ lie in pairwise distinct cosets by every proper subgroup of $\langle X\rangle$.
  2. (2)
    If $\mathcal{A}$ is nontrivial with trivial radical, then $\mathcal{A}=\operatorname{Cyc}(K,G)$ for some $K\leq\operatorname{Aut}(G)$ such that $|K|$ divides p1p-1 and $|X|=|K|$.
Lemma 6.9.
Let $p$ be an odd prime and $\mathcal{A}$ a cyclotomic $S$-ring over a cyclic $p$-group $G$ such that $|\operatorname{rad}(\mathcal{A})|\in\{1,p\}$. Then $|\operatorname{rad}(\mathcal{A}_{S})|=1$ for every $\mathcal{A}$-section $S\neq G$.
Proof.
Let $\mathcal{A}=\operatorname{Cyc}(K,G)$, where $K\leq\operatorname{Aut}(G)$. Since $G$ is a cyclic $p$-group of odd order, $\operatorname{Aut}(G)=A_{1}\times A_{2}$, where
$A_{1}\cong C_{p-1}~\text{and}~A_{2}=\{\sigma_{m}:~m\in\mathbb{Z},~m\equiv 1\mod~p\}\cong C_{|G|/p}.$
So $K=K_{1}\times K_{2}$, where $K_{1}=K\cap A_{1}$ and $K_{2}=K\cap A_{2}$. One can see that
$|K_{2}|=|\operatorname{rad}(\mathcal{A})|\in\{1,p\}.$
Therefore $|K_{2}|=1$ or $K_{2}=\{\sigma_{m}:~m\in\mathbb{Z},~m\equiv 1\mod~|G|/p\}$. This implies that $K^{G/L_{0}}$ and $K^{U_{0}}$, where $L_{0}$ and $U_{0}$ are the subgroups of $G$ of order and index $p$, respectively, are $p^{\prime}$-groups of order at most p1p-1. If $S=U/L$ is an $\mathcal{A}$-section and $S\neq G$, then $U\leq U_{0}$ or $L\geq L_{0}$. This yields that $\mathcal{A}_{S}=\operatorname{Cyc}((K^{G/L_{0}})^{S},S)$ or $\mathcal{A}_{S}=\operatorname{Cyc}((K^{U_{0}})^{S},S)$. As $|(K^{G/L_{0}})^{S}|\leq p-1$ and $|(K^{U_{0}})^{S}|\leq p-1$, we conclude that every basic set of $\mathcal{A}_{S}$ is of size at most p1p-1 and hence has a trivial radical. Thus, $|\operatorname{rad}(\mathcal{A}_{S})|=1$. ∎
Lemma 6.10.
Let $\mathcal{A}$ be a nontrivial generalized wreath product over a cyclic $p$-group $G$. Then $\mathcal{A}$ is the nontrivial $U/L$-wreath product for an $\mathcal{A}$-section $U/L$ of $G$ such that $L$ is the least nontrivial $\mathcal{A}$-subgroup and $|\operatorname{rad}(\mathcal{A}_{U})|=1$.
Proof.
Since $G$ is a cyclic $p$-group, there exists the least nontrivial $\mathcal{A}$-subgroup $L$. Let $U$ be the $\mathcal{A}$-subgroup generated by all basic set of $\mathcal{A}$ with trivial radical. Note that $U<G$ because $\mathcal{A}$ is a nontrivial generalized wreath product. Every basic set $X$ of $\mathcal{A}$ outside $U$ has a nontrivial radical and hence $\operatorname{rad}(X)\geq L$. Thus, $\mathcal{A}$ is the nontrivial $U/L$-wreath product as desired. ∎
A subset of $G$ is said to be regular if it consists of elements of the same order. The lemma below can be deduced from the description of $S$-rings over cyclic $p$-groups [27]. To make the text self-contained, we provide the proof here.
Lemma 6.11.
Let $\mathcal{A}$ be an $S$-ring over a cyclic $p$-group $G$ and $X\in\mathcal{S}(\mathcal{A})$. Suppose that $X$ is not regular. Then $X=U\setminus L$ for some $\mathcal{A}$-subgroups $U>L$ such that $|U:L|\geq p^{2}$ and $\mathcal{A}=\mathcal{A}_{L}\wr\mathcal{A}_{G/L}=\mathcal{A}_{U}\wr\mathcal{A}_{G/U}$.
Proof.
Let $L=\operatorname{rad}(X)$. The image of $T\subseteq G$ under the canonical epimorphism from $G$ to $G/L$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muT\mkern-1.5mu}\mkern 1.5mu$. The $S$-ring $\mathcal{A}_{\langle\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\rangle}$ has a trivial radical. So $\mathcal{A}_{\langle\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\rangle}$ is trivial or cyclotomic by Lemma 6.2. As $X$ is not regular, $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu$ also is not regular. Therefore $\mathcal{A}_{\langle\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\rangle}$ can not be cyclotomic and hence it is trivial. Thus, $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\cup\{L\}$ is an $\mathcal{A}_{G/L}$-subgroup. Since $L=\operatorname{rad}(X)$, we conclude that $X=U\setminus L$ for some $\mathcal{A}$-subgroup $U$ as required. One can see that $|U:L|\geq p^{2}$ because otherwise $X$ is regular which contradicts to the assumption of the lemma.
Now let us prove the second part of the lemma. If $U=G$, then $\mathcal{A}$ is the $L/L$-wreath product (possibly, trivial) and the trivial $U/U$-wreath product. Further, we assume that $U<G$ and hence $\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu<\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu$. The latter implies that $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ is nontrivial. Note that $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu=(\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu)^{\#}$ is a nonregular basic set of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ with trivial radical. So $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ can not be cyclotomic and hence $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ is a nontrivial generalized wreath product by Lemma 6.2. In view of Lemma 6.10, $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$ is the nontrivial $W/V$-wreath product, where $\operatorname{rad}(\mathcal{A}_{W})$ is trivial and $V$ is the least nontrivial $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}$-subgroup. Since $\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)$ is trivial, we conclude that $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\subseteq W$. Hence $\mathcal{A}_{W}$ is not cyclotomic. Due to Lemma 6.2, the $S$-ring $\mathcal{A}_{W}$ is trivial. Therefore $W=V=\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu$. Thus,
$\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu}=\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}\wr\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu/\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}.$
In view of the above equality, to prove the lemma it suffices to show that $L\leq\operatorname{rad}(Y)$ for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus U}$. Assume the contrary that $N=\operatorname{rad}(Y)<L$ for some $Y\in\mathcal{S}(\mathcal{A})_{G\setminus U}$. The image of $T\subseteq G$ under the canonical epimorphism from $G$ to $G/N$ is denoted by $\widetilde{T}$. As $Y$ lies outside $U$, we have $X\subseteq U\leq\langle Y\rangle$. So $\widetilde{X}$ is a nonregular basic set of $\mathcal{A}_{\langle\widetilde{Y}\rangle}$. This implies that $\mathcal{A}_{\langle\widetilde{Y}\rangle}$ can not be cyclotomic. Therefore $\mathcal{A}_{\langle\widetilde{Y}\rangle}$ is trivial by Lemma 6.2. On the other hand, $\widetilde{X}$ and $\widetilde{Y}$ are nontrivial basic sets of $\mathcal{A}_{\langle\widetilde{Y}\rangle}$, a contradiction. ∎
6.3. $S$-rings over some other cyclic groups
Lemma 6.12.
Every nontrivial $S$-ring over a cyclic group of order a product of two distinct primes is cyclotomic or a nontrivial wreath product.
Proof.
An $S$-ring over a group of prime order is cyclotomic by Lemma 6.7 and hence a tensor product of $S$-rings over groups of prime orders is cyclotomic by Eq. (4). Therefore the required statement follows from Lemma 6.2. ∎
Lemma 6.13.
Let $\mathcal{A}$ be a nontrivial $S$-ring over a cyclic group $G$ of order $2p^{k}$, where $p$ is an odd prime and $k\geq 1$. Suppose that $|\operatorname{rad}(\mathcal{A})|=1$. Then $\mathcal{A}$ is normal unless $\mathcal{A}=\mathbb{Z}L\otimes\mathcal{T}_{U}$, where $L,U\leq G$ are such that $|L|=2$ and $|U|=p^{k}$.
Proof.
Assume that $\mathcal{A}$ is nonnormal. Then $\mathcal{A}=\mathbb{Z}_{L}\otimes\mathcal{A}_{U}$ by Lemma 6.2. If $|\operatorname{rad}(\mathcal{A}_{U})|>1$, then $|\operatorname{rad}(\mathcal{A})|>1$, a contradiction to the assumption of the lemma. So $|\operatorname{rad}(\mathcal{A}_{U})|=1$. If $\mathcal{A}_{U}\neq\mathcal{T}_{U}$, then $\mathcal{A}_{U}$ is normal by Lemma 6.2 and consequently so is $\mathcal{A}$ by Eq. (3), a contradiction to the assumption of the lemma. Thus, $\mathcal{A}_{U}=\mathcal{T}_{U}$ and hence $\mathcal{A}=\mathbb{Z}_{L}\otimes\mathcal{T}_{U}$. ∎
7. $S$-rings over $E_{4}\times C_{n}$
The main goal of this section is to give a description of $S$-rings over the group $E_{4}\times C_{n}$, where $n\in\{p^{k},pq\}$, $p$ and $q$ are odd primes, and $k\geq 1$. We say that an $S$-ring $\mathcal{A}$ over $E_{4}\times C_{n}$ is dense if $E_{4}$ and $C_{n}$ are $\mathcal{A}$-subgroups and nondense otherwise.
Theorem 7.1.
Let $n\in\{p^{k},pq\}$, where $p$ and $q$ are odd primes and $k$ is a positive integer, and $\mathcal{A}$ a nontrivial $S$-ring over the group $G=H\times D$, where $H\cong E_{4}$ and $D\cong C_{n}$. Then $\mathcal{A}$ is cyclotomic, or a nontrivial tensor product, or a nontrivial $S$-wreath product for some $\mathcal{A}$-section $S=U/L$ and one of the following statements holds:
  1. (1)
    $|S|\leq 2$;
  2. (2)
    $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{U}$ or $\mathcal{A}_{S}$ is $\otimes$-complemented in $\mathcal{A}_{G/L}$;
  3. (3)
    $\mathcal{A}$ is nondense and at least one of the $S$-rings $\mathcal{A}_{U}$, $\mathcal{A}_{G/L}$ is a circulant cyclotomic $S$-ring with trivial radical;
  4. (4)
    $\mathcal{A}$ is dense, $L\leq D$, $U\geq H$, $\mathcal{A}_{U}$ is cyclotomic, and $|\operatorname{rad}(\mathcal{A}_{U\cap D})|=1$ unless $n=3^{k}$ and $|\operatorname{rad}(\mathcal{A}_{U\cap D})|=3$.
Observe that $U\cap D$ in Statement $(4)$ of Theorem 7.1 is cyclic and hence $\operatorname{rad}(\mathcal{A}_{U\cap D})$ is defined correctly.
Clearly, Theorem 1.3 immediately follows from Theorem 7.1.
We divide the proof of Theorem 7.1 into two cases depending on whether $\mathcal{A}$ is nondense or dense. We prove that every nondense nontrivial $S$-ring over $G$ is a nontrivial tensor or generalized wreath product and every dense $S$-ring over $G$ is cyclotomic or a nontrivial generalized wreath product. The nondense case is divided into two subcases depending on whether $H$ or $D$ is not an $\mathcal{A}$-subgroup (Theorems 7.3 and 7.8, respectively), whereas the dense case is divided into two subcases depending on whether $\operatorname{rad}(\mathcal{A}_{D})$ is trivial or not (Theorems 7.12 and 7.13, respectively).
In the first subsection, we study a structure of a basic set of $\mathcal{A}$. The second and third subsections deal with the above mentioned cases when $\mathcal{A}$ is nondense and when $\mathcal{A}$ is dense, respectively. In the fourth subsection, we give a proof of Theorem 7.1. The fifth subsection contains some auxiliary statements to be used in the proof of Theorem 1.2.
Throughout this section, $p$ and $q$ are odd primes, $k$ is a positive integer, $n\in\{p^{k},pq\}$, $D\cong C_{n}$, $P=D_{p}=\{g\in D:~g^{p}=e\}$, $H\cong E_{4}$, $G=H\times D$, and $\mathcal{A}$ is an $S$-ring over $G$.
7.1. Structure of a basic set
Lemma 7.2.
Let $n=p^{k}$ and $X\in\mathcal{S}(\mathcal{A})$. Suppose that $X\nsubseteq D$. Then one of the following statements holds:
  1. (1)
    $P\leq\operatorname{rad}(X)$;
  2. (2)
    $\langle(X^{[p]})_{H}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $H$;
  3. (3)
    $X\cap D\neq\varnothing$ and $X\cup\{e\}$ is an $\mathcal{A}$-subgroup of $G$.
Proof.
Assume that $P\nleq\operatorname{rad}(X)$, i.e. Statement $(1)$ of the lemma does not hold. Then the union $Y$ of all $hX_{h,l}=X\cap hD_{l}^{*}$, $h\in H$, $l\in\operatorname{Div}(n)$, such that $X_{h,l}\neq\varnothing$ and $P\nleq\operatorname{rad}(X_{h,l})$ is nonempty. As $D\cong C_{p^{k}}$ and $X_{h,l}\subseteq D$, the condition $P\nleq\operatorname{rad}(X_{h,l})$ is equivalent to $|\operatorname{rad}(X_{h,l})|=1$. Lemma 2.5 implies that each $X_{h,l}$ is an orbit of some subgroup $K$ of $\operatorname{Aut}(D_{l})$. From Lemma 6.8(1) it follows that $|K|\leq p-1$ and hence
(5) $|X_{h,l}|\leq p-1$
for all $h\in H$ and $l\in\operatorname{Div}(n)$ such that $X_{h,l}\neq\varnothing$ and $P\nleq\operatorname{rad}(X_{h,l})$.
Put $H_{0}=Y_{H}$, $Y^{[p]^{0}}=Y$, $Y^{[p]^{1}}=Y^{[p]}$, and $Y^{[p]^{i}}=(Y^{[p]^{i-1}})^{[p]}$ for $i\geq 2$. Due to the definition of $Y$ and Eq. (5), we have $Y^{[p]}=X^{[p]}$ and $Y^{[p]^{i}}=\{y^{p}:~y\in Y^{[p]^{i-1}}\}$. Therefore $(Y^{[p]^{i}})_{H}=H_{0}$ for every $i\geq 1$, in particular, $(Y^{[p]})_{H}=(X^{[p]})_{H}=H_{0}$. By Lemma 2.4, the set $Y^{[p]^{i}}$ is an $\mathcal{A}$-set for every $i\geq 1$. So $Y^{[p]^{k}}=H_{0}$ is an $\mathcal{A}$-set. Thus, $\langle H_{0}\rangle$ is an $\mathcal{A}$-subgroup. If $H_{0}\neq\{e\}$, then Statement $(2)$ of the lemma holds.
In view of the previous paragraph, we may assume that $H_{0}=\{e\}$ and hence
$Y\subseteq X\cap D\neq\varnothing.$
Let $l=p^{j}$ be the greatest integer such that $X_{e,l}\neq\varnothing$ and $P\nleq\operatorname{rad}(X_{e,l})$. By the supposition of the lemma, $X\setminus D\neq\varnothing$. So
(6) $X\setminus D_{l}\neq\varnothing~\text{and}~P\leq\operatorname{rad}(X\setminus D_{l}).$
If $j\neq 1$, then $Y^{[p]^{j-1}}$ is a nonemepty $\mathcal{A}$-set inside $P^{\#}$. Therefore $P=\langle Y^{[p]^{j-1}}\rangle$ is an $\mathcal{A}$-subgroup. Now Lemma 2.1 and Eq. (6) imply that $P\leq\operatorname{rad}(X)$ in contrast to the assumption $P\nleq\operatorname{rad}(X)$. If $j=1$, then $X_{e,l}\subseteq P^{\#}$. Eq. (6) and Lemma 2.2 applied to $X$ and $P$ yield that $\operatorname{rad}(X)\leq P$ and $X=\langle X\rangle\setminus\operatorname{rad}(X)$. By the assumption, $P\nleq\operatorname{rad}(X)$ and consequently $\operatorname{rad}(X)$ is trivial. Thus, $X=\langle X\rangle\setminus\{e\}$ and Statement $(3)$ of the lemma holds. ∎
7.2. Nondense $S$-rings
Theorem 7.3.
Suppose that $H\cong E_{4}$ is not an $\mathcal{A}$-subgroup. Then $\mathcal{A}$ is trivial, or a nontrivial tensor product, or a nontrivial $U/L$-wreath product for some $\mathcal{A}$-section $U/L$ satisfying one of the following conditions:
  1. (1)
    $|U/L|\leq 2$;
  2. (2)
    $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{U}$;
  3. (3)
    $4\nmid|U|$ and $|\operatorname{rad}(\mathcal{A}_{U})|=1$.
We divide the proof of Theorem 7.3 into two propositions.
Proposition 7.4.
If $H$ is not an $\mathcal{A}$-subgroup, then Theorem 7.3 holds unless $\mathcal{A}$ is a nontrivial $U/L$-wreath product for some $\mathcal{A}$-section $U/L$ such that $4\nmid|U|$ and $|L|\neq 2$.
Proof.
Let $n=pq$. Observe that the subgroup of $G$ of order $4p$ or the subgroup of $G$ of order $4q$ is not an $\mathcal{A}$-subgroup because otherwise $H$ is an intersection of $\mathcal{A}$-subgroups which contradicts to the supposition of the proposition. Without loss of generality, we may assume that the subgroup of $G$ of order $4q$ is not an $\mathcal{A}$-subgroup. Then
(7) $|H_{1}|\in\{1,2,q,2q\},$
where $H_{1}$ is a maximal $\mathcal{A}$-subgroup whose order is not divisible by $p$.
Let $P_{1}$ be the least $\mathcal{A}$-subgroup containing $P\cong C_{p}$. One of the statements of Lemma 5.3 holds for $\mathcal{A}$, $H_{1}$, and $P_{1}$. If Statement $(1)$ of Lemma 5.3 holds, then $H_{1}P_{1}=G$ and $\mathcal{A}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$. If $H_{1}\cap P_{1}$ is trivial, then $\mathcal{A}=\mathcal{A}_{H_{1}}\otimes\mathcal{A}_{P_{1}}$ is a nontrivial tensor product and we are done. Otherwise, $\mathcal{A}$ is the nontrivial $U/L=H_{1}/(H_{1}\cap P_{1})$-wreath product. Together with Eq. (7), this implies that $|U/L|\leq 2$ as desired or $|H_{1}|=2q$ and $|H_{1}\cap P_{1}|=2$. In the latter case, Lemma 6.12 yields that $\mathcal{A}_{H_{1}}=\mathcal{A}_{H_{1}\cap P_{1}}\otimes\mathcal{A}_{Q}$, where $Q$ is an $\mathcal{A}$-subgroup of order $q$, or $\mathcal{A}_{H_{1}}=\mathcal{A}_{H_{1}\cap P_{1}}\wr\mathcal{A}_{H_{1}/(H_{1}\cap P_{1})}$. In the first case, $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{U}$ and $U/L$ satisfies Condition $(2)$ from Theorem 7.3 as required. In the second one, $\mathcal{A}$ is the $L/L=(H_{1}\cap P_{1})/(H_{1}\cap P_{1})$-wreath product and the section $L/L$ satisfies Condition $(1)$ from Theorem 7.3 as required.
If Statement $(2)$ of Lemma 5.3 holds, then $\mathcal{A}$ is trivial as desired whenever $H_{1}$ is trivial and the $U/L=H_{1}/H_{1}$-wreath product satisfying Condition $(1)$ from Theorem 7.3 otherwise. If Statement $(3)$ of Lemma 5.3 holds, then $\mathcal{A}$ is the nontrivial $U/L=(H_{1}P_{1})/P_{1}$-wreath product. Due to Eq. (7), the section $U/L$ satisfies Condition $(1)$ from Theorem 7.3 or $4\nmid|U|$ and $|L|\neq 2$ as required.
Now let $n=p^{k}$, $k\geq 1$. As in the previous paragraph, $P_{1}$ denotes the least $\mathcal{A}$-subgroup containing $P$. One of Statements (1)(1)-$(3)$ of Lemma 7.2 holds for every basic set of $\mathcal{A}$ outside $D$. We divide the rest of the proof into two cases depending on existence of a basic set outside $D$ satisfying Statement $(3)$ of Lemma 7.2.
Case 1. Suppose that there is a basic set of $\mathcal{A}$ outside $D$, satisfying Statement $(3)$ of Lemma 7.2. Then $\mathcal{A}_{P_{1}}=\mathcal{T}_{P_{1}}$ and $P_{1}\nleq D$. If $P_{1}=G$, then $\mathcal{A}$ is trivial and we are done. Further, we assume that $G\setminus P_{1}\neq\varnothing$.
Observe that the condition of Lemma 7.2 holds for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$, i.e. $X\nsubseteq D$. Indeed, otherwise $\langle X\rangle$ is an $\mathcal{A}$-subgroup inside $D$. So $\langle X\rangle\geq P_{1}$ and hence
$D\geq\langle X\rangle\geq P_{1}\nleq D,$
a contradiction.
The above paragraph implies that one of Statements (1)(1)-$(3)$ from Lemma 7.2 holds for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$. Statement $(3)$ can not hold for $X\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$ because $P_{1}$ is a unique minimal $\mathcal{A}$-subgroup containing $P$. If Statement $(1)$ holds for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$, then $P_{1}\leq\operatorname{rad}(X)$ and hence $\mathcal{A}$ is the $U/L=P_{1}/P_{1}$-wreath product and $U/L$ satisfies Condition $(1)$ from Theorem 7.3 as required.
In view of the previous paragraph, we may assume that there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus P_{1}}$ for which Statement $(2)$ of Lemma 7.2 holds and $P_{1}\nleq\operatorname{rad}(X)$. Since $H$ is not an $\mathcal{A}$-subgroup, $A=\langle(X^{[p]})_{H}\rangle$ is an $\mathcal{A}$-subgroup of order $2$. Clearly, $\mathcal{A}_{A}=\mathbb{Z}A$. Note that $A\cap P_{1}$ is trivial because $\mathcal{A}_{P_{1}}=\mathcal{T}_{P_{1}}$. Therefore
(8) $\mathcal{A}_{A\times P_{1}}=\mathbb{Z}A\otimes\mathcal{T}_{P_{1}}$
by Lemma 3.2(2). If $A\times P_{1}=G$, then $\mathcal{A}$ is a nontrivial tensor product by Eq. (8) and we are done. So we may assume further that $A\times P_{1}<G$.
If $P\leq\operatorname{rad}(Y)$ for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus(A\times P_{1})}$, then $\mathcal{A}$ is the nontrivial $U/L=(A\times P_{1})/P_{1}$-wreath product and $U/L$ satisfies Condition $(1)$ from Theorem 7.3 as desired.
If there is $Y\in\mathcal{S}(\mathcal{A})_{G\setminus(A\times P_{1})}$ such that $P\nleq\operatorname{rad}(Y)$, then Statement $(2)$ of Lemma 7.2 holds for $Y$, i.e $\langle(Y^{[p]})_{H}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $H$. Eq. (8) and $P_{1}\nleq D$ imply that $A$ is a unique nontrivial $\mathcal{A}$-subgroup of $H$ and hence $\langle(Y^{[p]})_{H}\rangle=A$. The latter equality yields that $Y^{[p]}\subseteq A\times D$. The set $Y^{[p]}$ is an $\mathcal{A}$-set by Lemma 2.4 and consequently $\langle Y^{[p]}\rangle$ is an $\mathcal{A}$-subgroup of $A\times D$. However, $A$ is a unique nontrivial $\mathcal{A}$-subgroup of $A\times D$ by Eq. (8) and $P_{1}\nleq D$. Therefore $\langle Y^{[p]}\rangle=A$. From the latter equality, it follows that $Y\cap(A\times P_{1})\neq\varnothing$. As $A\times P_{1}$ is an $\mathcal{A}$-subgroup, we obtain $Y\subseteq A\times P_{1}$ in contrast to the definition of $Y$.
Case 2. Suppose that Statement $(1)$ or Statement $(2)$ of Lemma 7.2 holds for every basic set of $\mathcal{A}$ outside $D$. Clearly, $P\nleq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})$ such that $X\cap P\neq\varnothing$. So $N\geq P_{1}$, where $N$ is the $\mathcal{A}$-subgroup generated by all basic sets $X$ of $\mathcal{A}$ such that $P\nleq\operatorname{rad}(X)$.
We may assume that $|N|$ is divisible by $4$: otherwise $\mathcal{A}$ is the nontrivial $U/L=N/P_{1}$-wreath product and $4\nmid|U|$ and $|L|\neq 2$ as required. Assume that $|N|$ is divisible by $4$ or, equivalently, $N\geq H$. Then there is $Y\in\mathcal{S}(\mathcal{A})$ outside $D$ such that $P\nleq\operatorname{rad}(Y)$ because otherwise $N\leq D$ and hence $N\ngeq H$. Lemma 7.2 implies that $A=\langle(Y^{[p]})_{H}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $H$. As $H$ is not an $\mathcal{A}$-subgroup, $|A|=2$.
Since $N\geq H$, there is $Z\in\mathcal{S}(\mathcal{A})_{N}$ such that $P\nleq\operatorname{rad}(Z)$ and $Z_{H}\nsubseteq A$ because otherwise $N\leq A\times D$ and hence $N\ngeq H$. Note that $Z\nsubseteq H$ and hence
$Z_{D}\setminus\{e\}\neq\varnothing$
because otherwise $H=\langle A,Z\rangle$ is an $\mathcal{A}$-subgroup which contradicts to the assumption of the proposition. The group $\langle(Z^{[p]})_{H}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $H$ by Lemma 7.2. If $\langle(Z^{[p]})_{H}\rangle\neq A$, then $H=\langle A,(Z^{[p]})_{H}\rangle$ is an $\mathcal{A}$-subgroup, a contradiction to the assumption of the proposition. Therefore $\langle(Z^{[p]})_{H}\rangle=A$. This implies that $Z\cap(A^{\#}\times D^{\#})\neq\varnothing$,
(9) $P\nleq\operatorname{rad}(Z\cap(A\times D)),$
and
(10) $P\leq\operatorname{rad}(Z\setminus(A\times D)).$
Let $V=\langle Z\rangle$. The $S$-ring $\mathcal{A}_{V/A}$ is an $S$-ring over a cyclic group $V/A\cong C_{2p^{i}}$, $i\in\{1,\ldots,k\}$. Every basic set of a cyclotomic $S$-ring or a nontrivial tensor product over $V/A$ is regular. However, $Z$ and hence $Z^{\pi}$, where $\pi$ is the canonical epimorphism from $G$ to $G/A$, are nonregular because $Z_{D}\setminus\{e\}\neq\varnothing$ and $Z_{H}\nsubseteq A$. Therefore $\mathcal{A}_{V/A}$ can not be cyclotomic or a nontrivial tensor product. Lemma 6.2 implies that $\mathcal{A}$ is trivial or $|\operatorname{rad}(\mathcal{A}_{V/A})|>1$. Observe that $|\operatorname{rad}(Z^{\pi})|=1$ by Eq. (9) and the assumption that $H$ is not an $\mathcal{A}$-subgroup and hence the latter is impossible. Thus, $\mathcal{A}_{V/A}$ is trivial.
Let us prove that
$A\leq\operatorname{rad}(Z).$
Since $\mathcal{A}_{V/A}$ is trivial, $Z$ contains an element $h_{0}$ of order $2$. Let $z\in Z_{D}\setminus\{e\}\neq\varnothing$ and $h\in H$ such that $hz\in Z$. For every $i\in\mathbb{Z}$ coprime to $p$, there exists $m_{i}\in\mathbb{Z}$ such that $m_{i}\equiv 1\mod~2$ and $m_{i}\equiv i\mod~|z|$. One can see that $h_{0}^{m_{i}}=h_{0}\in X^{(m_{i})}\cap X$ and hence $X^{(m_{i})}=X$ by Lemma 2.3. Therefore $(hz)^{m_{i}}=hz^{i}\in X^{(m_{i})}=X$. This implies that
(11) $hD_{|z|}^{*}\subseteq Z$
for all $z\in Z_{D}\setminus\{e\}$ and $h\in H$ such that $hz\in Z$.
Observe that $Z\cap(A\times P^{\#})\neq\varnothing$ because $\mathcal{A}_{V/A}=\mathcal{T}_{V/A}$. Eq. (11) yields that $P^{\#}\subseteq Z$ or $aP^{\#}\subseteq Z$. If both of these two inclusions hold, then $A\leq\operatorname{rad}(Z)$ by Lemma 2.1. Thus, we may assume that exactly one of the above inclusions holds. Then $P^{\#}\subseteq Z$ or $P^{\#}\subseteq aZ$. Let $Z^{\prime}\in\{Z,aZ\}$ be such that $P^{\#}\subseteq Z^{\prime}$. From Eqs. (10) and (11), it follows that $Z^{\prime}\setminus P\neq\varnothing$ and $P\leq\operatorname{rad}(Z^{\prime}\setminus P)$. Therefore $|\operatorname{rad}(Z)|=1$ and $Z^{\prime}=\langle Z\rangle^{\#}$ by Lemma 2.2. We conclude that Statement $(3)$ of Lemma 7.2 holds for $Z^{\prime}$, a contradiction to the assumption of this case. Thus, $A\leq\operatorname{rad}(Z)$. Due to triviality of $\mathcal{A}_{V/A}$, we obtain
(12) $Z=V\setminus A~\text{and}~\mathcal{A}_{V}=\mathbb{Z}A\wr\mathcal{T}_{V/A}.$
The latter equality yields that $P_{1}=V$.
If $V=G$, then due to Eq. (12), $\mathcal{A}$ is the $U/L=A/A$-wreath product and $U/L$ satisfies Condition $(1)$ from Theorem 7.3. Further, we assume that $V<G$. If there exists $T\in\mathcal{S}(\mathcal{A})_{G\setminus V}$ such that $T\subseteq D$, then $\langle T\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $D$. So $\langle T\rangle\cap V\neq\varnothing$ and hence $V\subseteq\langle T\rangle\leq D$, a contradiction to the definition of $Z$. Therefore every basic set outside $V$ is not a subset of $D$ and hence Statement $(1)$ or $(2)$ of Lemma 7.2 holds for every such basic set. If Statement $(1)$ of Lemma 7.2 holds for every basic set $X$ outside $V$, i.e. $P\leq\operatorname{rad}(X)$, then $V=P_{1}\leq\operatorname{rad}(X)$. This implies that $\mathcal{A}$ is the $U/L=V/V$-wreath product and $U/L$ satisfies Condition $(1)$ from Theorem 7.3. If Statement $(2)$ of Lemma 7.2 holds for some $T\in\mathcal{S}(\mathcal{A})_{G\setminus V}$, then $\langle(T^{[p]})_{H}\rangle=A$ because $A$ is a unique nontrivial $\mathcal{A}$-subgroup of $H$. Together with Lemma 2.4, this yields that $W=\langle T^{[p]}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $A\times D$. Since $T$ lies outside $V$, we have $(T^{[p]})_{D}\neq\{e\}$ and hence $|W|$ is divisible by $p$. Therefore $W\cap Z\neq\varnothing$. Thus,
$Z\subseteq W\leq A\times D,$
a contradiction to the definition of $Z$. ∎
Proposition 7.5.
If $H$ is not an $\mathcal{A}$-subgroup and $\mathcal{A}$ is a nontrivial $U/L$-wreath product for an $\mathcal{A}$-section $U/L$ such that $4\nmid|U|$ and $|L|\neq 2$, then Theorem 7.3 holds.
Proof.
Note that $U$ is cyclic because $4\nmid|U|$. If $|\operatorname{rad}(\mathcal{A}_{U})|=1$, then $U/L$ satisfies Condition $(3)$ from Theorem 7.3 and we are done. From now on throughout the proof of this proposition, we assume that
(13) $|\operatorname{rad}(\mathcal{A}_{U})|>1.$
Suppose that $|U|$ is odd. Then $|U|=pq$ or $|U|$ is a $p$-power. In the former case, Eq. (13) and Lemma 6.12 imply that $\mathcal{A}_{U}=\mathcal{A}_{L}\wr\mathcal{A}_{U/L}$. So $\mathcal{A}=\mathcal{A}_{L}\wr\mathcal{A}_{G/L}$ and Theorem 7.3 holds as required. In the latter one, $\mathcal{A}_{U}$ is the nontrivial $U_{1}/L_{1}$-wreath product for an $\mathcal{A}_{U}$-section $U_{1}/L_{1}$ of $U$ such that $L_{1}$ is the least nontrivial $\mathcal{A}_{U}$-subgroup and $|\operatorname{rad}(\mathcal{A}_{U_{1}})|=1$ by Eq. (13) and Lemma 6.10. By the definitions of $U_{1}$ and $L_{1}$, we have $U_{1}\leq U$ and $L_{1}\leq L$. This implies that $\mathcal{A}$ is the $U_{1}/L_{1}$-wreath product and $U_{1}/L_{1}$ satisfies Condition $(3)$ from Theorem 7.3 as required.
Now suppose that $|U|$ is even. Then $|U|=2m$, where $m$ is odd, because $4\nmid|U|$. Let $A$ and $W$ be the Sylow $2$-subgroup and the Hall $2^{\prime}$-subgroup of $U$, respectively. If $|W|=1$, then $|L|=|U|=2$, a contradiction to the assumption $|L|\neq 2$ of the lemma. In what follows, we assume that
$|W|>1.$
The least $\mathcal{A}_{U}$-subgroup containing $A$ and a maximal $\mathcal{A}_{U}$-subgroup of $W$ are denoted by $A_{1}$ and $W_{1}$, respectively. Assume that $|W_{1}|=1$. Then $\mathcal{A}_{U}$ is trivial or $\mathcal{A}_{U}=\mathcal{A}_{A_{1}}\wr\mathcal{A}_{U/A_{1}}$ by Lemma 5.3 applied to $\mathcal{A}=\mathcal{A}_{U}$, $H_{1}=W_{1}$, and $P_{1}=A_{1}$. In the former case, we obtain a contradiction to Eq. (13). In the latter one, $A_{1}$ is the least nontrivial $\mathcal{A}_{U}$-subgroup because otherwise $A_{1}$ has a nontrivial $\mathcal{A}_{U}$-subgroup of odd order and hence $|W_{1}|>1$, a contradiction to the assumption $|W_{1}|=1$. So $A_{1}\leq L$. Together with $\mathcal{A}=\mathcal{A}_{U}\wr_{U/L}\mathcal{A}_{G/L}$ and $\mathcal{A}_{U}=\mathcal{A}_{A_{1}}\wr\mathcal{A}_{U/A_{1}}$, this yields that $\mathcal{A}=\mathcal{A}_{A_{1}}\wr\mathcal{A}_{G/A_{1}}$ and Theorem 7.3 holds as desired. In what follows, we assume that
$|W_{1}|>1.$
Lemma 7.6.
With the above notation, Theorem 7.3 holds unless $L\geq M$ for some minimal nontrivial $\mathcal{A}_{W_{1}}$-subgroup $M$.
Proof.
If $|L\cap W_{1}|>1$, then $L\cap W_{1}$ being an $\mathcal{A}_{W_{1}}$-subgroup contains a minimal nontrivial $\mathcal{A}_{W_{1}}$-subgroup and we are done. So we may assume that $|L\cap W_{1}|=1$. Since $|L|\neq 2$ (the assumption of the proposition) and $|W_{1}|>1$, this can happen only if $n=pq$, $U=W_{1}\times L$, $|W_{1}|$ is prime, $|L|$ is twice prime, and $L$ does not have a nontrivial $\mathcal{A}_{L}$-subgroup of odd order. Lemma 6.12 applied to $\mathcal{A}_{L}$ implies that
$\mathcal{A}_{L}=\mathcal{T}_{L}~\text{or}~\mathcal{A}_{L}=\mathcal{A}_{L_{0}}\wr\mathcal{A}_{L/L_{0}}$
for an $\mathcal{A}_{L}$-subgroup $L_{0}$ of order $2$.
In the former case, Lemma 6.2 applied to $\mathcal{A}_{U}$ yields that $\mathcal{A}_{U}=\mathcal{A}_{W_{1}}\otimes\mathcal{T}_{L}$. Therefore $|\operatorname{rad}(\mathcal{A}_{U})|=1$, a contradiction to Eq. (13). In the latter one, Eq. (13) and Lemma 6.2 applied to $\mathcal{A}_{U}$ yield that $\mathcal{A}_{U}$ is the nontrivial $(W_{1}L_{0})/L_{0}$-wreath product. As $L_{0}\leq L$ and $\mathcal{A}$ is the $U/L$-wreath product, the $S$-ring $\mathcal{A}$ is the nontrivial $(W_{1}L_{0})/L_{0}$-wreath product. Since $|L_{0}|=2$, we have $\mathcal{A}_{L_{0}}=\mathbb{Z}L_{0}$. So $\mathcal{A}_{W_{1}L_{0}}=\mathcal{A}_{W_{1}}\otimes\mathbb{Z}L_{0}$ by Lemma 3.2(2). Therefore $\mathcal{A}_{L_{0}}$ is $\otimes$-complemented in $\mathcal{A}_{W_{1}L_{0}}$. Thus, the section $(W_{1}L_{0})/L_{0}$ satisfies Condition $(2)$ from Theorem 7.3 and we are done. ∎
By Lemma 7.6, we may assume till the end of the proof that
$L\geq M$
for some minimal nontrivial $\mathcal{A}_{W_{1}}$-subgroup $M$. One can see that $\mathcal{A}_{W_{1}}$ is the $V/M$-wreath product (possibly, trivial), where $V$ is an $\mathcal{A}_{W_{1}}$-subgroup with $|\operatorname{rad}(\mathcal{A}_{V})|=1$. Indeed, this follows from Lemma 6.2 and Lemma 6.12 if $|W_{1}|=pq$ and from Lemma 6.2 and Lemma 6.10 if $|W_{1}|$ is a $p$-power. Lemma 5.4 implies that
(14) $\mathcal{A}_{W_{1}A_{1}}=\mathcal{A}_{W_{1}}\star\mathcal{A}_{A_{1}}=(\mathcal{A}_{V}\wr_{V/M}\mathcal{A}_{W_{1}/M})\star\mathcal{A}_{A_{1}}.$
Let $W_{1}A_{1}=U$. If $W_{1}\cap A_{1}$ is trivial, then $\mathcal{A}_{A_{1}}$ is primitive and hence $\mathcal{A}_{A_{1}}=\mathcal{T}_{A_{1}}$ by Lemma 6.1. Eq. (14) implies that $\mathcal{A}_{U}=\mathcal{A}_{W_{1}}\otimes\mathcal{A}_{A_{1}}$ and consequently $\mathcal{A}_{U}$ is the $U_{1}/L_{1}=(VA_{1})/M$-wreath product (possibly, trivial). As $M\leq L$, we conclude that $\mathcal{A}$ is the $U_{1}/L_{1}$-wreath product. Since $|\operatorname{rad}(\mathcal{A}_{V})|=1$ and
$\mathcal{A}_{U_{1}}=\mathcal{A}_{V}\otimes\mathcal{A}_{A_{1}}\cong\mathcal{A}_{V}\otimes\mathcal{T}_{A_{1}},$
we obtain $|\operatorname{rad}(\mathcal{A}_{U_{1}})|=1$. Thus, $U_{1}/L_{1}$ satisfies Condition $(3)$ from Theorem 7.3 and we are done.
Now let $W_{1}A_{1}<U$. This implies that
$W_{1}<W.$
In this case, $|W_{1}|$ is a $p$-power and hence $M$ is the unique least $\mathcal{A}_{W_{1}}$-subgroup. Lemma 5.3 applied to $\mathcal{A}=\mathcal{A}_{U}$, $H_{1}=W_{1}$, and $P_{1}=A_{1}$ yields that $\mathcal{A}_{U}=\mathcal{A}_{W_{1}}\wr\mathcal{A}_{U/W_{1}}$ or
(15) $\mathcal{A}_{U}=\mathcal{A}_{W_{1}A_{1}}\wr_{(W_{1}A_{1})/A_{1}}\mathcal{A}_{U/A_{1}}.$
In the former case, $\mathcal{A}$ is the nontrivial $V/M$-wreath product because $\mathcal{A}_{W_{1}}=\mathcal{A}_{V}\wr_{V/M}\mathcal{A}_{W_{1}/M}$ and $M\leq L$. Since $|\operatorname{rad}(\mathcal{A}_{V})|=1$, the section $V/M$ satisfies Condition $(3)$ from Theorem 7.3 and we are done.
Suppose that Eq. (15) holds. If $\mathcal{A}_{A_{1}}\neq\mathcal{T}_{A_{1}}$, then $A_{1}$ has a nontrivial $\mathcal{A}_{A_{1}}$-subgroup by Lemma 6.1 applied to $A_{1}$. This $\mathcal{A}_{A_{1}}$-subgroup is of odd order because $A_{1}$ is the least $\mathcal{A}_{U}$-subgroup containing $A$ and hence $|A_{1}\cap W_{1}|>1$. In view of Eqs. (14) and (15), we conclude that $\mathcal{A}_{U}$ is the nontrivial $W_{1}/(W_{1}\cap A_{1})$-wreath product. Since $M$ is the unique least $\mathcal{A}_{W_{1}}$-subgroup, we have $M\leq A_{1}\cap W_{1}$ and consequently $\mathcal{A}_{U}$ is the $W_{1}/M$-wreath product. As $\mathcal{A}_{W_{1}}=\mathcal{A}_{V}\wr_{V/M}\mathcal{A}_{W_{1}/M}$, we conclude that $\mathcal{A}_{U}$ is the nontrivial $V/M$-wreath product. Observe that $\mathcal{A}$ is the $V/M$-wreath product because $M\leq L$. The section $V/M$ satisfies Condition $(3)$ from Theorem 7.3 because $|\operatorname{rad}(\mathcal{A}_{V})|=1$ and we are done.
Now let $\mathcal{A}_{A_{1}}=\mathcal{T}_{A_{1}}$. Then $|W_{1}\cap A_{1}|=1$. Eq. (14) implies that
(16) $\mathcal{A}_{W_{1}A_{1}}=\mathcal{A}_{W_{1}}\otimes\mathcal{T}_{A_{1}}.$
If $A_{1}\neq A$, then $|W_{1}A_{1}|=2pq$, $U=W_{1}A_{1}$, and $L\in\{W_{1},A_{1}\}$. So $U/L$ satisfies Condition $(2)$ from Theorem 7.3 by Eq. (16) and we are done. Further, we assume that $A_{1}=A$.
Lemma 7.7.
With the above notation, let $X\in\mathcal{S}(\mathcal{A})$ be such that $X_{D}\nsubseteq W_{1}$. Then $A\leq\operatorname{rad}(X)$.
Proof.
If $X\subseteq U$, then $X\subseteq U\setminus(W_{1}A_{1})$. So $\operatorname{rad}(X)\geq A_{1}\geq A$ by Eq. (15). In the sequel, we assume that $X\nsubseteq U$. This implies that $L\leq\operatorname{rad}(X)$ because $\mathcal{A}$ is the $U/L$-wreath product. Let $R=\langle X\rangle$. Since $X_{D}\nsubseteq W_{1}$, we obtain
$R\cap(W\setminus W_{1})\neq\varnothing.$
Together with Eq. (15), this implies that $R\geq A_{1}\geq A$. The image of $Y\subseteq G$ under the canonical epimorphism from $G$ to $G/\operatorname{rad}(X)$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu$. Clearly, $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)|=1$ and hence $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is indecomposable.
Assume the contrary that $A\nleq\operatorname{rad}(X)$. Then $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu\geq\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$. Due to Eq. (15), we have $\operatorname{rad}(X)\cap D\leq W_{1}$ and
(17) $\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu\leq\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu)$
for every $Y\in\mathcal{S}(\mathcal{A})_{R}$ inside $R\cap(U\setminus(W_{1}A))$. Observe that $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is nontrivial because $\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$ is a proper nontrivial $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$-subgroup.
Suppose that $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is cyclic. Then $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is isomorphic to $C_{p^{i}}$ or $C_{2p^{i}}$, $i\geq 1$. As $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is indecomposable, we conclude that $\operatorname{rad}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu})$ is trivial. So $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is cyclotomic or a nontrivial tensor product by Lemma 6.2. However, every basic set of such $S$-ring over $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ has trivial radical which contradicts to Eq. (17).
Now suppose that $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is noncyclic. Then $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu\cong E_{4}\times C_{p^{i}}$, $i\geq 1$. In view of Eq. (17), the Hall $2^{\prime}$-subgroup of $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is not an $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$-subgroup. So the Sylow $2$-subgroup of $\widehat{G}$ is not an $\widehat{\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}}-$subgroup by Lemma 4.1(1). Since $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is indecomposable and nontrivial, so is $\widehat{\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}}$. Therefore $\widehat{\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}}$ is a nontrivial tensor product by Proposition 7.4 applied to $\widehat{\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}}$. Thus,
(18) $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}=\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{1}}\otimes\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{2}}$
for some $\mathcal{A}_{R}$-subgroups $R_{1}$ and $R_{2}$ by Lemma 4.1(3). As $\operatorname{rad}(X)\geq L$ and $|L|\neq 2$, we may assume without loss of generality that $|\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{1}|$ is a power of $2$. Due to Eqs. (17) and (18), $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{1}\ngeq\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$. Therefore $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{2}$ is a cyclic group containing $\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$. Lemma 6.2 and Eq. (17) yield that $|\operatorname{rad}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{2}})|\neq 1$. Together with Eq. (18), this implies that $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is decomposable, a contradiction. ∎
From Lemma 7.7 it follows that
$\mathcal{A}=\mathcal{A}_{W_{1}A}\wr_{(W_{1}A)/A}\mathcal{A}_{G/A}~\text{or}~\mathcal{A}=\mathcal{A}_{W_{1}H}\wr_{(W_{1}H)/A}\mathcal{A}_{G/A}.$
In the first case, $\mathcal{A}_{A}$ is $\otimes$-complemented in $\mathcal{A}_{W_{1}}$ by Eq. (16). So the section $(W_{1}A)/A$ satisfies Condition $(2)$ from Theorem 7.3 and we are done.
In the second case, $W_{1}H$ is an $\mathcal{A}$-subgroup. Let $a$ be a nontrivial element of $A$, $b\in H\setminus A$, and $T$ a basic set of $\mathcal{A}$ containing $b$. Clearly, $W_{1}H=W_{1}A\times\langle b\rangle$, $W_{1}A$ is an $\mathcal{A}_{W_{1}H}$-subgroup, and $W_{1}$ is the largest $\mathcal{A}_{W_{1}H}$-subgroup of odd order. Therefore the $S$-ring $\mathcal{A}_{W_{1}H}$, the basic set $T$, and the subgroup $W_{1}$ satisfy [14, Hypothesis $5.1$]. The structure of a basic set containing an involution of an $S$-ring satisfying [14, Hypothesis $5.1$] is described in [14, Lemma $5.5$]. Applying this lemma to $T$ and using the inclusion $\{a\}=A^{\#}\in\mathcal{S}(\mathcal{A})$ one can deduce that
$T=W_{0}Ab~\text{or}~T=W_{0}b$
for some $\mathcal{A}_{W_{1}}$-subgroup $W_{0}$. To complete the proof of the lemma, let us consider the above two cases separately.
Let $T=W_{0}Ab$. As $H=\langle Ab\rangle$ is not an $\mathcal{A}$-subgroup, $T\neq Ab$ and hence $|W_{0}|>1$. So $W_{0}$ satisfies the condition of [14, Lemma $5.4$]. Applying this lemma to $\mathcal{A}_{W_{1}H}$, we conclude that $\mathcal{A}_{W_{1}H}$ is the $(W_{1}A)/(W_{0}A)$-wreath product. Therefore $\mathcal{A}$ is the $(W_{1}A)/A$-wreath product. Again, the section $(W_{1}A)/A$ satisfies Condition $(2)$ from Theorem 7.3 by Eq. (16) and we are done.
Now let $T=W_{0}b$, then $W_{2}=\langle T,W_{1}\rangle$ is an $\mathcal{A}_{W_{1}H}$-subgroup such that $|W_{2}\cap A|=1$ and $W_{1}H=W_{2}\times A$. Lemma 3.2(2) implies that $\mathcal{A}_{W_{1}H}=\mathcal{A}_{W_{2}}\otimes\mathcal{A}_{A}$. Therefore $(W_{1}H)/A$ satisfies Condition $(2)$ from Theorem 7.3 as desired. ∎
Theorem 7.3 immediately follows from Propositions 7.4 and 7.5.
Theorem 7.8.
Suppose that $D$ is not an $\mathcal{A}$-subgroup. Then $\mathcal{A}$ is trivial, or a nontrivial tensor product, or a nontrivial $U/L$-wreath product for some $\mathcal{A}$-section $U/L$ satisfying one of the following conditions:
  1. (1)
    $|U/L|\leq 2$;
  2. (2)
    $\mathcal{A}_{U/L}$ is $\otimes$-complemented in $\mathcal{A}_{G/L}$;
  3. (3)
    $4\nmid|G/L|$ and $|\operatorname{rad}(\mathcal{A}_{G/L})|=1$.
Proof.
Let $\widehat{\mathcal{A}}$ be the $S$-ring dual to $\mathcal{A}$ over $\widehat{G}\cong G$. Since $D$ is not an $\mathcal{A}$-subgroup, the group $D^{\bot}\cong E_{4}$ is not an $\widehat{\mathcal{A}}-$subgroup by Lemma 4.1(1). So Theorem 7.3 holds for $\widehat{\mathcal{A}}$. If $\widehat{\mathcal{A}}$ is trivial, then so is $\mathcal{A}$ and we are done. If $\widehat{\mathcal{A}}$ is a nontrivial tensor product, then so is $\mathcal{A}$ by Lemma 4.1(3) and we are done.
[…]
Lemma 4.1(2) implies that $\widehat{\mathcal{A}_{G/L}}=\widehat{\mathcal{A}}_{\widehat{U}}$ and $\widehat{\mathcal{A}_{U/L}}=\widehat{\mathcal{A}}_{\widehat{U}/\widehat{L}}$. So if $\widehat{L}$ is $\otimes$-complemented in $\widehat{\mathcal{A}}_{\widehat{U}}$, then $\mathcal{A}_{U/L}$ is $\otimes$-complemented in $\mathcal{A}_{G/L}$ by Lemma 4.1(3) as required.
[…]
7.3. Dense $S$-rings
The lemma below is straightforward.
Lemma 7.9.
Every $S$-ring over $H\cong E_{4}$ is normal and equals $\operatorname{Cyc}(K,H)$, where $K\leq\operatorname{Aut}(H)$ is of order at most $3$.
Remark 7.10.
In Lemma 7.9, we have $\operatorname{Cyc}(K_{H},H)=\mathbb{Z}H$ if $|K_{H}|=1$, $\operatorname{Cyc}(K_{H},H)\cong\mathbb{Z}C_{2}\wr\mathbb{Z}C_{2}$ if $|K_{H}|=2$, and $\operatorname{Cyc}(K_{H},H)=\mathcal{T}_{H}$ if $|K_{H}|=3$.
Further throughout this subsection, we assume that $H$ and $D$ are $\mathcal{A}$-subgroups. Denote two distinct elements from $H^{\#}$ by $a$ and $b$. Given $X\in\mathcal{S}(\mathcal{A})$, put $\lambda_{X}=|X\cap Hx|$, where $x\in X$. Due to Lemma 2.1, $\lambda_{X}$ does not depend on $x$. Clearly, $1\leq\lambda_{X}\leq|X_{H}|\leq 3$.
Lemma 7.11.
Let $\mathcal{A}$ be dense and $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$. Then $X=X_{H}\times X_{D}$ or $\lambda_{X}=1$.
Proof.
Assume that $X\neq X_{H}\times X_{D}$. Then $\lambda_{X}<|X_{H}|$. We are done if $\lambda_{X}=1$. So it remains to consider only the case when $1<\lambda_{X}<|X_{H}|\leq 3$. Since $H\cong E_{4}$, this is possible only if $X_{H}=H^{\#}=\{a,b,ab\}$ and $\lambda_{X}=2$.
Assume that two latter equalities hold. Then $Y=(X_{H}\times X_{D})\setminus X$ is an $\mathcal{A}$-set containing exactly one element from $Hx$ for every $x\in X_{D}$. Together with Lemma 3.2(1) and the equality $X_{H}=H^{\#}$, this implies that $Y$ is exactly a basic set of $\mathcal{A}$ with
(19) $\lambda_{Y}=1.$
Let $x,x^{\prime}\in X_{D}$ be of the same order. Then there is a positive integer $m$ coprime to $2n$ such that $x^{m}=x^{\prime}$. Since $\lambda_{X}=2$, the sets $H_{x}=Xx^{-1}\cap H$ and $H_{x^{\prime}}=X(x^{\prime})^{-1}\cap H$ are subsets of $H^{\#}$ of size $2$. Therefore there is $h\in H_{x}\cap H_{x^{\prime}}$. One can see that $(hx)^{m}=hx^{\prime}\in X^{(m)}\cap X$. So $X^{(m)}=X$ by Lemma 2.3. This yields that
$(H_{x}x)^{m}=H_{x}x^{m}=H_{x}x^{\prime}=H_{x^{\prime}}x^{\prime}$
and hence $H_{x}=H_{x^{\prime}}$.
Let $l$ be the maximum of orders of the elements from $X_{D}$. Due to the above discussion, there is $H_{0}\subseteq H$ such that $|H_{0}|=2$ and $H_{x}=H_{0}$ for every $x\in(X_{D}\cap D_{l}^{*})$. So
(20) $X_{D}\cap D_{l}^{*}\subseteq Y_{h_{0}},$
where $h_{0}$ is a unique nontrivial element from $H\setminus H_{0}$. Let $h\in H_{0}$. From Eq. (19) it follows that $Y_{h_{0}}\cap Y_{h}=\varnothing$. Together with $Y_{h_{0}}\cup Y_{h}\subseteq X_{D}$ (Lemma 3.2(1)) and Eq. (20), this implies that
(21) $Y_{h}\subseteq X_{D}\setminus D_{l}^{*}.$
Observe that $|Y_{h_{0}}|=|Y_{h}|$ by Lemma 2.1. In view of Eqs. (20) and (21), to obtain a contradiction, it suffices to show that
(22) $|X_{D}\cap D_{l}^{*}|>|X_{D}\setminus D_{l}^{*}|.$
The set $X_{D}$ is a basic set of $\mathcal{A}_{D}$ (Lemma 3.2(1)). If $|X_{D}\setminus D_{l}^{*}|=0$, then Eq. (22) is obvious. Suppose that $|X_{D}\setminus D_{l}^{*}|>0$, i.e. $X_{D}$ is nonregular. If $n=pq$, then Lemma 6.12 implies that $\mathcal{A}_{D}$ is trivial or a nontrivial generalized wreath product. So $X_{D}=D^{\#}$ or $X_{D}$ is a union of some $L$-cosets, where $L=\operatorname{rad}(X_{D})$ is a nontrivial $\mathcal{A}_{D}$-subgroup. In the former case, $|X_{D}\cap D_{l}^{*}|=(p-1)(q-1)$ and $|X_{D}\setminus D_{l}^{*}|=p+q-2$, whereas in the latter one $|X_{D}\cap D_{l}^{*}|=r(|L|-1)$ and $|X_{D}\setminus D_{l}^{*}|=r$, where $r=|X_{D}/L|$. If $n=p^{k}$, then by Lemma 6.11, we obtain $X_{D}=U\setminus L$ for some $\mathcal{A}$-subgroups $U\cong C_{p^{l}}>L$ and consequently $|X_{D}\cap D_{l}^{*}|=p^{l-1}(p-1)$ and $|X_{D}\setminus D_{l}^{*}|=p^{l-1}-|L|$. Since $p$ and $q$ are odd, Eq. (22) holds in all cases and we are done. ∎
Theorem 7.12.
Let $\mathcal{A}$ be dense. Suppose that $|\operatorname{rad}(\mathcal{A}_{D})|=1$ and $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$. Then:
  1. (1)
    $\mathcal{A}$ is normal, cyclotomic, and $2$-minimal;
  2. (2)
    $|\operatorname{Aut}(\mathcal{A})^{H}|\leq 12$.
Proof.
By the condition $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$ and Lemma 3.2(2), we have $\mathcal{A}_{H}\neq\mathbb{Z}H$. Then Lemma 7.9 and Remark 7.10 imply that $\mathcal{A}_{H}=\operatorname{Cyc}(K_{H},H)$, where $K_{H}\leq\operatorname{Aut}(H)$ is of order $2$ or $3$, and a unique nonsingleton basic set of $\mathcal{A}_{H}$ is of size $|K_{H}|$. We divide the rest of the proof into two cases depending on whether $\mathcal{A}_{D}$ is cyclotomic or not.
Case 1: $\mathcal{A}_{D}$ is noncyclotomic. The condition $|\operatorname{rad}(\mathcal{A}_{D})|=1$ of the theorem and Lemma 6.2 imply that $\mathcal{A}_{D}$ is trivial whenever $n=p^{k}$. The same arguments and Lemma 6.12 imply that $\mathcal{A}_{D}$ is trivial in case $n=pq$. Thus, $\mathcal{A}_{D}$ is trivial in any case and hence $D^{\#}\in\mathcal{S}(\mathcal{A}_{D})$. Moreover, $n$ is not prime by Lemma 6.7. Therefore $n=pq$ or $n=p^{k}$ for $k\geq 2$.
Let $n=pq$. Then $H$ is a maximal $\mathcal{A}$-subgroup whose order is not divisible by $p$ and $D$ is the least $\mathcal{A}$-subgroup whose order is divisible by $p$. By Lemma 5.3 applied to $\mathcal{A}$, $H$, and $D$, we have $\mathcal{A}=\mathcal{A}_{H}\star\mathcal{A}_{D}$. Since $H\cap D$ is trivial, we obtain $\mathcal{A}=\mathcal{A}_{H}\otimes\mathcal{A}_{D}$, a contradiction to the assumption of the theorem.
Now let $n=p^{k}$, $k\geq 2$. Since $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$ by the condition of the theorem, there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$ such that
(23) $X\neq X_{H}\times X_{D}.$
Note that $X_{H}\in\mathcal{S}(\mathcal{A}_{H})$ and $X_{D}=D^{\#}$ by Lemma 3.2(1). Eq. (23) implies that $|X_{H}|>1$. So $|X_{H}|\in\{2,3\}$ by Lemma 7.9. Eq. (23) and Lemma 7.11 yield that $\lambda_{X}=1$. If $P\nleq\operatorname{rad}(X_{h,p^{l}})$ (see Notation for $X_{h,p^{l}}$) for some $h\in X_{H}$ and $l\in\{2,\ldots,k\}$, then the set $Y=X^{[p]}$, which is an $\mathcal{A}$-set by Lemma 2.4, lies outside $H\cup D$ and $Y_{D}\subsetneq D^{\#}$, a contradiction to Lemma 3.2(1). Therefore
(24) $P\leq\operatorname{rad}(X_{h}\setminus X_{h,p})$
for every $h\in X_{H}$.
Lemma 2.5 applied to $X$, $h\in X_{H}$, and $l=p$ implies that the sets $X_{h,p}$, $h\in X_{H}$, are orbits of some $K\leq\operatorname{Aut}(P)$. This yields that all of these sets are of the same size $|K|$. Since $X_{D}=D^{\#}$, we have
$\bigcup\limits_{h\in X_{H}}X_{h,p}=P^{\#}$
and hence
$\sum\limits_{h\in X_{H}}|X_{h,p}|=p-1.$
Thus,
(25) $|X_{h,p}|=(p-1)/|X_{H}|.$
Further, we are going to compute $\sum\limits_{h\in X_{H}}\underline{X_{h,p}}^{2}$. Recall that $|X_{H}|\in\{2,3\}$. If $|X_{H}|=2$ and (p1)/2(p-1)/2 is odd, then $\operatorname{Cyc}(K,P)$ is an antisymmetric $S$-ring of rank $3$. Using the formulas for the intersection numbers of an antisymmetric association scheme of rank $3$ (see, e.g. [3, Exercise 2.7.57]), one can compute that
(26) $\sum\limits_{h\in X_{H}}\underline{X_{h,p}}^{2}=\frac{p-1}{2}\underline{P}^{\#}.$
If $|X_{H}|=2$ and (p1)/2(p-1)/2 is even or $|X_{H}|=3$, then $X_{h,p}=X_{h,p}^{-1}$ for every $h\in X_{H}$ and the required sum is a special case of the sum computed in [31, Lemma 5.2]. So applying this lemma, we obtain
(27) $\sum\limits_{h\in X_{H}}\underline{X_{h,p}}^{2}=(p-1)e+(\frac{p-1}{|X_{H}|}-1)\underline{P}^{\#}.$
Put
$\xi=\sum\limits_{h\in X_{H}}\underline{X_{h}}^{2}\in\mathbb{Z}D.$
One can compute directly that
$\underline{X}^{2}=\xi+\sum\limits_{h,h^{\prime}\in X_{H}}hh^{\prime}\underline{X_{h}}\cdot\underline{X_{h^{\prime}}}\in\mathcal{A}.$
Observe that only elements from $G\setminus D$ enter the second sum on the right-hand side of the above equality. Since $D$ is an $\mathcal{A}$-subgroup, we conclude that $\xi\in\mathcal{A}$. Further, one can compute $\xi$ as follows:
$\xi=\sum\limits_{h\in X_{H}}(\underline{X_{h}\setminus X_{h,p}}+\underline{X_{h,p}})^{2}=\sum\limits_{h\in X_{H}}(\underline{X_{h}\setminus X_{h,p}})^{2}+2\sum\limits_{h\in X_{H}}\underline{X_{h}\setminus X_{h,p}}\cdot\underline{X_{h,p}}+\sum\limits_{h\in X_{H}}\underline{X_{h,p}}^{2}.$
Each element appearing in the first of three last sums enters with a coefficient divisible by $p$ due to Eq. (24) and hence
$\sum\limits_{h\in X_{H}}(\underline{X_{h}\setminus X_{h,p}})^{2}=p\eta$
for some $\eta\in\mathbb{Z}D$. Recall that $X_{h,p}\subseteq P\leq\operatorname{rad}(X_{h}\setminus X_{h,p})$ for every $h\in X_{H}$, where the latter inequality holds by Eq. (24). Therefore
$\sum\limits_{h\in X_{H}}\underline{X_{h}\setminus X_{h,p}}\cdot\underline{X_{h,p}}=\sum\limits_{h\in X_{H}}|X_{h,p}|\underline{X_{h}\setminus X_{h,p}}=2(p-1)/|X_{H}|\sum\limits_{h\in X_{H}}(\underline{X_{h}}-\underline{X_{h,p}}),$
where the first equality follows from Eq. (1), whereas the second one from Eq. (25). Thus,
$\xi=p\eta+2(p-1)/|X_{H}|\sum\limits_{h\in X_{H}}(\underline{X_{h}}-\underline{X_{h,p}})+\sum\limits_{h\in X_{H}}\underline{X_{h,p}}^{2}.$
The above expression for $\xi$ implies that every element from $D\setminus P$ enters $\xi$ with the coefficient equal to $2(p-1)/|X_{H}|$ modulo $p$. On the other hand, this expression and Eqs. (26) and (27) imply that every element from $P^{\#}$ enters $\xi$ with coefficient equal to $(p-1)/|X_{H}|$ or $(p-1)/|X_{H}|-1$ modulo $p$. It is easy to see that none of the above numbers is equal to $2(p-1)/|X_{H}|$ modulo $p$. However, each element from $D^{\#}$ must enter $\xi$ with the same coefficient because $D^{\#}\in\mathcal{S}(\mathcal{A})$, a contradiction.
Case 2: $\mathcal{A}_{D}$ is cyclotomic. In this case, all basic sets of $\mathcal{A}_{D}$ are regular. Let $K_{D}\leq\operatorname{Aut}(D)$ be such that $\mathcal{A}_{D}=\operatorname{Cyc}(K_{D},D)$.
Claim 1.
Let $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$. Then $X^{(r)}\in\mathcal{S}(\mathcal{A})$ for every $r\in\mathbb{Z}$.
Proof.
Without loss of generality, we may assume that $r$ is prime. If $r$ is coprime to $2n$, then the claim follows from Lemma 2.3. Lemma 3.2(1) implies that $X_{D}\in\mathcal{S}(\mathcal{A}_{D})$. If $r=2$, then $X^{(r)}=X^{(2)}=X_{D}^{(2)}\in\mathcal{S}(\mathcal{A}_{D})$ by Lemma 2.3.
It remains to consider the case when $r$ is an odd prime divisor of $n$. Put $Y=X^{(r)}$. Let $R$ be the (unique) subgroup of $D$ of order $r$. Since $\mathcal{A}_{D}$ is cyclotomic, $R$ is an $\mathcal{A}$-subgroup. Observe that $R\nleq\operatorname{rad}(X_{D})$. Indeed, if $n=p^{k}$, then this follows from Lemma 6.8(2), whereas if $n=pq$, then this immediately follows from the the fact that $\mathcal{A}$ is cyclotomic with trivial radical. Therefore $R\nleq\operatorname{rad}(X)$ and hence $Y=X^{[r]}$. Lemma 2.4 yields that $Y$ is an $\mathcal{A}$-set.
If $Y_{D}$ is trivial, then $Y=X_{H}$ and we are done by Lemma 3.2(1). Further, we assume that $Y_{D}$ is nontrivial. Let us show that
(28) $Y_{D}\in\mathcal{S}(\mathcal{A}_{D})$
One can see that $Y_{D}=X_{D}^{(r)}$. If $n=p^{k}$, then $|Y_{D}|=|X_{D}|$ because all the elements of $X_{D}$ lie in pairwise distinct $R$-cosets (Lemma 6.8(1)). From Lemma 6.8(2) it follows that all nontrivial basic set of $\mathcal{A}_{D}$ has the same size and hence Eq. (28) holds. If $n=pq$, then $R$ and the Hall $r^{\prime}$-subgroup $U$ of $D$ are $\mathcal{A}$-subgroups because $\mathcal{A}_{D}$ is cyclotomic. So $H$, $U$, and $R$ are $\mathcal{A}$-subgroups such that $G=H\times U\times R$. As $R\nleq\operatorname{rad}(X)$, one can see that $Y=X_{H\times U}$ and consequently $Y_{D}=(X_{H\times U})_{D}=X_{U}$. Therefore $Y_{D}\in\mathcal{S}(\mathcal{A}_{U})$ by Lemma 3.2(1) applied to $G_{1}=H\times R$ and $G_{2}=U$. Thus, Eq. (28) holds.
To complete the proof, let us show that $Y\in\mathcal{S}(\mathcal{A})$. Assume the contrary that $Y\notin\mathcal{S}(\mathcal{A})$. Then there exists $Z\in\mathcal{S}(\mathcal{A})$ such that $Z\subsetneq Y$. Recall that $\lambda_{X}=|Hx\cap X|$ does not depend on $x\in X$. We have $\lambda_{X}\neq 1$. Indeed, if $\lambda_{X}=1$, then $\lambda_{Y}=1$ and hence $|Y|=|Y_{D}|$. Together with $Z\subsetneq Y$, this implies that $Z_{D}\subsetneq Y_{D}$. Due to Eq. (28), we conclude that $Z_{D}\notin\mathcal{S}(\mathcal{A}_{D})$, a contradiction to Lemma 3.2(1).
Since $\lambda_{X}\neq 1$, we have $X=X_{H}\times X_{D}$ by Lemma 7.11 and $\lambda_{X}=|X_{H}|\geq 2$. So $Y=Y_{H}\times Y_{D}=X_{H}\times Y_{D}$. As $Z_{H},Y_{H}\in\mathcal{S}(\mathcal{A}_{H})$ and $Z_{H}\subseteq Y_{H}$, we have $Z_{H}=Y_{H}$. Besides, $Z_{D}=Y_{D}$ by Eq. (28). Since $Z\subsetneq Y$, we have $\lambda_{Z}=1$ by Lemma 7.11.
If $n=p^{k}$, then given $x\in X_{D}$, all the elements from $Y_{D}x$ lie in the same $\langle Y_{D}\rangle$-coset. So by Lemma 6.8(1), they lie in pairwise distinct basic sets of $\mathcal{A}_{D}$. Together with $|X_{D}|=|Y_{D}|$, this implies that there are at least $|X_{D}|$ pairwise distinct $T_{D}\in\mathcal{S}(\mathcal{A}_{D})$ with $c_{X_{D}Y_{D}}^{T_{D}}\geq 1$. Therefore there are at least $|X_{D}|$ pairwise distinct $T\in\mathcal{S}(\mathcal{A})$ with $c_{X_{D}Z}^{T}\geq 1$. Every such $T$ is rationally conjugate to $X$ (Lemma 2.3) and hence $|T|=|X|=|X_{H}||X_{D}|$. Thus, at least $|X_{H}||X_{D}|^{2}$ elements enter the element $\underline{X_{D}}\cdot\underline{Z}$. On the other hand, $|Z|=|Z_{D}|=|Y_{D}|=|X_{D}|$ because $\lambda_{Z}=1$ and consequently exactly $|X_{D}|^{2}$ elements enter the element $\underline{X_{D}}\cdot\underline{Z}$, a contradiction to $|X_{H}|=|K_{H}|\geq 2$.
If $n=pq$, then $X^{\pi}\cap Z^{\pi}\neq\varnothing$, where $\pi$ is the canonical epimorphism from $G$ to $G/R$. Since $X^{\pi},Z^{\pi}\in\mathcal{S}(\mathcal{A}_{G/R})$, we conclude that $X^{\pi}=Z^{\pi}$. On the other hand, $|X^{\pi}\cap H^{\pi}x|=|X_{H}|\geq 2$ for every $x\in X^{\pi}$, whereas $|Z^{\pi}\cap H^{\pi}z|=1$ for every $z\in Z^{\pi}$ by Lemma 2.1, a contradiction. ∎
Since $\mathcal{A}_{H}=\operatorname{Cyc}(K_{H},H)$, where $K_{H}\leq\operatorname{Aut}(H)$ is of order $2$ or $3$, and $\mathcal{A}_{D}=\operatorname{Cyc}(K_{D},D)$, one can choose a basic set $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$ such that
$|X_{H}|=|K_{H}|>1$
and $X_{D}$ consists of generators of $D$. Clearly, $X_{H}\in\operatorname{Orb}(K_{H},H)$ and $X_{D}\in\operatorname{Orb}(K_{D},D)$.
Claim 2.
Let $Y\in\mathcal{S}(\mathcal{A})\setminus(\mathcal{S}(\mathcal{A}_{H})\cup\mathcal{S}(\mathcal{A}_{D}))$. Then $Y=X^{(r)}$ for some $r\in\mathbb{Z}$ or $|Y_{H}|=1$ and $Y=Y_{H}\times Y_{D}$.
Proof.
Let $Y_{H}=X_{H}$. Since $Y\notin\mathcal{S}(\mathcal{A}_{H})$ by the condition of the claim and $X_{D}$ consists of generators of $D$, we conclude that $Y\cap X^{(r)}\neq\varnothing$ for some $r\in\mathbb{Z}$. Claim $1$ implies that $X^{(r)}\in\mathcal{S}(\mathcal{A})$ and hence $Y=X^{(r)}$ as required. Now let $Y_{H}\neq X_{H}$. Since $X_{H},Y_{H}\subseteq H^{\#}$ and $|X_{H}|>1$, we obtain $|Y_{H}|=1$ and hence $Y=Y_{H}\times Y_{D}$ as required. ∎
If $X=X_{H}\times X_{D}$, then $Y=Y_{H}\times Y_{D}$ for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$ by Claim $2$ and hence $\mathcal{A}=\mathcal{A}_{H}\otimes\mathcal{A}_{D}$ which contradicts to the supposition of the theorem. Therefore $X\neq X_{H}\times X_{D}$. Together with Lemma 7.11, this implies that
$\lambda_{X}=1.$
Further, we are going to construct $K\leq\operatorname{Aut}(G)$ such that
(29) $X\in\operatorname{Orb}(K,G).$
Since $X_{D}\in\operatorname{Orb}(K_{D},D)$, the set $X_{D}$ is regular. So $X_{h}=X_{h,p^{k}}$ for every $h\in X_{H}$. Lemma 2.5 implies that $X_{h}$ is an orbit of some $K^{0}_{D}\leq\operatorname{Aut}(D)$ for every $h\in H$. Clearly, $K^{0}_{D}\leq K_{D}$, $K_{D}$ acts transitively on the set $\{X_{h}:~h\in X_{H}\}=\operatorname{Orb}(K^{0}_{D},X_{D})$, and $K^{0}_{D}$ is the kernel of this action. As $\lambda_{X}=1$, we have $|K_{D}:K^{0}_{D}|=|X_{H}|=|K_{H}|\leq 3$. Therefore $K_{H}\cong K_{D}/K^{0}_{D}\cong C_{|X_{H}|}$. Let $\sigma_{0}$ be a generator of $K_{H}$. Since $|K_{D}:K^{0}_{D}|=|K_{H}|\leq 3$, there is $\tau_{0}\in K_{D}$ such that
(30) $X_{h}^{\tau_{0}}=X_{h^{\sigma_{0}}}$
for every $h\in X_{H}$. Let $\psi$ be an isomorphism from $K_{H}$ to $K_{D}/K^{0}_{D}$ such that $\sigma_{0}^{\psi}=K^{0}_{D}\tau_{0}$ and
$K=\{(\sigma,\tau)\in K_{H}\times K_{D}:~(\sigma)^{\psi}=K^{0}_{D}\tau\}.$
Eq. (30) yields that $X$ is $K$-invariant. As $K_{H}$ and $K_{D}$ are transitive on $X_{H}$ and $X_{D}$, respectively, so is $K$ on $X$. Thus, Eq. (29) holds.
To prove that $\mathcal{A}$ is cyclotomic, it is enough to show that $\mathcal{A}=\operatorname{Cyc}(K,G)$. One can see that $\mathcal{A}_{H}=\operatorname{Cyc}(K,G)_{H}$ ($\mathcal{A}_{D}=\operatorname{Cyc}(K,G)_{D}$, respectively) because $K^{H}=K_{H}$ ($K^{D}=K_{D}$, respectively). Let $Y\in\mathcal{S}(\mathcal{A})\setminus(\mathcal{S}(\mathcal{A}_{H})\cup\mathcal{S}(\mathcal{A}_{D}))$. Then $Y=X^{(r)}$ for some $r\in\mathbb{Z}$ or $|Y_{H}|=1$ and $Y=Y_{H}\times Y_{D}$ by Claim $2$. In the former case, $Y\in\operatorname{Orb}(K,G)$ by Eq. (29), whereas in the latter one, by the definition of $K$. Thus, every basic set of $\mathcal{A}$ is an orbit of $K$, i.e. $\mathcal{A}=\operatorname{Cyc}(K,G)$ as desired.
Now let us prove that $\mathcal{A}$ is $2$-minimal. Let $x\in X$. It is enough to show that the pointwise stabilizer $\operatorname{Aut}(\mathcal{A})_{e,x}$ is trivial. Indeed, then $X$ is a faithful regular orbit of $\operatorname{Aut}(\mathcal{A})_{e}$ and $\mathcal{A}$ is $2$-minimal by [23, Lemma 8.2]. It suffices to verify that every $f\in\operatorname{Aut}(\mathcal{A})_{e,x}$ is trivial.
Claim 3.
If $y^{f}=y$ for some $y\in G\setminus H$, then $(Hy)^{f}=Hy$ and $f^{Hy}=\operatorname{id}_{Hy}$.
Proof.
As $H$ is an $\mathcal{A}$-subgroup, each $H$-coset is a block of $\operatorname{Aut}(\mathcal{A})$. One can see that $y=y^{f}\in Hy\cap(Hy)^{f}$. So $(Hy)^{f}=Hy$. Since $\mathcal{A}=\operatorname{Cyc}(K,G)$, we have $\lambda_{Y}=1$ for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus(H\cup D)}$. Therefore all the elements of $Hy$ lie in pairwise distinct basic sets. Together with $(Hy)^{f}=Hy$, this implies that $f^{Hy}=\operatorname{id}_{Hy}$. ∎
The element $x$ can be uniquely presented in the form $x=hx_{0}$, where $h\in H^{\#}$ and $x_{0}\in D$ is a generator of $D$. From Claim $3$ it follows that $x_{0}^{f}=x_{0}$. Observe that $D^{f}=D$ because $D$ is an $\mathcal{A}$-subgroup and hence $f^{D}\in\operatorname{Aut}(\mathcal{A}_{D})$. Lemma 6.5 yields that $\operatorname{Aut}(\mathcal{A}_{D})=G_{r}\rtimes K_{D}\leq\operatorname{Hol}(G)$. So $f^{D}\in K_{D}\leq\operatorname{Aut}(D)$. Together with $x_{0}^{f}=x_{0}$, this implies that $f^{D}=\operatorname{id}_{D}$. Therefore $f$ fixes an element from every $H$-coset. Thus, $f$ is trivial by Claim $3$.
Since $\mathcal{A}=\operatorname{Cyc}(K,G)$ and $\mathcal{A}$ is $2$-minimal, we have $\operatorname{Aut}(\mathcal{A})=G_{r}\rtimes K$. This implies that $\mathcal{A}$ is normal and $|\operatorname{Aut}(\mathcal{A})^{H}|=|H||K^{H}|\leq 12$ which completes the proof of the theorem. ∎
Theorem 7.13.
Let $\mathcal{A}$ be dense. Suppose that $\mathcal{A}_{D}$ is the nontrivial $U/L$-wreath product, where $L$ is a minimal nontrivial $\mathcal{A}_{D}$-subgroup and $U$ is an $\mathcal{A}_{D}$-subgroup with $|\operatorname{rad}(\mathcal{A}_{U})|=1$. Then one of the following statements holds:
  1. (1)
    $\mathcal{A}$ is the nontrivial $(H\times U)/L$-wreath product;
  2. (2)
    $n=3^{k}$, $\mathcal{A}$ is the $(H\times W)/L$-wreath product (possibly, trivial) for an $\mathcal{A}_{D}$-subgroup $W>U$ such that $|\operatorname{rad}(\mathcal{A}_{W})|=3$ and $\mathcal{A}_{H\times W}$ is cyclotomic.
Proof.
At first, suppose that $n=pq$. As $\mathcal{A}_{D}$ is the nontrivial $U/L$-wreath product, we may assume without loss of generality that $|U|=|L|=p$. Then $H$ is a maximal $\mathcal{A}$-subgroup whose order is not divisible by $p$. So $\mathcal{A}$ is the nontrivial $(H\times U)/L$-wreath product by Lemma 5.3 applied to $H_{1}=H$ and $P_{1}=L$ and Statement $(1)$ of the theorem holds as desired.
Now suppose that $n=p^{k}$. Since $\mathcal{A}_{D}$ is the nontrivial $U/L$-wreath product, we have $k\geq 2$.
Claim 4.
If there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times U)}$ such that $L\nleq\operatorname{rad}(X)$, then $n=3^{k}$, $\mathcal{A}_{H}=\mathcal{T}_{H}$, $L=P$, and every such $X$ is equal to
(31) $h_{1}\{x\}\cup h_{2}\{xx_{0}\}\cup h_{3}\{xx_{0}^{2}\}~\text{or}~h_{1}\{x,x^{-1}\}\cup h_{2}\{xx_{0},x^{-1}x_{0}^{2}\}\cup h_{3}\{xx_{0}^{2},x^{-1}x_{0}\},$
where $h_{1},h_{2},h_{3}$ are pairwise distinct nonidentity elements of $H$, $x_{0}$ is a generator of $P$, and $x\in D$. In particular, $|X|=|X_{D}|\leq 6$.
Proof.
We have $X\neq X_{H}\times X_{D}$ because otherwise $L\leq\operatorname{rad}(X_{D})$ and hence $L\leq\operatorname{rad}(X)$ that contradicts to the assumption of the lemma. Hence $|X_{H}|\in\{2,3\}$ by Lemma 7.9 and
$\lambda_{X}=|X\cap Hx|=1~\text{for every}~x\in X$
by Lemma 7.11. The number $\mu_{X}=|X\cap Lx|$ does not depend on $x\in X$ by Lemma 2.1. Since $\lambda_{X}=1$, we conclude that $\mu_{X}=|L|$ or $\mu_{X}=|L|/|X_{H}|$. In the former case, $L\leq\operatorname{rad}(X)$, a contradiction to the assumption of the claim. In the latter one, since $|X_{H}|\in\{2,3\}$ and $|L|$ is a $p$-power for an odd prime $p$, we obtain $|X_{H}|=p=3$. This implies that $n=3^{k}$, $\mathcal{A}_{H}=\mathcal{T}_{H}$, and $X_{H}=H^{\#}$.
Let $X_{D}$ be regular. Then all the $X_{h}$, $h\in H^{\#}$, are orbits of some $K\leq\operatorname{Aut}(\langle X_{D}\rangle)$ by Lemma 2.5. Since $L\nleq\operatorname{rad}(X)$, we obtain $P\nleq\operatorname{rad}(X_{h})$ and hence $|\operatorname{rad}(X_{h})|=1$ for every $h\in H^{\#}$. Lemma 6.8(1) implies that $|K|=|X_{h}|\leq 2$ for every $h\in H^{\#}$. Therefore each $X_{h}$ is a singleton or each $X_{h}$ is of the form $\{x,x^{-1}\}$, $x\in D$. Due to $\lambda_{X}=1$, all the $X_{h}$ are pairwise disjoint and consequently
$\sum\limits_{h\in H^{\#}}|X_{h}|=|X_{D}|\in\{3,6\}.$
As $L\leq\operatorname{rad}(X_{D})$, we have $L=\operatorname{rad}(X_{D})=P$ and hence $X_{D}=Px$ or $X_{D}=Px\cup Px^{-1}$ for some $x\in D$. In the former case, each $X_{h}$ is a singleton and hence $X$ is of the first form from Eq. (31), whereas in the latter one, each $X_{h}$ consists of two mutually inverse elements and hence $X$ is of the second form from Eq. (31) as desired.
Now let $X_{D}$ be nonregular. Then $X_{D}=V\setminus N$ for some $\mathcal{A}_{D}$-subgroups $V>N\geq U$ such that $|V:N|\geq 9$ by Lemma 6.11. Let $m$ be the minimum of orders of the elements from $X_{D}$. If $P\nleq\operatorname{rad}(X_{h,l})$ for some $h\in H^{\#}$ and $l>m$, then
$\varnothing\neq X^{[p]}_{D}\subsetneq X_{D}.$
However, $X^{[p]}$ is an $\mathcal{A}$-set by Lemma 2.4 and hence $X^{[p]}_{D}$ is a union of some basic sets of $\mathcal{A}_{D}$ by Lemma 3.2(1), a contradiction to the above inclusion. Therefore
(32) $P\leq\operatorname{rad}(X_{h,l})$
for all $h\in H^{\#}$ and $l>m$. Due to Lemma 2.5, all $X_{h,m}$, $h\in H^{\#}$, are orbits of some $K\leq\operatorname{Aut}(D_{m})$. As $L\nleq\operatorname{rad}(X)$, Eq. (32) implies that $P\nleq\operatorname{rad}(X_{h,m})$ and hence $|\operatorname{rad}(X_{h,m})|=1$ for every $h\in H^{\#}$. From Lemma 6.8(1) it follows that $|X_{h,m}|\leq 2$ for every $h\in H^{\#}$. Together with $X_{D}=V\setminus N$, this yields that
$3^{m}-3^{m-1}=|D_{m}^{*}|=|\bigcup\limits_{h\in H^{\#}}X_{h,m}|\leq 6.$
Therefore $m\leq 2$. If $m=1$, then $U$ is trivial. However, this contradicts to the assumption of the theorem that $\mathcal{A}$ is the nontrivial $U/L$-wreath product. If $m=2$, then $U=L=P$ is an $\mathcal{A}$-subgroup. By Lemma 2.1 and Eq. (32), we obtain $P=L\leq\operatorname{rad}(X)$, a contradiction to the assumption of the claim. ∎
If $L\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times U)}$, then $\mathcal{A}$ is the nontrivial $(H\times U)/L$-wreath product and Statement $(1)$ of the theorem holds.
Suppose that $L\nleq\operatorname{rad}(X)$ for some $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times U)}$. As $L$ is a minimal nontrivial $\mathcal{A}_{D}$-subgroup, we obtain $P\nleq\operatorname{rad}(X)$. So the set
$\mathcal{W}=\{X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times U)}:~P\nleq\operatorname{rad}(X)\}$
is nonempty. Put
$W=\langle X_{D}:~X\in\mathcal{W}\rangle.$
Observe that $X_{D}$ is an $\mathcal{A}_{D}$-set for every $X\in\mathcal{W}$ by Lemma 3.2(1) and hence $W$ is an $\mathcal{A}_{D}$-subgroup. By the definition, $W>U$. Claim $4$ implies that $n=3^{k}$, $\mathcal{A}_{H}=\mathcal{T}_{H}$, $L=P$, and $|X_{D}|\leq 6$ for every $X\in\mathcal{W}$. Together with $\operatorname{rad}(X_{D})\geq L$, the latter yields that $\operatorname{rad}(X_{D})=P$ for every $X\in\mathcal{W}$ and consequently $\operatorname{rad}(\mathcal{A}_{W})=P$. By the definition of $W$, the radical of every basic set outside $H\times W$ contains $P$. Therefore $\mathcal{A}$ is the $(H\times W)/P$-wreath product (possibly, trivial).
In view of the above paragraph, to prove Statement $(2)$ of the theorem, it suffices to show that the $S$-ring $\mathcal{A}_{H\times W}$ is cyclotomic. Let $X\in\mathcal{W}$ (recall that $\mathcal{W}$ is nonempty) be such that $X_{D}$ contains a generator of $W$. By Claim $4$, the set $X$ is one of the forms from Eq. (31), where $x\in W$ is a generator of $W$. Let $\sigma\in\operatorname{Aut}(H)$ be such that $\sigma=(h_{1}h_{2}h_{3})$, and let $\tau_{1},\tau_{2}\in\operatorname{Aut}(W)$ be such that $x^{\tau_{1}}=xx_{0}$ and $x^{\tau_{2}}=x^{-1}$. Put
$K_{1}=\langle\varphi\rangle\leq\operatorname{Aut}(H\times W)~\text{and}~K_{2}=K_{1}\times\langle\psi\rangle\leq\operatorname{Aut}(H\times W),$
where $\varphi,\psi\in\operatorname{Aut}(H\times W)$ are such that $\varphi^{H}=\sigma$, $\varphi^{D}=\tau_{1}$, $\psi^{H}=\operatorname{id}_{H}$, $\psi^{D}=\tau_{2}$. Clearly, $K_{1}\cong C_{3}$ and $K_{2}\cong C_{6}$. By the definitions of $K_{1}$ and $K_{2}$, we have $X\in\operatorname{Orb}(K,H\times W)$ and $|X|=|K|$, where $K=K_{1}$ if $X$ is of the first form from Eq. (31) and $K=K_{2}$ otherwise. It is easy to see that $H^{\#},X_{D}\in\operatorname{Orb}(K,H\times W)$. Put $X^{0}=X$ and $X^{i}=(X^{i-1})^{[p]}$ for $i\in\{1,\ldots,l-1\}$, where $l$ is such that $|W|=3^{l}$. Eq. (31) implies that $X^{i}=H^{\#}x^{p^{i}}$ or $X^{i}=H^{\#}x^{p^{i}}\cup H^{\#}x^{-p^{i}}$ for $i\in\{1,\ldots,l-1\}$. Each $X^{i}$ is an $\mathcal{A}$-set by Lemma 2.4. Moreover, using Lemma 2.1 and Lemma 3.2(1), it is easy to verify that each $X^{i}$ and $X^{i}_{D}$ are exactly basic sets of $\mathcal{A}_{H\times W}$ and $\mathcal{A}_{W}$, respectively. Due to the definition of $K$, each $X^{i}$ and $X^{i}_{D}$ are orbits of $K$. Every $Y\in\mathcal{S}(\mathcal{A}_{H\times W})_{G\setminus H}$ is rationally conjugate to $X^{i}$ or $X^{i}_{D}$ for some $i\in\{0,\ldots,l-1\}$ by Lemma 2.3 and hence $Y\in\operatorname{Orb}(K,H\times W)$. Thus, every basic set of $\mathcal{A}_{H\times W}$ is an orbit of $K$, i.e. $\mathcal{A}_{H\times W}=\operatorname{Cyc}(K,G)$ as required. ∎
Lemma 7.14.
Suppose that $\mathcal{A}$ is schurian, $\mathcal{A}_{H}=\mathcal{T}_{H}$, and $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$. Then $|\operatorname{Aut}(\mathcal{A})^{H}|=12$.
Proof.
One can see that $\operatorname{Aut}(\mathcal{A})^{H}\approx_{2}\operatorname{Aut}(\mathcal{A}_{H})\cong\operatorname{Sym}(4)$ because $\mathcal{A}$ is schurian. Moreover, $|\operatorname{Aut}(\mathcal{A})^{H}|\in\{12,24\}$ by Remark 2.7.
Since $\mathcal{A}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{D}$, there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus(U\cup D)}$ such that $X\neq X_{H}\times X_{D}$. Lemma 7.11 implies that $\lambda_{X}=1$. From Lemma 3.2(1) it follows that $X_{D}\in\mathcal{S}(\mathcal{A}_{D})$ and $X_{H}=H^{\#}$. Let $L=\operatorname{rad}(X)$, $U=\langle X\rangle$, and $W=\langle X_{D}\rangle$. Since $D$ is an $\mathcal{A}$-subgroup, we have $L\leq D$. The image of $T\subseteq G$ under the canonical epimorphism from $G$ to $G/L$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muT\mkern-1.5mu}\mkern 1.5mu$. The argument from the previous paragraph applied to $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$ implies that $|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|\in\{12,24\}$. To prove the lemma, it is enough to show that
$|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=12.$
Indeed, then
$|\operatorname{Aut}(\mathcal{A})^{H}|=|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|\leq|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=12$
as required, where the first equality holds because $|L\cap H|=1$ and hence the mapping $f\mapsto\mkern 1.5mu\overline{\mkern-1.5muf\mkern-1.5mu}\mkern 1.5mu$ from $\operatorname{Aut}(\mathcal{A})^{H}$ to $\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}$, where $\mkern 1.5mu\overline{\mkern-1.5muf\mkern-1.5mu}\mkern 1.5mu$ is the permutation induced by $f$ on $\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu$, is an isomorphism.
By the definition of $L$, we have
(33) $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)|=1.$
As $\lambda_{X}=1$ and $L\leq D$, we obtain $\lambda_{\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu}=|\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\cap\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu|=1$ for every $\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu\in\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu$. So $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}\neq\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}\otimes\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muW\mkern-1.5mu}\mkern 1.5mu}$. The set $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D}$ is a basic set of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muD\mkern-1.5mu}\mkern 1.5mu}$. If $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D})|=1$, then $|\operatorname{rad}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muW\mkern-1.5mu}\mkern 1.5mu})|=1$ and Theorem 7.12(2) applied to $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$ yields that $|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=12$ as desired.
Suppose that $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D})|>1$. In view of Lemma 6.12 if $|W|=pq$ and Lemma 6.10 if $|W|$ is a $p$-power, Theorem 7.13 holds for $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$. If Statement $(1)$ of Theorem 7.13 holds for $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$, then $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)|>1$, a contradiction to Eq. (33). Therefore Statement $(2)$ of Theorem 7.13 holds for $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$. Due to Claim $4$ and Eq. (33), the basic sets of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu}$ inside $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{H}\times\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D}$ are of the form
$\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{1}\cup\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{2}\cup\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{3},~\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{2}\cup\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{3}\cup\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{1},~\mkern 1.5mu\overline{\mkern-1.5muZ\mkern-1.5mu}\mkern 1.5mu=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{3}\cup\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{1}\cup\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{2},$
where the $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{i}$ are pairwise distinct elements of $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{H}$ and the $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{i}$ are pairwise disjoint subsets of $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{D}$ of size at most $2$. If $|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=24$, then there is $f\in\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})$ such that $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}^{f}=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}$, $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}^{f}=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}$, and $\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}^{f}=\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}$. Eq. (2) implies that
$\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{1}^{f}=(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}\cap\mkern 1.5mu\overline{\mkern-1.5muD\mkern-1.5mu}\mkern 1.5mu)^{f}=\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{1}^{f}\cap\mkern 1.5mu\overline{\mkern-1.5muD\mkern-1.5mu}\mkern 1.5mu=\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{1}$
and
$\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{1}^{f}=(\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}\cap\mkern 1.5mu\overline{\mkern-1.5muD\mkern-1.5mu}\mkern 1.5mu)^{f}=\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{3}^{f}\cap\mkern 1.5mu\overline{\mkern-1.5muD\mkern-1.5mu}\mkern 1.5mu=\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5muh\mkern-1.5mu}\mkern 1.5mu_{2}\cap\mkern 1.5mu\overline{\mkern-1.5muD\mkern-1.5mu}\mkern 1.5mu=\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu_{3},$
a contradiction. Thus, $|\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muU\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muH\mkern-1.5mu}\mkern 1.5mu}|=12$ and we are done. ∎
7.4. Proof of Theorem 7.1
Suppose that $H$ or $D$ is not an $\mathcal{A}$-subgroup. Then Theorem 7.1 holds for $\mathcal{A}$ or $\mathcal{A}$ is a nontrivial $S$-wreath product for an $\mathcal{A}$-section $S=U/L$ such that $4\nmid|W|$ and $|\operatorname{rad}(\mathcal{A}_{W})|=1$, where $W=U$ if $H$ is not an $\mathcal{A}$-subgroup and $W=G/L$ if $D$ is not an $\mathcal{A}$-subgroup: indeed, this follows from Theorem 7.3 in the former case and from Theorem 7.8 in the latter one.
Note that $|W_{H}|\leq 2$, where $W_{H}$ is the Sylow $2$-subgroup of $W$, because $|W|$ is not divisible by $4$ and consequently $\mathcal{A}_{W_{H}}=\mathbb{Z}W_{H}$. So Lemma 3.2(2) yields that
(34) $\mathcal{A}_{W}=\mathcal{A}_{W_{H}}\otimes\mathcal{A}_{W_{D}},$
where $W_{D}$ is the Hall $2^{\prime}$-subgroup of $W$ (here $W_{H}$ and $W_{D}$ can be trivial). Since $|\operatorname{rad}(\mathcal{A}_{W})|=1$, we conclude that $|\operatorname{rad}(\mathcal{A}_{W_{D}})|=1$.
If $\mathcal{A}_{W_{D}}=\mathcal{T}_{W_{D}}$, then due to Eq. (34), we obtain $L\in\{W_{H},W_{D},W\}$ whenever $W=U$ and $S\in\{W_{H},W_{D},\{L\}\}$ whenever $W=G/L$. This yields that Statement $(1)$ or $(2)$ from Theorem 7.1 holds. If $\mathcal{A}_{W_{D}}\neq\mathcal{T}_{W_{D}}$, then $\mathcal{A}_{W_{D}}=\operatorname{Cyc}(K_{D},W_{D})$ for some $K_{D}\leq\operatorname{Aut}(W_{D})$ by Lemma 6.2. As $|W_{H}|\leq 2$, we have $\mathcal{A}_{W_{H}}=\operatorname{Cyc}(K_{H},W_{H})$, where $K_{H}$ is trivial. Therefore $\mathcal{A}_{W}=\operatorname{Cyc}(K_{H}\times K_{D},W)$ by Eqs. (4) and (34) and Statement $(3)$ from Theorem 7.1 holds.
Now suppose that $H$ and $D$ are $\mathcal{A}$-subgroups. If $|\operatorname{rad}(\mathcal{A}_{D})|=1$, then $\mathcal{A}=\mathcal{A}_{H}\otimes\mathcal{A}_{D}$ or $\mathcal{A}$ is cyclotomic by Theorem 7.12(1) and we are done. If $|\operatorname{rad}(\mathcal{A}_{D})|>1$, then $\mathcal{A}_{D}$ is a nontrivial $U/L$-wreath product for some $\mathcal{A}_{D}$-section $U/L$ such that $L$ is a minimal nontrivial $\mathcal{A}_{D}$-subgroup and $|\operatorname{rad}(\mathcal{A}_{U})|=1$. Indeed, this follows from Lemma 6.12 if $n=pq$ and from Lemma 6.10 if $n=p^{k}$. So Theorem 7.13 holds for $\mathcal{A}$. If Statement $(2)$ of Theorem 7.13 holds for $\mathcal{A}$, then $n=3^{k}$, $\mathcal{A}$ is the $S=(H\times W)/L$-wreath product (possibly, trivial) for an $\mathcal{A}_{D}$-subgroup $W>U$ such that $|\operatorname{rad}(\mathcal{A}_{W})|=3$, and $\mathcal{A}_{H\times W}$ is cyclotomic. In this case, $H\times W=G$ and $\mathcal{A}$ is cyclotomic or $H\times W<G$, $\mathcal{A}$ is the nontrivial $S$-wreath product. Thus, Statement $(4)$ from Theorem 7.1 holds.
Suppose that Statement $(1)$ of Theorem 7.13 holds for $\mathcal{A}$, i.e. $\mathcal{A}$ is the nontrivial $S=V/L$-wreath product, where $V=H\times U$. If $\mathcal{A}_{V}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{U}$, then $\mathcal{A}_{V}$ is cyclotomic by Theorem 7.12(1). So Statement $(4)$ from Theorem 7.1 holds for $\mathcal{A}$ and $S$. Further, we assume that
(35) $\mathcal{A}_{V}=\mathcal{A}_{H}\otimes\mathcal{A}_{U}.$
Then $\mathcal{A}_{H}=\operatorname{Cyc}(K_{H},H)$ for some $K_{H}\leq\operatorname{Aut}(H)$ by Lemma 7.9. If $\mathcal{A}_{U}=\operatorname{Cyc}(K_{U},U)$ for some $K_{U}\leq\operatorname{Aut}(U)$, then $\mathcal{A}_{V}=\operatorname{Cyc}(K_{H}\times K_{U},V)$ by Eqs. (4) and (35). Therefore Statement $(4)$ from Theorem 7.1 holds for $\mathcal{A}$ and $S$. Otherwise, $\mathcal{A}_{U}=\mathcal{T}_{U}$ by Lemma 6.2. Then $L=U$. Together with Eq. (35), this implies that $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{V}$. Thus, Statement $(2)$ from Theorem 7.1 holds for $S$ and we are done.
7.5. Auxiliary statements
Lemma 7.15.
Let $\mathcal{B}$ be an $S$-ring over a subgroup $V$ of $E_{4}\times C_{p^{k}}$, where $p\geq 5$ and $k\geq 2$. Suppose that $p^{2}~|~|V|$, $\mathcal{B}$ is indecomposable, and there is a $\mathcal{B}$-subgroup of order or index $p$. Then $\mathcal{B}$ is normal.
Proof.
Clearly, $\mathcal{B}$ is nontrivial. Firstly, suppose that $V$ is cyclic. If $|V|$ is a $p$-power, then $\mathcal{B}$ is normal by the supposition of this lemma and Lemma 6.2. If $|V|=2p^{l}$ for some $l\geq 2$ and $\mathcal{B}$ is not normal, then $\mathcal{B}=\mathcal{T}_{V_{1}}\otimes\mathbb{Z}V_{2}$, where $V_{1},V_{2}\leq V$ with $|V_{1}|=p^{l}$ and $|V_{2}|=2$, by Lemma 6.13. However, the latter equality contradicts to the supposition of the lemma that there is a $\mathcal{B}$-subgroup of order or index $p$.
Now suppose that $V$ is noncyclic, i.e. 4||V|4~|~|V|. Let
$\mathcal{B}=\mathcal{B}_{V_{1}}\otimes\mathcal{B}_{V_{2}}$
for some proper nontrivial $\mathcal{B}$-subgroups $V_{1}$ and $V_{2}$ such that $V=V_{1}\times V_{2}$. Without loss of generality, we may assume that $|V_{1}|=p^{l}$ and $|V_{2}|=4$, or $|V_{1}|=2p^{l}$ and $|V_{2}|=2$ for some $l\geq 2$. In particular, $V_{1}$ is cyclic. Note that $|\operatorname{rad}(\mathcal{B}_{V_{1}})|=1$ because otherwise $\mathcal{B}_{V_{1}}$ is decomposable by Lemma 6.2 and hence so is $\mathcal{B}$, which contradicts to the supposition of the lemma. If $|V_{1}|=p^{l}$ ($|V_{2}|=2p^{l}$, respectively), then the last condition of this lemma and Lemma 6.2 (Lemma 6.13, respectively) imply that $\mathcal{B}_{V_{1}}$ is normal. Then $\mathcal{B}_{V_{2}}$ is also normal. Indeed, this is clear if $|V_{2}|=2$ and follows from Lemma 7.9 if $|V_{2}|=4$. Therefore $\mathcal{B}$ is normal by Eq. (3).
In view of the above paragraph, we may assume that $\mathcal{B}$ is not a nontrivial tensor product. Since $\mathcal{B}$ is also nontrivial and indecomposable, Theorems 7.3 and 7.8 can not hold for $\mathcal{B}$. So $\mathcal{B}$ is dense. In particular, the Sylow $p$-subgroup $V_{1}$ of $S$ is a $\mathcal{B}$-subgroup. If $|\operatorname{rad}(\mathcal{B}_{V_{1}})|>1$, then $\mathcal{B}$ is decomposable or $p=3$ by Theorem 7.13, a contradiction to the supposition of the lemma. Therefore $|\operatorname{rad}(\mathcal{B}_{V_{1}})|=1$. Thus, $\mathcal{B}$ is normal by Theorem 7.12(1). ∎
Recall that $\mathcal{A}$ is an $S$-ring over $G=H\times D$, where $H\cong E_{4}$ and $D\cong C_{n}$ for $n\in\{pq,p^{k}\}$.
Lemma 7.16.
Let $n=p^{k}$. The following statements hold.
  1. (1)
    If $P\cong C_{p}\leq G$ is an $\mathcal{A}$-subgroup, then $\operatorname{Aut}(\mathcal{A})^{P}\leq\operatorname{Hol}(P)$ or $\mathcal{A}$ is the $U/P$-wreath product (possibly, trivial) for an $\mathcal{A}$-subgroup $U$ such that $|U/P|\leq 4$.
  2. (2)
    If $U\cong E_{4}\times C_{p^{k-1}}\leq G$ is an $\mathcal{A}$-subgroup, then $\operatorname{Aut}(\mathcal{A})^{G/U}\leq\operatorname{Hol}(G/U)$ or $\mathcal{A}$ is the $U/L$-wreath product (possibly, trivial) for an $\mathcal{A}$-subgroup $L$ such that $|U/L|\leq 4$.
Proof.
If $p=3$, then $\operatorname{Hol}(C_{3})\cong\operatorname{Sym}(3)$ and the lemma is clear. Let $p\geq 5$. Let us prove Statement $(1)$. If $P\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times P)}$, then $\mathcal{A}$ is the $U/P$-wreath product for an $\mathcal{A}$-subgroup $U\leq H\times P$ and we are done. Further, we assume that there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus(H\times P)}$ such that
$\operatorname{rad}(X)\ngeq P.$
Let $R=\operatorname{rad}(X)$ and $V=\langle X\rangle$. The image of $Y\subseteq G$ under the canonical epimorphism from $G$ to $G/R$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu$. As $R\ngeq P$, we have $R\leq H$ and $|R\cap P|=1$ (note that $R$ can be trivial). The latter equality implies that the kernel of the action of $\operatorname{Aut}(\mathcal{A})^{P}$ on $\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu$ is trivial and hence it suffices to show that
$\operatorname{Aut}(\mathcal{A})^{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}\leq\operatorname{Hol}(\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu).$
Since $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu$ generates $\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu$ and $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)|=1$, the $S$-ring $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$ is indecomposable. Clearly, $\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu$ is an $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$-subgroup of order $p$. Moreover, $p^{2}~|~|\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu|$ because $X\nsubseteq H\times P$ and $R\leq H$. So $\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu$ satisfies all the conditions of Lemma 7.15 and consequently $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$ is normal. Therefore
$\operatorname{Aut}(\mathcal{A})^{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}\leq\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu})^{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}\leq(\operatorname{Hol}(\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu))^{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}\leq\operatorname{Hol}(\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu),$
where the second inequality follows from the normality of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$, and we are done.
Now let us prove Statement $(2)$. If $U\cap D\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus U}$, then $\mathcal{A}$ is the $U/L$-wreath product for an $\mathcal{A}$-subgroup $L\geq(U\cap D)$ and we are done. Further, we assume that there is $X\in\mathcal{S}(\mathcal{A})_{G\setminus U}$ such that
$\operatorname{rad}(X)\ngeq(U\cap D).$
Let $V=\langle X\rangle$. Clearly, $|G:V|\in\{1,2,4\}$ and $|V:(V\cap U)|=|G:U|=p$. Observe that $\operatorname{Aut}(\mathcal{A})^{G/U}$ and $\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}$ are permutation isomorphic in the sense of [4, p. 17]. Indeed, the kernel of the action of $\operatorname{Aut}(\mathcal{A})$ on $G/U$ contains $H_{r}$ and hence for every $f\in\operatorname{Aut}(\mathcal{A})$, there is $f^{\prime}\in\operatorname{Aut}(\mathcal{A})$ such that $f^{G/U}=(f^{\prime})^{G/U}$ and $V^{f^{\prime}}=V$. As $V\cap U\leq U$, for every $\Delta\in G/U$, there is $\Delta^{\prime}\in V/(V\cap U)$ such that $\Delta^{\prime}\subseteq\Delta$. Moreover, such $\Delta^{\prime}$ is unique because $|V:(V\cap U)|=|G:U|=p$. Therefore the mappings $f^{G/U}\mapsto(f^{\prime})^{V/(V\cap U)}$ from $\operatorname{Aut}(\mathcal{A})^{G/U}$ to $\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}$ and $\Delta\mapsto\Delta^{\prime}$ from $G/U$ to $V/(V\cap U)$ define a permutation isomorphism from $\operatorname{Aut}(\mathcal{A})^{G/U}$ to $\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}$. Thus, it is enough to show that
$\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}\leq\operatorname{Hol}(V/(V\cap U)).$
Let $R=\operatorname{rad}(X)$. Again, the image of $Y\subseteq G$ under the canonical epimorphism from $G$ to $G/R$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu$. Clearly, the $S$-ring $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$ is indecomposable and $\mkern 1.5mu\overline{\mkern-1.5muV\cap U\mkern-1.5mu}\mkern 1.5mu$ is an $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$-subgroup of index $p$. Since $R\ngeq(U\cap D)$, we conclude that $p^{2}~|~|\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu|$. So $\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu$ satisfies all the conditions of Lemma 7.15 and hence $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$ is normal. Therefore
$\operatorname{Aut}(\mathcal{A})^{V/(V\cap U)}\leq\operatorname{Aut}(\mathcal{A}^{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu})^{V/(V\cap U)}\leq\operatorname{Aut}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu})^{V/(V\cap U)}\leq$
$\leq(\operatorname{Hol}(\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu))^{V/(V\cap U)}\leq\operatorname{Hol}(V/(V\cap U)),$
where the third inequality follows from the normality of $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muV\mkern-1.5mu}\mkern 1.5mu}$, and we are done. ∎
It should be mentioned that if $\mathcal{A}_{P}$ ($\mathcal{A}_{G/U}$, respectively) in Statement $(1)$ (Statement $(2)$, respectively) of Lemma 7.16 is nontrivial, then $\mathcal{A}_{P}$ ($\mathcal{A}_{G/U}$, respectively) is normal by Lemma 6.2 because $P$ ($G/U$, respectively) is of prime order $p$. Therefore the claim of Lemma 7.16 easily follows from the definition of a normal $S$-ring. We also note that if the generalized wreath product in Statement $(1)$ or $(2)$ of Lemma 7.16 is trivial, then $|G|=4p$.
8. Proof of Theorem 1.2
Throughout this section, we keep the notation from Section 7. We start with a lemma concerned with a schurity of proper subgroups of $G$.
Lemma 8.1.
Every proper subgroup of $G$ is Schur.
Proof.
Let $N<G$. If $N$ is cyclic, then $N$ is Schur by [10, Theorem 1.1], whereas if $N\cong E_{4}$, then $N$ is Schur by [11, Theorem 1.2]. Otherwise, $N\cong E_{4}\times C_{r^{m}}$, where $r\in\{p,q\}$ and $k>m\geq 1$, is Schur by induction on $m$ whose base follows from [11, Theorem 1.5]. ∎
Let $\mathcal{A}$ be a nontrivial $S$-ring over $G$. Let us prove that $\mathcal{A}$ is schurian. Theorem 7.1 holds for $\mathcal{A}$. We are done if $\mathcal{A}$ is cyclotomic. If $\mathcal{A}$ is a nontrivial tensor product, then $\mathcal{A}$ is schurian by Lemma 3.1 and Lemma 8.1. So we may assume that $\mathcal{A}$ is a nontrivial $S$-wreath product for some $\mathcal{A}$-section $S=U/L$ and one of Statements (1)(1)-$(4)$ from Theorem 7.1 holds. The $S$-rings $\mathcal{A}_{U}$ and $\mathcal{A}_{G/L}$ are schurian by Lemma 8.1. If Statement $(1)$ holds, i.e. $|S|\leq 2$, then $\mathcal{A}_{S}$ is $2$-minimal by Lemma 2.6 and hence $\mathcal{A}$ is schurian by Lemma 3.6. If Statement $(2)$ holds, i.e. $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{U}$ or $\mathcal{A}_{S}$ is $\otimes$-complemented in $\mathcal{A}_{G/L}$, then $\mathcal{A}$ is schurian by Lemma 3.7.
Now suppose that Statement $(3)$ or $(4)$ holds. Then one can choose $W\in\{U,G/L\}$ such that $\mathcal{A}_{W}$ is cyclotomic. This yields that $\mathcal{A}_{S}$ is also cyclotomic. Let $S_{H}$ and $S_{D}$ be the Sylow $2$-subgroup and the Hall $2^{\prime}$-subgroup of $S$, respectively. Note that $S_{H}$ and $S_{D}$ are $\mathcal{A}_{S}$-subgroups because $\mathcal{A}_{S}$ is cyclotomic. If $n=pq$ and $|S_{D}|=pq$, then $S=S_{D}$, $|U|=|G/L|=2pq$, and $|L|=|G/U|=2$. So $|W_{H}|=2$, where $W_{H}$ is the Sylow $2$-subgroup of $W$. Since $W$ is cyclotomic, $W_{H}$ is an $\mathcal{A}_{W}$-subgroup. Clearly, $\mathcal{A}_{W_{H}}=\mathbb{Z}W_{H}$. Together with Lemma 3.2(2), this implies that $\mathcal{A}_{W}\cong\mathbb{Z}W_{H}\otimes\mathcal{A}_{S}$. Therefore $\mathcal{A}_{L}$ is $\otimes$-complemented in $\mathcal{A}_{U}$ or $\mathcal{A}_{S}$ is $\otimes$-complemented in $\mathcal{A}_{G/L}$ and hence $\mathcal{A}$ is schurian by Lemma 3.7.
In view of the above paragraph, we may assume further that $|S_{D}|$ is a $p$-power. The Hall $2^{\prime}$-subgroup of $W$ is denoted by $W_{D}$. Note that $W_{D}$ is an $\mathcal{A}_{W}$-subgroup because $\mathcal{A}_{W}$ is cyclotomic.
Lemma 8.2.
With the above notation, $\mathcal{A}_{S_{D}}$ is a cyclotomic $S$-ring with trivial radical.
Proof.
One can see that $\mathcal{A}_{S_{D}}$ is cyclotomic because so is $\mathcal{A}_{W}$. Suppose that Statement $(3)$ of Theorem 7.1 holds. Then $|\operatorname{rad}(\mathcal{A}_{W})|=1$ and $|W_{H}|\leq 2$. Observe that $|\operatorname{rad}(\mathcal{A}_{W_{D}})|=1$. Indeed, Lemma 3.2(2) and Lemma 6.2 imply that $\mathcal{A}_{W}=\mathbb{Z}W_{H}\otimes\mathcal{A}_{W_{D}}$ and hence $|\operatorname{rad}(\mathcal{A}_{W_{D}})|=|\operatorname{rad}(\mathcal{A}_{W})|=1$. If $S_{D}=W_{D}$, then obviously $|\operatorname{rad}(\mathcal{A}_{S_{D}})|=1$. Otherwise, $|\operatorname{rad}(\mathcal{A}_{S_{D}})|=1$ by Lemma 6.9.
Now suppose that Statement $(4)$ of Theorem 7.1 holds. Then $W=U$, $\{e\}<L\leq D$, and $|\operatorname{rad}(\mathcal{A}_{W_{D}})|\in\{1,3\}$. As $\{e\}<L\leq D$, we conclude that $S_{D}\neq W_{D}$. Therefore $|\operatorname{rad}(\mathcal{A}_{S_{D}})|=1$ by Lemma 6.9. ∎
Lemma 8.3.
With the above notation, $\mathcal{A}_{S}$ is $2$-minimal unless $\mathcal{A}_{S}=\mathcal{A}_{S_{H}}\otimes\mathcal{A}_{S_{D}}$ and at least one of the following statements holds:
  1. (1)
    $|S_{H}|=4$ and $\mathcal{A}_{S_{H}}=\mathcal{T}_{S_{H}}$;
  2. (2)
    $|S_{D}|=p$ and $\mathcal{A}_{S_{D}}=\mathcal{T}_{S_{D}}$.
Proof.
Assume that $\mathcal{A}_{S}$ is not $2$-minimal. Clearly, $|S_{H}|\in\{1,2,4\}$. Let us prove that
(36) $\mathcal{A}_{S}=\mathcal{A}_{S_{H}}\otimes\mathcal{A}_{S_{D}}$
If $|S_{H}|\leq 2$, then $\mathcal{A}_{S_{H}}=\mathbb{Z}S_{H}$ and Eq. (36) follows from Lemma 3.2(2). If $|S_{H}|=4$ and Eq. (36) does not hold, then $\mathcal{A}_{S}$ is $2$-minimal by Theorem 7.12(1) and Lemma 8.2, a contradiction to the assumption.
As $\mathcal{A}_{S}$ is not $2$-minimal, Corollary 3.4 yields that at least one of the $S$-rings $\mathcal{A}_{S_{H}}$, $\mathcal{A}_{S_{D}}$ is not $2$-minimal. If $\mathcal{A}_{S_{H}}$ is not $2$-minimal, then Statement $(1)$ of the lemma holds by Lemma 2.6. If $\mathcal{A}_{S_{D}}$ is not $2$-minimal, then $\mathcal{A}_{S_{D}}$ is not normal by Lemma 6.5. Note that $|\operatorname{rad}(\mathcal{A}_{S_{D}})|=1$ by Lemma 8.2 and hence $\mathcal{A}_{S_{D}}=\mathcal{T}_{S_{D}}$ by Lemma 6.2. Since $\mathcal{A}_{S_{D}}$ is cyclotomic, we conclude that $|S_{D}|=p$. Thus, Statement $(2)$ of the lemma holds. ∎
It is worth to mention that the tensor product from Lemma 8.3 can be trivial, i.e. $S_{H}$ or $S_{D}$ can be trivial. In view of Lemma 3.6 and Lemma 8.3, we may assume further that
$\mathcal{A}_{S}=\mathcal{A}_{S_{H}}\otimes\mathcal{A}_{S_{D}}$
and Statement $(1)$ or $(2)$ of Lemma 8.3 holds. The canonical epimorphism from $G$ to $G/L$ is denoted by $\pi$ and the image of $X\subseteq G$ under $\pi$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu$.
Lemma 8.4.
If Statement $(3)$ of Theorem 7.1 holds, then $\mathcal{A}$ is schurian.
Proof.
By the condition of the lemma, $\mathcal{A}_{W}$ is a circulant cyclotomic $S$-ring with trivial radical. Clearly, $|W_{H}|\leq 2$ in this case. So $|S_{H}|\leq 2$. This implies that Statement $(2)$ of Lemma 8.3 holds for $\mathcal{A}_{S}$, i.e. $|S_{D}|=p$ and $\mathcal{A}_{S_{D}}=\mathcal{T}_{S_{D}}$. Let $V\in\{U,G/L\}\setminus\{W\}$. If $4\nmid|V|$, then $V$ is cyclic and 2||L|2~|~|L|. The latter yields that $2\nmid|S|$ and hence $|S|=|S_{D}|=p$. Therefore $\mathcal{A}$ is schurian by Lemma 6.6.
Now suppose that 4||V|4~|~|V|. Note that
$|V|=4p^{l}$
for some $l\geq 1$. This is obvious if $n=p^{k}$ and follows from $|S_{D}|=p$ and $V<G$ if $n=pq$. One can see that
$|S_{H}|=2$
because $4\nmid|W|$ and 4||V|4~|~|V|. Clearly, $\mathcal{A}_{S_{H}}=\operatorname{Cyc}(\operatorname{Aut}(S_{H}),S_{H})$ and $\mathcal{A}_{S_{D}}=\mathcal{T}_{S_{D}}=\operatorname{Cyc}(\operatorname{Aut}(S_{D}),S_{D})$. So
$\mathcal{A}_{S}=\mathcal{A}_{S_{H}}\otimes\mathcal{A}_{S_{D}}=\operatorname{Cyc}(\operatorname{Aut}(S_{H})\times\operatorname{Aut}(S_{D}),S)=\operatorname{Cyc}(\operatorname{Aut}(S),S)$
by Eq. (4). The latter is equivalent to
$\mathcal{A}_{S}=V(\operatorname{Hol}(S),S).$
As $\mathcal{A}_{W}$ is cyclotomic, there exists $K_{1}\leq\operatorname{Hol}(W)$ such that $\mathcal{A}_{W}=V(K_{1},W)$. It is easy to see that $K_{1}^{S}\leq\operatorname{Hol}(S)$ and $\mathcal{A}_{S}=V(K_{1}^{S},S)$. Since $\mathcal{A}_{S}$ is Cayley minimal (Lemma 6.4) and $\mathcal{A}_{S}=V(\operatorname{Hol}(S),S)$, we conclude that
(37) $K_{1}^{S}=\operatorname{Hol}(S).$
Let $\operatorname{Aut}(\mathcal{A}_{V})^{S_{D}}\leq\operatorname{Hol}(S_{D})$. Then $\operatorname{Aut}(\mathcal{A}_{V})^{S_{D}}=\operatorname{Hol}(S_{D})$ because $\operatorname{Hol}(S_{D})\in\mathcal{K}^{\min}(\mathcal{A}_{S_{D}})$ and $\mathcal{A}_{S_{D}}=V(\operatorname{Aut}(\mathcal{A}_{V})^{S_{D}},S_{D})$. Note that $\operatorname{Aut}(\mathcal{A}_{V})^{S_{H}}=\operatorname{Hol}(S_{H})$ because $|S_{H}|=2$. Therefore
$\operatorname{Aut}(\mathcal{A}_{V})^{S}=\operatorname{Aut}(\mathcal{A}_{V})^{S_{H}}\times\operatorname{Aut}(\mathcal{A}_{V})^{S_{D}}=\operatorname{Hol}(S_{H})\times\operatorname{Hol}(S_{D})=\operatorname{Hol}(S).$
Together with Eq. (37), this yields that $K_{1}^{S}=\operatorname{Aut}(\mathcal{A}_{V})^{S}$. Thus, $\mathcal{A}$ is schurian by Lemma 3.5.
In view of the above paragraph, we may assume that
$\operatorname{Aut}(\mathcal{A}_{V})^{S_{D}}\nleq\operatorname{Hol}(S_{D}).$
If $V=G/L$, then $\mathcal{A}_{V}$ is the $\mkern 1.5mu\overline{\mkern-1.5muU_{1}\mkern-1.5mu}\mkern 1.5mu/S_{D}$-wreath product for an $\mathcal{A}_{G/L}$-subgroup $\mkern 1.5mu\overline{\mkern-1.5muU_{1}\mkern-1.5mu}\mkern 1.5mu\geq S$ with $|\mkern 1.5mu\overline{\mkern-1.5muU_{1}\mkern-1.5mu}\mkern 1.5mu/S_{D}|\leq 4$ by Lemma 7.16(1) applied to $\mathcal{A}_{G/L}$, whereas if $V=U$, then $\mathcal{A}_{V}$ is the $U_{2}/L_{2}$-wreath product, where $U_{2}=S_{H}^{\pi^{-1}}$ and $L_{2}\leq L$ is an $\mathcal{A}_{U}$-subgroup with $|S_{H}^{\pi^{-1}}/L_{2}|\leq 4$, by Lemma 7.16(2) applied to $\mathcal{A}_{U}$.
Suppose that $|V|\neq 4p$. Then $\mathcal{A}$ is the nontrivial $S_{1}=U_{1}/L_{1}$-wreath product, where $U_{1}=\mkern 1.5mu\overline{\mkern-1.5muU_{1}\mkern-1.5mu}\mkern 1.5mu^{\pi^{-1}}$ and $L_{1}=(S_{D})^{\pi^{-1}}$, whenever $V=G/L$, and $\mathcal{A}$ is the nontrivial $S_{2}=U_{2}/L_{2}$-wreath product whenever $V=U$. As $|\mkern 1.5mu\overline{\mkern-1.5muU_{1}\mkern-1.5mu}\mkern 1.5mu/S_{D}|\leq 4$ ($|S_{H}^{\pi^{-1}}/L_{2}|\leq 4$, respectively), we obtain $|S_{1}|\leq 4$ ($|S_{2}|\leq 4$, respectively). If $|S_{i}|=4$, $i\in\{1,2\}$, then $\mathcal{A}_{S_{i}}\neq\mathcal{T}_{S_{i}}$ because $S_{D}$ is an $\mathcal{A}_{G/L}$-subgroup of order $2$. Therefore $\mathcal{A}_{S_{i}}$, $i\in\{1,2\}$, is $2$-minimal by Lemma 2.6. Thus, $\mathcal{A}$ is schurian by Lemma 3.6.
Now suppose that $|V|=4p$. If $\mathcal{A}_{V}$ is dense, then $\mathcal{A}$ is cyclotomic by [11, Case 33, p. 115]. Therefore $\mathcal{A}_{V}=V(K_{0},V)$ for some $K_{0}\leq\operatorname{Hol}(V)$. Clearly, $K_{0}^{S}\leq\operatorname{Hol}(S)$ and $\mathcal{A}_{S}=V(K_{0}^{S},S)$. Since $\mathcal{A}_{S}$ is Cayley minimal (Lemma 6.4) and $\mathcal{A}_{S}=V(\operatorname{Hol}(S),S)$, we conclude that $K_{0}^{S}=\operatorname{Hol}(S)$. Together with Eq. (37) and Lemma 3.5, this implies that $\mathcal{A}$ is schurian.
Let $\mathcal{A}$ be nondense. If $V=G/L$, then $\mathcal{A}_{V}=\mathcal{A}_{G/L}$ is the $S/S_{D}$-wreath product by [11, Case 11, p. 114] and hence $\mathcal{A}$ is the $S_{1}=U/L_{1}$-wreath product, where $L_{1}=(S_{D})^{\pi^{-1}}$. If $V=U$, then $\mathcal{A}_{V}=\mathcal{A}_{U}$ is the $S_{2}=H/L$-wreath product by [11, Case 22, p. 114] (in this case, $|L|=2$) and hence $\mathcal{A}$ is the $S_{2}$-wreath product. One can see that $|S_{1}|=|U/L_{1}|=|S/S_{D}|=|S_{H}|=2$ and $|S_{2}|=|H/L|=2$. Thus, $\mathcal{A}$ is schurian by Lemma 3.6. ∎
In view of Lemma 8.4, we may assume that Statement $(4)$ of Theorem 7.1 holds, i.e. $\mathcal{A}$ is dense, $U\geq H$, $\mathcal{A}_{U}$ is cyclotomic, $|\operatorname{rad}(\mathcal{A}_{U\cap D})|=1$ unless $n=3^{k}$ and $|\operatorname{rad}(\mathcal{A}_{U\cap D})|=3$, and $L\leq D$. In this case, $|S_{H}|=4$.
Lemma 8.5.
There exist $K_{1}\leq\operatorname{Aut}(\mathcal{A}_{U})$ and $K_{0}\leq\operatorname{Aut}(\mathcal{A}_{G/L})$ such that $K_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{U})$, $K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{G/L})$, and $K_{1}^{S_{H}}=K_{0}^{S_{H}}\in\mathcal{K}^{\min}(\mathcal{A}_{S_{H}})$.
Proof.
If $\mathcal{A}_{S_{H}}\neq\mathcal{T}_{S_{H}}$, then $\mathcal{A}_{S_{H}}$ is $2$-minimal by Lemma 2.6. So $\operatorname{Aut}(\mathcal{A}_{U})^{S_{H}}=\operatorname{Aut}(\mathcal{A}_{G/L})^{S_{H}}=\operatorname{Aut}(\mathcal{A}_{S_{H}})$ and the lemma holds for $K_{1}=\operatorname{Aut}(\mathcal{A}_{U})$ and $K_{0}=\operatorname{Aut}(\mathcal{A}_{G/L})$.
Further, we assume that $|S_{H}|=4$ and $\mathcal{A}_{S_{H}}=\mathcal{T}_{S_{H}}$. By Remark 2.7, there is a unique proper $2$-equivalent subgroup $\operatorname{Alt}(S_{H})\cong\operatorname{Alt}(4)$ of $\operatorname{Aut}(\mathcal{A}_{S_{H}})\cong\operatorname{Sym}(4)$ of order $12$ such that $\operatorname{Alt}(S_{H})\in\mathcal{K}^{\min}(\mathcal{A}_{S_{H}})$. If $\mathcal{A}_{U}\neq\mathcal{A}_{H}\otimes\mathcal{A}_{U\cap D}$, then put $K_{1}=\operatorname{Aut}(\mathcal{A}_{U})$. In this case, $K_{1}^{S_{H}}=\operatorname{Alt}(S_{H})$ by Lemma 7.14. Otherwise, $\operatorname{Aut}(\mathcal{A}_{U})=\operatorname{Aut}(\mathcal{A}_{H})\times\operatorname{Aut}(\mathcal{A}_{U\cap D})$. Let $K_{H}\leq\operatorname{Sym}(H)$ be such that $K_{H}\cong\operatorname{Alt}(S_{H})$. The group $K_{H}$ is $2$-equivalent to $\operatorname{Aut}(\mathcal{T}_{H})$ (see Remark 2.7) and hence $K_{1}=K_{H}\times\operatorname{Aut}(\mathcal{A}_{U\cap D})$ is $2$-equivalent to $\operatorname{Aut}(\mathcal{A}_{U})$. By the definition of $K_{1}$,
$K_{1}^{S_{H}}=(K_{H}\times\operatorname{Aut}(\mathcal{A}_{U\cap D}))^{S_{H}}=(K_{H})^{S_{H}}=\operatorname{Alt}(S_{H}).$
The groups $S_{H}$ and $D/L$ are $\mathcal{A}_{G/L}$-subgroups. If $\mathcal{A}_{G/L}\neq\mathcal{A}_{S_{H}}\otimes\mathcal{A}_{D/L}$, then put $K_{0}=\operatorname{Aut}(\mathcal{A}_{G/L})$; otherwise put $K_{0}=\operatorname{Alt}(S_{H})\times\operatorname{Aut}(\mathcal{A}_{D/L})$. The argument which is the similar to the argument from the previous paragraph implies that $K_{0}$ is $2$-equivalent to $\operatorname{Aut}(\mathcal{A}_{G/L})$ and $K_{0}^{S_{H}}=\operatorname{Alt}(S_{H})=K_{1}^{S_{H}}$ as required. ∎
Let $K_{1}$ and $K_{0}$ are defined as in Lemma 8.5.
Lemma 8.6.
With the above notation, $K_{1}^{S_{D}}=K_{0}^{S_{D}}\in\mathcal{K}^{\min}(\mathcal{A}_{S_{D}})$ unless $\mathcal{A}$ is the $(H\times L)/L$- or $U/(U\cap D)$-wreath product.
Proof.
By Lemma 6.2 and Lemma 8.2, the $S$-ring $\mathcal{A}_{S_{D}}$ is normal or trivial over a group of order $p$. If $\mathcal{A}_{S_{D}}$ is normal, then $\mathcal{A}_{S_{D}}$ is $2$-minimal by Lemma 6.5. If $|S_{D}|=3$, then $\mathcal{A}_{S_{D}}$ is obviously $2$-minimal. So $K_{1}^{S_{D}}=K_{0}^{S_{D}}=\operatorname{Aut}(\mathcal{A}_{S_{D}})\in\mathcal{K}^{\min}(\mathcal{A}_{S_{D}})$ in both these cases as required. Further, we assume that
$|S_{D}|=p\geq 5~\text{and}~\mathcal{A}_{S_{D}}=\mathcal{T}_{S_{D}}.$
This assumption implies that $|(U\cap D):L|=p$.
If $\operatorname{Aut}(\mathcal{A}_{U})^{S_{D}}\leq\operatorname{Hol}(S_{D})$ and $\operatorname{Aut}(\mathcal{A}_{G/L})^{S_{D}}\leq\operatorname{Hol}(S_{D})$, then $K_{1}^{S_{D}}\leq\operatorname{Hol}(S_{D})$ and $K_{0}^{S_{D}}\leq\operatorname{Hol}(S_{D})$. Since $\operatorname{Hol}(S_{D})\in\mathcal{K}^{\min}(\mathcal{A}_{S_{D}})$, we obtain $K_{1}^{S_{D}}=K_{0}^{S_{D}}=\operatorname{Hol}(S_{D})$ as desired.
If $\operatorname{Aut}(\mathcal{A}_{U})^{S_{D}}\nleq\operatorname{Hol}(S_{D})$ or $\operatorname{Aut}(\mathcal{A}_{G/L})^{S_{D}}\nleq\operatorname{Hol}(S_{D})$, then $\mathcal{A}_{U}$ is the $(H\times L)/L$-wreath product or $\mathcal{A}_{G/L}$ is the $S/S_{D}$-wreath product, respectively, by Lemma 7.16. In the former case, $\mathcal{A}$ is the $(H\times L)/L$-wreath product, whereas in the latter one, $\mathcal{A}$ is the $U/(U\cap D)$-wreath product and we are done. ∎
Return to the proof of Theorem 1.2. By Lemma 8.5, we have $K_{1}^{S_{H}}=K_{0}^{S_{H}}\in\mathcal{K}^{\min}(\mathcal{A}_{S_{H}})$. If $\mathcal{A}$ is not the $T$-wreath product, where $T=U_{1}/L_{1}\in\{(H\times L)/L,U/(U\cap D)\}$, then Lemma 8.6 implies that $K_{1}^{S_{D}}=K_{0}^{S_{D}}\in\mathcal{K}^{\min}(\mathcal{A}_{S_{D}})$. Therefore $K_{1}^{S}=K_{0}^{S}$ by Lemma 3.3. Thus, $\mathcal{A}$ is schurian by Lemma 3.5.
Suppose that $\mathcal{A}$ is the $T$-wreath product. Clearly, $|T|=|H|=4$ and Statement $(4)$ of Theorem 7.1 holds for $\mathcal{A}$ and $T$. By Lemma 8.5, there are $M_{1}\leq\operatorname{Aut}(\mathcal{A}_{U_{1}})$ and $M_{0}\leq\operatorname{Aut}(\mathcal{A}_{G/L_{1}})$ such that $M_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{U_{1}})$, $M_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{G/L_{1}})$, and $M_{1}^{T}=M_{0}^{T}$. Again, $\mathcal{A}$ is schurian by Lemma 3.5.
9. Proof of Theorem 1.4
Since a section of a Schur group is a Schur group, to prove Theorem 1.4, it suffices to prove that the group
$G\cong C_{8}\times C_{2}\times C_{p},$
where $p$ is an odd prime, is not a Schur group. Let $A$, $B$, and $C$ be subgroups of $G$ such that
$A\cong C_{8},~B\cong C_{2},~P\cong C_{p},~\text{and}~G=A\times B\times P.$
Denote generators of $A$ and $B$ by $a$ and $b$, respectively. Put
$A_{1}=\langle a^{2}\rangle\cong C_{4},~A_{2}=\langle a^{4}\rangle\cong C_{2},~U_{1}=A_{1}\times P\cong C_{4p},~\text{and}~U=A_{1}\times B\times P=B\times U_{1}\cong C_{2}\times C_{4p}.$
Let us construct a nonschurian $S$-ring over $G$. The group $\operatorname{Aut}(P)$ is a cyclic group of order p1p-1. Let $M_{0}$ be the subgroup of $\operatorname{Aut}(P)$ of index $2$. The canonical epimorphism from $\operatorname{Aut}(P)$ to $\operatorname{Aut}(P)/M_{0}$ is denoted by $\pi$. Since $\operatorname{Aut}(A_{1})$ is a group of order $2$, there exists a unique isomorphism $\varphi$ from $\operatorname{Aut}(A_{1})\cong C_{2}$ to $\operatorname{Aut}(P)/M_{0}\cong C_{2}$. Put
$\mathcal{A}_{1}=\operatorname{Cyc}(M_{1},U_{1}),$
where $M_{1}=\{(\delta,\tau)\in\operatorname{Aut}(A_{1})\times\operatorname{Aut}(P):~\delta^{\varphi}=\tau^{\pi}\}\leq\operatorname{Aut}(U_{1})$, and
$\mathcal{A}_{2}=\mathcal{T}_{B}\otimes\mathcal{T}_{P}.$
One can see that $P$ is an $\mathcal{A}_{1}$- and $\mathcal{A}_{2}$-section and $(\mathcal{A}_{1})_{P}=(\mathcal{A}_{2})_{P}=\mathcal{T}_{P}$. Therefore one can form the $S$-ring
$\mathcal{A}_{12}=\mathcal{A}_{1}\wr_{U_{1}/A_{1}}\mathcal{A}_{2}$
over $U$.
Put
$\mathcal{A}_{3}=\operatorname{Cyc}(\langle\delta_{0}\rangle,A),$
where $\delta_{0}\in\operatorname{Aut}(A)$ is such that $a^{\delta_{0}}=a^{-1}$. It is a straightforward to check that the partition of $(A/A_{2})\times B\cong C_{4}\times C_{2}$ into the sets
$\{A_{2}\},~\{A_{2}a^{2}\},~\{A_{2}ab\},~\{A_{2}a^{3}b\},~\{A_{2}a,A_{2}a^{3}\},~\{A_{2}b,A_{2}a^{2}b\}$
defines the $S$-ring $\mathcal{A}_{4}$ over $(A/A_{2})\times B$ such that
$\mathcal{A}_{4}\cong\mathbb{Z}C_{4}\wr_{C_{2}}\mathbb{Z}E_{4}.$
One can see that $A/A_{2}$ is an $\mathcal{A}_{3}$- and $\mathcal{A}_{4}$-section and $(\mathcal{A}_{3})_{A/A_{2}}=(\mathcal{A}_{4})_{A/A_{2}}\cong\mathbb{Z}C_{2}\wr\mathbb{Z}C_{2}$. Therefore one can form the $S$-ring
$\mathcal{A}_{34}=\mathcal{A}_{3}\wr_{A/A_{2}}\mathcal{A}_{4}$
over $A\times B$.
The section $U/P=A_{1}\times B$ is an $\mathcal{A}_{12}$- and $\mathcal{A}_{34}$-section and
$(\mathcal{A}_{12})_{U/P}=(\mathcal{A}_{34})_{U/P}=(\mathbb{Z}A_{2}\wr\mathbb{Z}(A_{1}/A_{2}))\wr\mathbb{Z}B\cong(\mathbb{Z}C_{2}\wr\mathbb{Z}C_{2})\wr\mathbb{Z}C_{2}.$
Therefore one can form the $S$-ring
$\mathcal{A}=\mathcal{A}_{12}\wr_{U/P}\mathcal{A}_{34}$
over $G$. The basic sets of $\mathcal{A}$ are the following:
$X_{0}=\{e\},~X_{1}=\{a^{4}\},~X_{2}=\{a^{2},a^{6}\},~X_{3}=P^{\#},~X_{4}=a^{4}P^{\#},$
$Y_{1}=a^{2}P_{1}\cup a^{6}P_{2},~Y_{2}=a^{2}P_{2}\cup a^{6}P_{1},$
$Z_{1}=A_{1}b,~Z_{2}=A_{1}bP^{\#},$
$T_{1}=\{a,a^{7}\}P,~T_{2}=\{a^{3},a^{5}\}P,~T_{3}=\{ab,a^{5}b\}P,~T_{4}=\{a^{3}b,a^{7}b\}P,$
where $P_{1}$ and $P_{2}$ are the nontrivial orbits of $M_{0}$ on $P$.
Proposition 9.1.
The $S$-ring $\mathcal{A}$ is nonschurian.
Proof.
Assume the contrary. Then
(38) $\mathcal{S}(\mathcal{A})=\operatorname{Orb}(K,G),$
where $K=\operatorname{Aut}(\mathcal{A})_{e}$. Let $x_{1}\in P_{1}$ and $x_{2}\in P_{2}$. Since the elements $bx_{1}$ and $bx_{2}$ lie in the same basic set $Z_{2}$ of $\mathcal{A}$, there exists $f\in K$ such that $(bx_{1})^{f}=bx_{2}$. From Eqs. (2) and (38) it follows that
$(P^{\#}\setminus\{x_{1}\})^{f}=(A_{1}P^{\#}x_{1}\cap P^{\#})^{f}=(Z_{2}bx_{1}\cap X_{3})^{f}=Z_{2}bx_{2}\cap X_{3}=A_{1}P^{\#}x_{2}\cap P^{\#}=P^{\#}\setminus\{x_{2}\}.$
Since $X_{3}=P^{\#}$ is an orbit of $K$, we have
$x_{1}^{f}=x_{2}.$
The above equality together with Eqs. (2) and (38) imply that
$((a^{2}P_{1}x_{1}\cup a^{6}P_{2}x_{1})\cap\{a^{2},a^{6}\})^{f}=(Y_{1}x_{1}\cap X_{2})^{f}=Y_{1}x_{2}\cap X_{2}=(a^{2}P_{1}x_{2}\cup a^{6}P_{2}x_{2})\cap\{a^{2},a^{6}\}.$
If $|M_{0}|$ is even, then $P_{1}=P_{1}^{-1}$ and $P_{2}=P_{2}^{-1}$ and hence the left-hand side and the right-hand side of the latter equality are equal to $\{a^{2}\}$ and $\{a^{6}\}$, respectively. If $|M_{0}|$ is odd, then $P_{2}=P_{1}^{-1}$ and hence the left-hand side and the right-hand side of the latter equality are equal to $\{a^{6}\}$ and $\{a^{2}\}$, respectively. Since $X_{2}=\{a^{2},a^{6}\}\in\operatorname{Orb}(K,G)$, we have
(39) $(a^{2})^{f}=a^{6}$
in both cases.
Now by Eqs. (2), (38), and (39), we have
(40) $(Pa)^{f}=(T_{1}a^{2}\cap T_{1})^{f}=T_{1}a^{6}\cap T_{1}=Pa^{7}.$
On the other hand, since $(bx_{1})^{f}=bx_{2}$, we have
$(Pa)^{f}=(T_{3}bx_{1}\cap T_{1})^{f}=T_{3}bx_{2}\cap T_{1}=Pa,$
a contradiction to Eq. (40). ∎
Theorem 1.4 follows from Proposition 9.1.
10. Proof of Theorem 1.5
Let $p$ be an odd prime, $P\cong C_{p}$, $H\cong E_{16}$, and $G=H\times P$. If $p=3$, then $G$ is a Schur group by the computational results [35]. Further, we assume that $p\geq 5$. In this case, we are going to construct a nonschurian $S$-ring over $G$ and thereby to prove Theorem 1.5.
Let $a,b,c,d\in H$ be such that $\langle a,b,c,d\rangle=H$, $A=\langle a\rangle$, $B=\langle b\rangle$, $C=\langle c\rangle$, and let $\delta_{0}\in\operatorname{Aut}(H)$ be such that
$\delta_{0}:(a,b,c,d)\mapsto(a,ab,bc,cd),$
$M_{0}=\langle\delta_{0}\rangle$, and
$\mathcal{A}_{0}=\operatorname{Cyc}(M_{0},H).$
It is straightforward to verify that $|\delta_{0}|=4$ and the basic sets of $\mathcal{A}_{0}$ are the following:
$\{e\},~\{a\},~Ab,~(A\times B)c,~\{e,b,c,abc\}d,~\{a,ab,ac,bc\}d.$
One can see that each of two latter basic sets has trivial radical and generates $H$. So $\mathcal{A}_{0}$ is indecomposable. Every basic set of $\mathcal{A}_{0}$ has the $2$-power size, i.e. $\mathcal{A}_{0}$ is a 22-$S$-ring in terms of [19]. Clearly, $\mathcal{A}_{0}$ is cyclotomic and hence schurian. Therefore $\mathcal{A}_{0}$ is $2$-minimal by [19, Lemma 5.6] stating that every schurian indecomposable $p$-$S$-ring over an elementary abelian group of order $p^{4}$ is $2$-minimal for every prime $p$. Thus,
(41) $K_{0}=H_{r}\rtimes M_{0}$
for every $K_{0}\leq\operatorname{Sym}(H)$ such that $K_{0}\geq H_{r}$ and $K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{0})$ (recall that $H_{r}$ is the group of $\operatorname{Sym}(H)$ induced by all right multiplications of $H$). In particular, $\operatorname{Aut}(\mathcal{A}_{0})=H_{r}\rtimes M_{0}$.
Let $V=A\times B\times C$ and $U=V\times P$. As $p\geq 5$, the group $M_{P}=\operatorname{Aut}(P)\cong C_{p-1}$ has a nontrivial subgroup $M_{P}^{0}$ of index $2$. Let $\delta_{1},\delta_{2}\in\operatorname{Aut}(V)$ be such that
$\delta_{1}:(a,b,c)\mapsto(a,ab,bc),~\delta_{2}:(a,b,c)\mapsto(a,b,bc).$
Put $M_{V}=\langle\delta_{1},\delta_{2}\rangle$. It can be verified directly that $|\delta_{1}|=4$, $|\delta_{2}|=2$, and $\delta_{2}\delta_{1}\delta_{2}=\delta_{1}^{-1}$. So $M_{V}\cong D_{8}$.
Let $M^{0}_{V}=\langle\delta_{1}^{2},\delta_{2}\delta_{1}\rangle$. Clearly, $M_{V}^{0}\cong E_{4}$ and hence $M_{V}/M_{V}^{0}\cong C_{2}$. Put
$M_{1}=\{(\delta,\tau)\in M_{V}\times M_{P}:~(\delta^{\pi_{1}})^{\psi}=\tau^{\pi_{2}}\}\leq\operatorname{Aut}(U),$
where $\pi_{1}$ and $\pi_{2}$ are the canonical epimorphisms from $M_{V}$ to $M_{V}/M_{V}^{0}$ and from $M_{P}$ to $M_{P}/M_{P}^{0}$, respectively, and $\psi$ is a unique isomorphism from $M_{V}/M_{V}^{0}\cong C_{2}$ to $M_{P}/M_{P}^{0}\cong C_{2}$, and
$\mathcal{A}_{1}=\operatorname{Cyc}(M_{1},U).$
The definition of $M_{1}$ implies that $V$ and $P$ are $\mathcal{A}_{1}$-subgroups and the basic sets of $\mathcal{A}_{1}$ are the following:
$\{e\},~\{a\},~Ab,~(A\times B)c,~P^{\#},~aP^{\#},~AbP^{\#},$
$AcP_{1}\cup AbcP_{2},~AcP_{2}\cup AbcP_{1},$
where $P_{1}$ and $P_{2}$ are the nontrivial orbits of $M_{P}^{0}$ on $P$.
One can see that $V$ is an $\mathcal{A}_{0}$- and $\mathcal{A}_{1}$-subgroup and $(\mathcal{A}_{0})_{V}=(\mathcal{A}_{1})_{V}\cong(\mathbb{Z}C_{2}\wr\mathbb{Z}C_{2})\wr\mathbb{Z}C_{2}$. Therefore one can form the $S$-ring
$\mathcal{A}=\mathcal{A}_{1}\wr_{S}\mathcal{A}_{0}$
over $G$, where $S=U/P$ and $\mathcal{A}_{0}$ is considered as an $S$-ring over $G/P\cong H$.
To complete the proof of Theorem 1.5, it suffices to prove the proposition below.
Proposition 10.1.
The $S$-ring $\mathcal{A}$ is nonschurian.
Proof.
Assume the contrary that $\mathcal{A}$ is schurian. Then there exist two groups $K_{1}\leq\operatorname{Sym}(U)$ and $K_{0}\leq\operatorname{Sym}(G/P)$ such that
$K_{1}\geq U_{r},~K_{0}\geq(G/P)_{r},~K_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{1}),~K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{0}),~\text{and}~K_{0}^{S}=K_{1}^{S}$
by Lemma 3.5. From Eq. (41) it follows that
(42) $K_{0}^{S}=S_{r}\rtimes M_{0}^{S}.$
Let $X\in\mathcal{S}(\mathcal{A}_{U})=\mathcal{S}(\mathcal{A}_{1})$ be such that
$X=AcP_{1}\cup AbcP_{2}$
and $x\in P_{1}$. Suppose that $f$ belongs to the one-point stabilizer $M_{x}$ of $x$ in $M$, where $M=(K_{1})_{e}$ is the one-point stabilizer of $e$ in $K_{1}$. Since $K_{0}^{S}=K_{1}^{S}$, Eq. (42) implies that $f^{S}\in M_{0}^{S}$. In view of Eq. (2) and $(A\times B)c,X^{-1}\in\mathcal{S}(\mathcal{A})$, we obtain
$(Ac)^{f}=(X^{-1}x\cap(A\times B)c)^{f}=X^{-1}x\cap(A\times B)c=Ac.$
Together with $f^{S}\in M_{0}^{S}$ and the definition of $M_{0}$, this shows that $f^{S}\in\langle\delta_{0}^{2}\rangle$ and hence $b^{f}=b$. Since $H$ and $P$ are $\mathcal{A}$-subgroups, $Hx$ and $Pb$ are blocks of $\operatorname{Aut}(\mathcal{A})$. As $x^{f}=x$ and $b^{f}=b$, we conclude that $(Hx)^{f}=Hx$ and $(Pb)^{f}=Pb$. Therefore
$(bx)^{f}=(Hx\cap Pb)^{f}=Hx\cap Pb=bx.$
Thus, $f$ belongs to the one-point stabilizer $M_{bx}$ of $bx$ in $M$ and consequently
(43) $M_{x}\leq M_{bx}.$
The definition of $M_{1}$ implies that the basic set of $\mathcal{A}$ containing $x$ is $P^{\#}$ and the basic set of $\mathcal{A}$ containing $bx$ is $Y=AbP^{\#}$. Clearly, |Y|=2(p1)|Y|=2(p-1). By the assumption, we have $P^{\#},Y\in\operatorname{Orb}(M,U)$. Therefore
$|M_{x}|=\frac{|M|}{p-1}>\frac{|M|}{2(p-1)}=\frac{|M|}{|Y|}=|M_{bx}|,$
in contrast to Eq. (43). ∎
Acknowledgment: The author is very grateful to Prof. Ilia Ponomarenko for a careful reading and valuable comments which help to improve the text significantly. The author would like to thank Dr. Matan Ziv-Av for the help with computer calculations.